### v4

# The Cone Restriction: An old Approach Revisited

Abstract
We consider Ou-Wang’s approach to cone restriction via polynomial partitioning. By restructuring their induction arguments into a recursive algorithm and applying the nested polynomial Wolff axioms, we refine the bounds on cone restriction estimate in higher dimensions.
1 Introduction
In this paper, we present a small improvement on cone restriction in higher dimensions. The Fourier restriction is one of the most significant open challenges in harmonic analysis, addresses the restriction of the Fourier transform to a compact manifold. Stein [1] conjectured that the restriction of Fourier transforms to a well-curved surface, such as sphere, paraboloids, or the cone studied in this paper, is a bounded $L^{p^{\prime}}\to L^{q^{\prime}}$ operator for some $p^{\prime}$ and $q^{\prime}$. The precise statement of cone restriction conjecture can be formulated as follows.
Conjecture 1.1.
Let $n\geq 3$ and $\mathcal{C}$ be the truncated cone in $\mathbb{R}^{n}$ defined by
$\mathcal{C}:=\{(\xi_{1},\xi_{2},...,\xi_{n}):\xi_{1}^{2}+\xi_{1}^{3}+...+\xi_{n-1}^{2}=\xi_{n}^{2}\}$ (1.1)
Then, for all $p,q>0$ satisfy
$p>\frac{2(n-1)}{n-2},\frac{n-2}{q^{\prime}}\geq\frac{n}{p},\frac{1}{p}+\frac{1}{p^{\prime}}=\frac{1}{q}+\frac{1}{q^{\prime}}=1$ (1.2)
there holds
$\|\hat{f}\|_{L^{q^{\prime}}(\mathcal{C};d\sigma)}\lesssim_{p,q}\|f\|_{p^{\prime}},\forall f:\mathbb{R}^{n}\to\mathbb{C}$
where $d\sigma$ denotes the Riemannian measure on $\mathcal{C}$.
Let $Ef$ be the associated Fourier extension operator on $\mathcal{C}$, i.e.,
[…]
By a short duality argument, we see that conjecture 1.1 is equivalent to
Conjecture 1.2.
Let $n\geq 3$. For all $p,q>0$ satisfy (1.2), there holds
$\|Ef\|_{p}\lesssim_{n,p,q}\|f\|_{q},\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$ (1.3)
Conjecture 1.2 has so far been proven only in dimensions $n=3,4,5$ with contributions from Barcelo [2], Wolff [3], and Ou-Wang [4]. Prior to this work, the best-known range for higher dimensions was established by Ou-Wang [4], with results as follows.
Theorem 1.3 ([4], Theorem 1).
For $n\geq 3$, (1.3) holds whenever
$q\geq p>\left\{\begin{array}[]{rcl}4&\mbox{if}&n=3\\ 2\frac{3n+1}{3n-3}&\mbox{if}&n>3\quad odd\\ 2\frac{3n}{3n-4}&\mbox{if}&n>3\quad even\end{array}\right.$ (1.4)
Theorem 1.4 ([4], Theorem 2).
For $n\geq 3$, (1.3) holds whenever the tuple $(p,q,k)$ is admissible in the sense that $2<q<p$, $2\leq k\leq n$ and
$\left\{\begin{array}[]{rcl}p>2\frac{n+2}{n},q^{\prime}\leq\frac{n-2}{n}p&\mbox{if}&k=2\\ p>2\frac{n+k}{n+k-2},p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\mbox{if}&k\geq 3\end{array}\right.$
The Fourier restriction problem has been the focus of extensive study for decades. For hypersurfaces with non-vanishing Gaussian curvature, such as the sphere and paraboloid, restriction estimates are closely tied to several central problems in harmonic analysis, discrete geometry, and PDE, including the Bochner–Riesz conjecture, the Kakeya conjecture, and the Schrödinger maximal estimate. Furthermore, Fourier restriction has applications beyond harmonic analysis, impacting fields such as geometric measure theory (e.g., Falconer’s distance set conjecture) and number theory (e.g., counting solutions to Diophantine equations). However, sharp restriction estimates are known only for very few surfaces and dimensions. For instance, the restriction conjecture for the paraboloid and sphere remains unsolved for $n\geq 3$ It is known that the restriction conjecture for paraboloids or spheres in $\mathbb{R}^{n}$ implies the corresponding conjecture for cones in $\mathbb{R}^{n+1}$ (see [5]), which highlights the potential applications of our results.
In this work, we show that:
Theorem 1.5.
For $n\geq 3$,
$\|Ef\|_{L^{p}(B_{R})}\lesssim_{p,\epsilon}R^{\epsilon}\|f\|_{2}^{2/p}\|f\|_{\infty}^{1-2/p}$ (1.5)
whenever
$p>2+\lambda n^{-1}+O(n^{-2})$ (1.6)
where $\lambda=2.596…$ and $B_{R}$ is an arbitrary ball in $\mathbb{R}^{n}$ of radius $R$.
By a real interpolation of restrict-type and a standard $\epsilon$-removal argument (see [6]), we see that
Corollary 1.6.
For $n\geq 3$, (1.3) holds whenever $q\geq p$ and (1.6) holds.
Guth [8, 9] introduced polynomial partitioning to the study of Fourier restriction and established the improved results for paraboloids. Later, his results were subsequently refined by Wang[10], Wang-Wu[11], Hickman–Rogers [12], Hickman–Zahl [13], and Guo-Wang-Zhang [14]. In [12, 13] the authors reformulated Guth’s induction-on-scale argument into a recursive algorithm, utilizing various Kakeya-type geometric results from [15, 16, 17], including the nested polynomial Wolff axiom, to bound the number of directions of the tubes those stay in varieties. Beyond paraboloid restriction, polynomial partitioning has also been applied to many related problems in harmonic analysis, such as Bochner-Riesz conjecture (see [18, 19]), the Schrödinger maximal estimate (see [20]), and cone restriction (see [4]).
In this article, we refine the results presented in [4] to establish an improved $k$-broad estimate for the cone by adapting the methods used in [12, 13]. Specifically, we will employ iterated polynomial partitioning and the nested polynomial Wolff axiom in Section 3, where we reformulate the core of the inductive proofs from [4] as a recursive process. Compared to the paraboloid case addressed in [12, 13], the main difference in our approach is as follows: in the paraboloid case, a leaf backtracks to the root. However, if we were to apply this strategy to the cone, the number of tube directions would not be well-bounded, as the number of sectors on a cone is significantly fewer than the number of caps on a paraboloid. Instead, we make a leaf backtrack to its $(n-1)$- dimensional ancestor. We will provide a more detailed discussion on this topic in Section 3.
This article is not self-contained and heavily references the work of Guth [9], Ou-Wang [4], Hickman-Rogers [12], and Hickman-Zahl [13]. In Section 2, we review some of the tools employed in this work, including the $k$-broad norm, wave packet decomposition, polynomial partitioning, transverse equidistribution estimates, and the nested polynomial Wolff axiom. In Section 3, we present a modified version of the polynomial structural decomposition described in [12, 13] by applying iterated polynomial partitioning and the nested polynomial Wolff axiom. Finally, Section 4 discusses the numerology of Theorem 1.5.
[…]
Acknowledgements The author would like to thank Xiaochun Li for suggesting the problem and for numerous insightful discussions. The author is also grateful to Larry Guth and Shunhua Shizuo for their valuable guidance in learning decoupling theory.
2 Preliminaries
2.1 Notational Conventions
In this paper, we work with Schwartz function $f$ or $g$ supported in $2\bar{B}^{n-1}\setminus B^{n-1}$. We write $x\in\mathbb{R}^{n}$ as $x=(x^{\prime},x_{n})$. The parameters $n,k,\epsilon$ are fixed. All implicit constants are allowed to depend on $n,k,\epsilon$. We inductively define $\delta$ and $\delta_{j}$ for $0\leq j\leq n$ as bellows:
$\delta:=\epsilon^{10},\quad\delta_{j}:=\delta_{j-1}^{10}\quad\mbox{for}\quad 1\leq j\leq n,\quad\delta:=\delta_{n}^{10}$
In addition, we say $A\lesssim_{L}B$, $B\gtrsim_{L}A$, or $A=O_{L}(B)$ if $A\leq C_{L}B$ for some constant $C_{L}$ depending only on $L$. We say $A\sim_{L}B$ if both $A\lesssim_{L}B$ and $A\gtrsim_{L}B$ hold. We say $A\ll B$ or $B\gg A$ if $CA\leq B$ for some choice of $C\geq 1$ that can be taken to be as large as desired provided it is admissible.We say $A\lessapprox B$ if $A\lesssim R^{O(\delta_{0})}B$. Let #$\#A$ denote the cardinality of a finite set $A$. Let $N_{r}E$ denote the $r$-neighbourhood of a set $E$. Let $B_{r}$ denote an arbitrary ball in $\mathbb{R}^{n}$ of radius $r$. Let $T_{x}Z$ denote the tangent space of $Z$ at $x$. We say $A^{\prime}\subseteq A$ is a refinement of $A$ if #$\#A\lessapprox\#A^{\prime}$. We say $A$ = RapDec($r$) if $A\lesssim_{N}r^{-N}$. We say $f$ is essentially supported in $\Omega$ if $f(x)\lesssim$ RapDec($R$) $\|f\|_{2}$. We say $Ef$ is essentially supported in $\Omega$ if $Ef(x)\lesssim$ RapDec($R$) $\|f\|_{2}$.
2.2 Wave Packet Decomposition
The wave packet decomposition is a crucial tool in our argument. Here we recall its definition and basic properties from section 2.2 and 5.2.1 of [4].
[…]
Figure 1: A sector of a truncated cone.
Given a function $f$ with supp $f\subseteq 2\bar{B}^{n-1}\setminus B^{n-1}$, one writes $f=\sum_{\theta\in\Theta[r]}f_{\theta}$ where $f_{\theta}=f\psi_{\theta}$ for $\psi_{\theta}$ a smooth partition of unity subordinate to the cover $\theta$. For each $\theta$, we break up $f_{\theta}$ as follows. First, we decompose $\mathbb{R}^{n-1}$ into finitely overlapping cubes of side length  $r^{1/2+\delta}$, centered at $v\in r^{1/2+\delta}Z^{n-1}$. Now, we can write $\hat{f_{\theta}}=\sum_{v}\hat{f_{\theta}}\eta_{v}$ , where $\eta_{v}$ is a smooth partition of unity subordinate to this cover. Second, for each $v$, we decompose $\hat{f_{\theta}}\eta_{v}$ into even finer pieces according to $\theta$. Cover the cube centered at $v$ by parallel tine plates $P_{\theta,v}^{l}$ with radius  $r^{1/2+\delta}$, thickness  $r^{\delta/2}$, and normal direction $\xi_{\theta}$. Here $\xi_{\theta}$ is the point on the central line of a sector $\theta$ with $|\xi_{\theta}|=1$. Hence, one can write
$f=\sum_{\theta,v,l\in\mathbb{T}[r]}(\hat{f_{\theta}}\eta_{v}\eta_{\theta,v}^{l})^{\vee}$
where $\eta_{\theta,v}^{l}$ is a smooth partition of unity subordinate to this cover, and $\mathbb{T}[r]$ denotes the set of wave packets of scale $r$. Since $(\hat{f_{\theta}}\eta_{v}\eta_{\theta,v}^{l})^{\vee}$ is essentially supported in $\theta$ (Strictly speaking, this should be $C\theta$ for some constant $C$. However, it is convenient to omit the constant,as it does not impact the main arguments. Therefore, we will continue to do so in the following discussion.), we can choose bump functions $\psi_{\theta}^{\prime}$ so that $\psi_{\theta}^{\prime}=1$ on a $r^{-1/2}$ neighborhood of supp $\psi_{\theta}$, and decays rapidly outside this region. Finally, we define
$f_{\theta,v}^{l}:=\psi_{\theta}^{\prime}(\hat{f_{\theta}}\eta_{v}\eta_{\theta,v}^{l})^{\vee}$
and then
$f=\sum_{(\theta,v,l)\in\mathbb{T}[r]}f_{\theta,v}^{l}+\mbox{RapDec}(R)\|f\|_{2}$ (2.1)
For any $\mathcal{T}\subseteq\mathbb{T}[r]$, define
$f_{\mathcal{T}}:=\sum_{(\theta,v,l)\in\mathcal{T}}f_{\theta,v}^{l}$
We say $f$ is concentrated on wave packets from $\mathcal{T}\subseteq\mathbb{T}[r]$ if
$\|f_{\mathbb{T}[r]\setminus\mathcal{T}}\|_{\infty}\lesssim\mbox{RapRec}(R)\|f\|_{2}$
[…]
$L^{2}$-OrthogonalityBy Plancherel theorem, the functions $f_{\theta,v}^{l}$ are almost orthogonal in $L^{2}$.
Lemma 2.1 ([4], Section 2.2).
Let $r\geq 1$, then for any set $\mathcal{T}\subseteq\mathbb{T}[r]$,
$\|f_{\mathcal{T}}\|_{2}^{2}\sim\sum_{(\theta,v,l)\in\mathcal{T}}\|f_{\theta,v}^{l}\|_{2}^{2}$
Like the case of paraboloid [12], there is also some local versions of the $L^{2}$-orthogonality for cone, which are a direct corollary of lemma 2.1:
Corollary 2.2.
Let $1\leq\rho\leq r$, then for any set $\mathcal{T}\subseteq\mathbb{T}[r]$ and $\rho^{-1/2}$-sector $\theta^{*}$,
$\|\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{f^{l}_{\theta,v}}\|^{2}_{L^{2}\left({\theta}^{*}\right)}\lesssim\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{{\left\|f^{l}_{\theta,v}\right\|}^{2}_{L^{2}\left(C{\theta}^{*}\right)}}$
$\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{{\left\|f^{l}_{\theta,v}\right\|}^{2}_{L^{2}\left({\theta}^{*}\right)}}\lesssim{\left\|\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{f^{l}_{\theta,v}}\right\|}^{2}_{L^{2}\left(C{\theta}^{*}\right)}$
for some constant $C$.
Corollary 2.3.
Let $1\leq\rho\leq r$, then for any set $\mathcal{T}\subseteq\mathbb{T}[r]$,
${\mathop{\mathrm{max}}_{{\theta}^{*}\mathrm{:}{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{f^{l}_{\theta,v}}\right\|}^{2}_{L^{2}\left({\theta}^{*}\right)}\ }\lesssim{\mathop{\mathrm{max}}_{{\theta}^{*}\mathrm{:}{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{{\left\|f^{l}_{\theta,v}\right\|}^{2}_{L^{2}\left(C{\theta}^{*}\right)}}\ }$
${\mathop{\mathrm{max}}_{{\theta}^{*}\mathrm{:}{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{{\left\|f^{l}_{\theta,v}\right\|}^{2}_{L^{2}\left({\theta}^{*}\right)}}\ }\lesssim{\mathop{\mathrm{max}}_{{\theta}^{*}\mathrm{:}{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{f^{l}_{\theta,v}}\right\|}^{2}_{L^{2}\left(C{\theta}^{*}\right)}\ }$
for some constant $C$.
[…]
Spatial concentration For a wave packet $\left(\theta,v,l\right)$, the restriction of $Ef^{l}_{\theta,v}$ to the ball $B\left(0,r\right)$ is essentially supported on a tube. More precisely, let
$T^{l}_{\theta,v}:=N_{r^{\delta}}\left(P^{l}_{\theta,v}+rL\left(\theta\right)\right),\ \ rL\left(\theta\right):=\left\{tL\left(\theta\right):-r\leq t\leq r\right\}$
be a $r$-tube with long direction $L\left(\theta\right)$ and mini direction $M\left(\theta\right)$, here
$L\left(\theta\right):=\left(-{\xi}_{\theta},1\right),\ \ M\left(\theta\right):=\left({\xi}_{\theta},1\right)$
Figure 2: The spatial structure of a tube.
It is well-known that $Ef^{l}_{\theta,v}$ is essentially supported on $T^{l}_{\theta,v}$. Also, there is a standard $L^{2}$-estimate for $Ef$, see [9, 4] for details.
Lemma 2.4.
For a function $f:2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$, we have
${\left\|Ef\right\|}_{L^{2}\left(B_{R}\right)}\lesssim R^{1/2}{\left\|f\right\|}_{2}$
Also, if $R^{1/2+\delta}\leq r\leq R$ and $f$ is concentrated on wave packets $\left(\theta,v,l\right)$ with $T^{l}_{\theta,v}\cap B_{r}\neq\emptyset$, then
${\left\|Ef\right\|}_{L^{2}\left(B_{10r}\right)}\sim r^{1/2}{\left\|f\right\|}_{2}$
[…]
Comparing wave packet decompositions at different scales Let $r^{1/2+2\delta}\leq\rho\leq r$ and $y\in B\left(0,r\right)$. Suppose $g$ is concentrated on the $r$-scale wave packet $\left(\theta,v,l\right)$ such that $T_{\theta,v,l}\cap B\left(y,\rho\right)\neq\emptyset$. To apply induction on scale, we need decompose a function into wave packet at different scales. We break up $g$ into wave packets over the ball $B\left(y,\rho\right)$ at scale $\rho$ as follows. First, we apply a transform to recenter $B\left(y,\rho\right)$ at the origin, i.e., $\tilde{g}\left(\xi\right):=e^{i{\psi}_{y}\left(\xi\right)}g\left(\xi\right)$, here
${\psi}_{y}\left(\xi\right):=y_{1}{\xi}_{1}+\dots+y_{n-1}{\xi}_{n-1}+y_{n}\left|\xi\right|$
One can verify that $Eg\left(x\right)=E\tilde{g}\left(\tilde{x}\right)$ where $x=y+\tilde{x}$ for some $\tilde{x}\in B\left(0,\rho\right)$. Next, we break $\tilde{g}$ into scale $\rho$ wave packets
$\tilde{g}=\sum_{\left(\zeta,w,L\right)\in\mathbb{T}\left[\rho\right]}{{\tilde{g}}^{L}_{\zeta,w}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$
Let $T^{L}_{\zeta,w}\left(y\right)$denote the $\rho$-tube corresponding to $\left(\zeta,w,L\right)$ in the $x$ coordinate (contained in $B\left(y,\rho\right)$). Define ${\widetilde{\mathbb{T}}}_{\theta,v,l}$ as
${\widetilde{\mathbb{T}}}_{\theta,v,l}:=\left\{\left(\zeta,w,L\right):\ \mathrm{D}\mathrm{ist}\left(\theta,\zeta\right)\lesssim{\rho}^{-1/2},\mathrm{D}\mathrm{ist}\left(P^{L}_{\zeta,w},P^{l}_{\theta,v}+P^{0}_{\theta}-{\partial}_{\xi}{\psi}_{y}\left({\xi}_{\theta}\right)\lesssim r^{\delta}\right)\right\}$
here
$\mathrm{D}\mathrm{ist}\left(A,B\right):={\mathop{\mathrm{max}}_{a\in A}{\mathop{\mathrm{min}}_{b\in B}\mathrm{dist}\left(a,b\right)\ }\ }\ \mathrm{for}\ A,B\subseteq{\mathbb{R}}^{n}$
and $P^{0}_{\theta}$ is the tine plates centered at origin, with radius $\sim r^{1/2+\delta}$, thickness $\sim r^{\delta/2}$, and normal direction ${\xi}_{\theta}$.
Lemma 2.5 ([4], Lemma 5.4).
For any $\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{\theta,v,l}$, there holds
$\angle\left(L\left(\theta\right),L\left(\zeta\right)\right)\lesssim{\rho}^{-1/2},\angle\left(M\left(\theta\right),M\left(\zeta\right)\right)\lesssim{\rho}^{-1/2}$
$\mathrm{D}\mathrm{ist}\left(T^{L}_{\zeta,w}\left(y\right),\ \left[T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right]+P^{0}_{\theta}\right)\lesssim r^{\delta}$
here $\angle\left(v_{1},v_{2}\right)$denotes the angle between two non-zero vectors $v_{1},v_{2}\in{\mathbb{R}}^{n}$.
In [4] the relationship between the two wave packet decompositions was studied in detail. For example, if $g$ is concentrated on wave packets from $\mathcal{T}\subseteq\mathbb{T}\left[r\right]$, then which scale $\rho$ wave packets will $\tilde{g}$ be concentrated on? We state the results from [4] below.
Lemma 2.6 ([4], Lemma 5.3).
If $g$ is concentrated on wave packet from $\mathcal{T}\subseteq\mathbb{T}\left[r\right]$, then $\tilde{g}$ is concentrated on wave packets in
$\widetilde{\mathcal{T}}=\bigcup_{\left(\theta,v,l\right)\in\mathcal{T}}{{\widetilde{\mathbb{T}}}_{\theta,v,l}}$
[…]
${\mathrm{D}\mathrm{ist}}_{H}\left({\zeta}_{0},\theta\right)\lesssim{\rho}^{-1/2},\emptyset\neq T^{l}_{\theta,v}\cap B\left(y,\rho\right)\subseteq B\left(v_{0},r^{1/2+2\delta}\right)+\rho L\left({\zeta}_{0}\right)+\left\{y\right\}$
here ${\mathrm{D}\mathrm{ist}}_{H}$ denotes the Hausdorff distance. For a $r$-tube $T^{l}_{\theta,v}$, if there are multiple $\left({\zeta}_{0},v_{0}\right)$ satisfying the above conditions, then we assign $T^{l}_{\theta,v}$ to only one ${\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)$. Also, we also define
$g_{{\zeta}_{0},v_{0}}:=\sum_{\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)}{g^{l}_{\theta,v}}$
${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}:=\bigcup_{\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)}{{\widetilde{\mathbb{T}}}_{\theta,v,l}},\ \ {\tilde{g}}_{{\zeta}_{0},v_{0}}:=\sum_{\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}}{{\tilde{g}}^{L}_{\zeta,w}}$
One immediately concludes that
${\mathrm{D}\mathrm{ist}}_{H}\left({\zeta}_{0},\zeta\right)\lesssim{\rho}^{-\frac{1}{2}},P^{L}_{\zeta,w}\subseteq B\left(v_{0},r^{\frac{1}{2}+2\delta}\right),\ \ \forall\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}},\quad g=\sum_{{\zeta}_{0},v_{0}}{g_{{\zeta}_{0},v_{0}}}+\mbox{RapDec}\left(r\right){\left\|g\right\|}_{2}$ (2.2)
Since each $\left(\zeta,w,L\right)$ belongs to $\sim 1$ ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, the Lemma 2.1 implies
${\left\|g\right\|}^{2}_{2}\sim\sum_{{\zeta}_{0},v_{0}}{{\left\|g_{{\zeta}_{0},v_{0}}\right\|}^{2}_{2}},\ \ {\left\|\tilde{g}\right\|}^{2}_{2}\sim\sum_{{\zeta}_{0},v_{0}}{{\left\|{\tilde{g}}_{{\zeta}_{0},v_{0}}\right\|}^{2}_{2}}$ (2.3)
From the lemma 2.5 and 2.6, one has
Corollary 2.7 ([4], Lemma 5.6).
If $g$ is concentrated on wave packets in ${\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)$, then $\tilde{g}$ is concentrated on wave packets in ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$. On the other hand, if $\tilde{g}$ is concentrated on wave packets in ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, then inside $B\left(y,\rho\right)$$g$ is concentrated on wave packets in
$\bigcup_{\begin{array}[]{c}{\mathrm{D}\mathrm{ist}}_{H}\left({\zeta}_{0},\zeta\right)\lesssim{\rho}^{-1/2}\\ {\mathrm{D}\mathrm{ist}}_{H}\left(v_{0},v\right)\lesssim r^{1/2+2\delta}\end{array}}{{\mathbb{T}}_{\zeta,v}\left(y\right)}$
[…]
Tangential and transverse wave packets Here we recall some definitions and properties about the tangential and transverse wave packets from [9, 4].
Definition 2.8 (Transverse complete intersection, [9], Section 5.1).
Suppose $1\leq m\leq n$. Let $Z\left(P_{1},\dots,P_{n-m}\right)$ denote the set of common zeros of the polynomials $P_{1},\dots,P_{n-m}$ on ${\mathbb{R}}^{n}$. The variety $Z\left(P_{1},\dots,P_{n-m}\right)$ is said to be a transverse complete intersection if
$\nabla P_{1}\left(x\right)\wedge\dots\wedge\nabla P_{n-m}\left(x\right)\neq 0,\ \ \forall x\in Z\left(P_{1},\dots,P_{n-m}\right)$
The degree of the transverse complete intersection is defined as ${\mathop{\mathrm{max}}_{1\leq j\leq n-m}{\mathrm{deg}P_{j}\ }\ }$.
From the above definition, if $\mathrm{Z}:=Z\left(P_{1},\dots,P_{n-m}\right)$ is a transverse complete intersection, then $\mathrm{Z}$ is a smooth manifold of dimension $m$, in other words, $\mathrm{Z}$ has a tangent plane $T_{x}\mathrm{Z}$ at every point $x\in\mathrm{Z}$.
Definition 2.9 (Tangential and transverse tubes, [9], Section 5.2).
Let $\boldsymbol{\mathrm{Z}}$ be a transverse complete intersection of dimension $m$. For a fixed ball $B\left(y,r\right)$, a $r$-tube $T^{l}_{\theta,v}$ with $T^{l}_{\theta,v}\cap B\left(y,r\right)\cap\boldsymbol{\mathrm{Z}}\neq\emptyset$ is said to be tangent to $\mathrm{Z}$ in $B\left(y,r\right)$ if $T^{l}_{\theta,v}\cap B\left(y,2r\right)\subseteq N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{\mathrm{Z}}\right)$ and
$\angle\left(L\left(\theta\right),T_{x}\boldsymbol{\mathrm{Z}}\right)\lesssim r^{-1/2+{\delta}_{m}},\ \ \forall x\in\mathrm{Z}\cap B\left(y,r\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(T^{l}_{\theta,v})\right)$
Otherwise, we say that $T^{l}_{\theta,v}$ is transverse to $\boldsymbol{\mathrm{Z}}$ in $B\left(y,r\right)$. We use $\mathbb{T}\left[\boldsymbol{\mathrm{Z}},B\left(y,r\right)\right]$ to denote the $r$-tube tangent to $\boldsymbol{\mathrm{Z}}$ in $B\left(y,r\right)$.
In secton 3 we need the following result, which can be roughly taken as a continuum version of the Bezout’s theorem, to controls the transverse intersections between a tube and an algebraic variety,
Lemma 2.10 ([9], Lemma 5.7).
Let $T$ be a cylinder of radius $r$ with central line $l$ and suppose that $Z=Z\left(P_{1},\dots,P_{n-m}\right)\subseteq{\mathbb{R}}^{n}$ is a transverse complete intersection with degree at most $D$. For any $\alpha>0$, define
$Z_{>\alpha}:=\left\{x\in Z:\angle\left(l,T_{x}\boldsymbol{\mathrm{Z}}\right)>\alpha\right\}$
Then $Z_{>\alpha}\cap T$ is contained in a union of $\lesssim D^{n}$ balls of radius $\lesssim r{\alpha}^{-1}$.
2.3 Some Kakeya-type geometric results
First, we introduce some definitions about grains and multigrain from section 3 of [13].
Definition 2.11 ([13], Definition 3.3).
A grain is defined to be a pair $\left(Z,B\left(y,r\right)\right)$ where $Z$ is a transverse complete intersection and $B_{r}$ is a ball of some radius $r>0$. The dimension of a grain $\left(Z,B\left(y,r\right)\right)$ is that of the variety $Z$, while its degree is that of $Z$ and its scale is $r$.
Definition 2.12 ([13], Definition 3.4).
Let $S=\left(Z,B\left(y,r\right)\right)$ be a grain of dimension $m$. A function $f$ with $\mathrm{supp}\ f\subseteq 2{\overline{B}}^{n-1}\backslash B^{n-1}$ is said to be tangent to $S$ if it is concentrated on scale $r$ wave packets belong to the collection:
$\mathbb{T}\left[S\right]:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]:T^{l}_{\theta,v}\left(y\right)\ \mathrm{is}\ \mathrm{tangent\ to}\ Z\ \mathrm{in}\ B\left(y,r\right)\right\}$
Definition 2.13 ([13], Definition 3.5).
A multigrain is a tuple of grains
$\overrightarrow{S}=\left(S_{s},S_{s+1},\dots,S_{t}\right),\ \ S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$
satisfying
${\mathrm{dim}Z_{i}\ }=i,\quad\forall i$
$N_{r^{1/2+{\delta}_{s}}_{s}}Z_{s}\subseteq\dots\subseteq N_{r^{1/2+{\delta}_{t}}_{t}}Z_{t}$
$r_{s}\leq r_{s+1}\leq\dots\leq r_{t}$
The parameter $\left(t-s+1\right)$ is referred to as the level of multigrain. The degree of $\overrightarrow{S}$ is defined as ${\mathop{\mathrm{max}}_{1\leq i\leq l}Z_{i}\ }$. The multiscale of $\overrightarrow{S}$ is the tuple $\overrightarrow{r}=\left(r_{s},r_{s+1},\dots,r_{t}\right)$. Finally, we write
${\overrightarrow{S}}_{1}=\left(S_{s},S_{s+1},\dots,S_{t}\right)\preccurlyeq{\overrightarrow{S}}_{2}=\left(S_{s^{\prime}},S_{s^{\prime}+1},\dots,S_{t}\right),\ \ \mathrm{if\ }s^{\prime}\geq s$
The restriction conjecture is associated with a Kakeya-type problem concerning the continuum incidence theory of tubes. Specifically, given a multigrain ${\overrightarrow{S}}_{s,t}=\left(S_{s},\dots,S_{t}\right)$ with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$, if $\mathcal{L}$ is a set of $r^{-1/2}_{t}$-direction separated rays that satisfies
$L\cap B\left(y_{i},r_{i}\right)\cap N_{{\rho}_{j}}Z_{i}\neq\emptyset$
$L\cap B\left(y_{i},2r_{i}\right)\subseteq N_{{\rho}_{j}}Z_{i}\cap B\left(y_{i},2r_{i}\right)$
for any $L\in\mathcal{L}$, then what is the maximum possible value of $\#\mathcal{L}$? Here $\mathcal{L}$ is said to be $r^{-1/2}$-direction separated if
$\angle\left(L_{1},L_{2}\right)\gtrsim r^{-\frac{1}{2}},\ \ \mathrm{if}\ L_{1},L_{2}\in\mathcal{L}\boldsymbol{\ }\mathrm{and}\boldsymbol{\ }L_{1}\neq L_{2}$
[…]
Theorem 2.14 (Polynomial Wolff Axiom).
Let $r\geq 1$. Let $Z\subset{\mathbb{R}}^{n}$ be an algebraic variety of dimension $m$ and degree at most $d$. If $\mathcal{T}\subseteq\mathbb{T}\left[r\right]$satisfies: $\forall T\in\mathcal{T}$,
$T\cap B_{r}\cap N_{r^{1/2+{\delta}_{m}}}Z\neq\emptyset$
$T\cap 2B_{r}\subseteq N_{r^{1/2+{\delta}_{m}}}Z$
then
#$\#\mathcal{T}{\lesssim}_{d}r^{O\left({\delta}_{m}\right)}r^{\frac{m-1}{2}}$
[…]
Definition 2.15 (Nested Tubes).
Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$ and ${\mathrm{dim}Z_{i}\ }=i$. We define $\mathbb{T}\left[{\overrightarrow{S}}^{i}_{m}\right]$ by induction on $i\in\mathbb{N}\cap\left[m,n-1\right]$. For $i=m$, we set
$\mathbb{T}\left[{\overrightarrow{S}}^{m}_{m}\right]:=\mathbb{T}\left[S_{m}\right]$
For $m<i\leq n-1$, let $\mathbb{T}\left[{\overrightarrow{S}}^{i}_{m}\right]$ be the set of all $r_{i}$-tubes $T^{l}_{\theta,v}$ satisfying the following conditions:
(a). $T^{l}_{\theta,v}\in\mathbb{T}\left[Z_{i},B\left(y_{i},r_{i}\right)\right]$
(b). There exists $\exists T^{L}_{\zeta,w}\in\mathbb{T}\left[Z_{i-1},B\left(y_{i-1},r_{i-1}\right)\right]$ such that
$\quad{\mathrm{D}\mathrm{ist}}_{H}\left(w,\ \zeta\right)\lesssim r^{-1/2}_{i-1},\quad{\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w},\ T^{l}_{\theta,v}\cap B\left(y_{i},r_{i}\right)\right)\lesssim r^{1/2+\delta}_{i}$
Finally, we set
$\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\mathbb{T}\left[{\overrightarrow{S}}^{n-1}_{m}\right],\ \ \mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\left\{\theta:\exists T^{l}_{\theta,v}\in\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{m}\right]\ \mathrm{for\ some\ }v,l\right\}$
One can observe the second condition in definition 2.15 is weaker than lemma 2.5. However, this is good enough for our arguments. In [13, 19, 25] a non-trivial extension of Theorem 2.14, which is a corollary to lemma 2.11 in [17], was used to establish a new result on the Fourier restriction, the Bochner-Riesz problem, and the Stein’s square functions respectively. By setting $r_{0}=r_{1}$ in this extension (see [13], Lemma 3.7), one has
Lemma 2.16 (Nested polynomial Wolff axioms).
Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$,
${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d,\ \ \forall i\in\mathbb{N}\cap\left[m,n\right)$
then
#$\#\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]{\lesssim}_{d}r^{O\left({\delta}_{0}\right)}_{n-1}\left(\prod^{n-2}_{j=m}{r^{-\frac{1}{2}}_{j}}\right)r^{\frac{n-2}{2}}_{n-1}$
In [13, 19, 25] $S_{m}$ was passed back to $S_{n}:=\left({\mathbb{R}}^{n},B_{R}\right)$, and thus
$\mathrm{\Theta}\left[{\overrightarrow{S}}_{m}\right]{\lessapprox}_{d}\left(\prod^{n-1}_{j=m}{r^{-1/2}_{j}}\right)R^{\left(n-1\right)/2}$
Since $\mathrm{\Theta}\left[R\right]\sim R^{\left(n-1\right)/2}$ in the case of paraboloids, this bound gives a gain of $\prod^{n-1}_{j=m}{r^{-1/2}_{j}}$. However, in the case of cone this doesn’t guarantee a gain due to $\mathrm{\Theta}\left[R\right]\sim R^{\left(n-2\right)/2}$. This is the reason for passing $S_{m}$ back to $S_{n-1}$ rather than $S_{n}$ in our arguments.
2.4 $k$-broad Norms
[…]
$L\left(\tau\right):=\left\{L\left(\theta\right):\theta\subset\tau\right\}$
Cover $B_{R}$ by finitely overlapping $B_{K^{2}}$. Fix a parameter $1<A\lesssim K^{\epsilon}$. For each $B_{K^{2}}$, define
${\mu}_{Ef}\left(B_{K^{2}}\right):={\mathop{\mathrm{min}}_{V_{1},\dots,\ V_{A}\ \left(k-1\right)\mathrm{-}\mathrm{subspace\ of}\ {\mathbb{R}}^{n}}\left(\mathop{\mathrm{max}}_{\tau:\angle\left(L\left(\tau\right),V_{a}\right)\gtrsim K^{-2},\forall 1\leq a\leq A}\int_{B_{K^{2}}}{{\left|Ef_{\tau}\right|}^{p}}\right)\ }$
here
$\angle\left(L\left(\tau\right),V_{a}\right):={\mathop{\mathrm{min}}_{v\in G\left(\tau\right)}\angle\left(v,V_{a}\right)\ }$
Compared with the paraboloid case [9], the angel condition is modified to $\angle\left(G\left(\tau\right),V_{a}\right)\gtrsim K^{-2}$, due to some geometric feature of the cone. See [4] for details. For $U\subseteq{\mathbb{R}}^{n}$ the $k$-board norm over $U$ is defined as follows
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}:={\left(\sum_{B_{K^{2}}}{\frac{\left|B_{K^{2}}\cap U\right|}{\left|B_{K^{2}}\right|}}{\mu}_{Ef}\left(B_{K^{2}}\right)\right)}^{\frac{1}{p}}$
[…]
Basic Properties The parameter $A$ is used to make $k$-board norm behave like a regular $L^{p}$-norm. Here we recall some inequalities about .${\left\|.\right\|}_{BL^{p}_{k,A}}$.
Lemma 2.17 (Triangle inequality, [4], Lemma 3.1).
Suppose that $1\leq p<\infty$, $f=g+h$, and $A=A_{1}+A_{2}$, where $A,A_{1},A_{2}$ are non-negative integers. Then
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}\lesssim{\left\|Eg\right\|}_{BL^{p}_{k,A_{1}}\left(U\right)}+{\left\|Eh\right\|}_{BL^{p}_{k,A_{2}}\left(U\right)}$
Lemma 2.18 (Hölder’s inequality, [4], Lemma 3.2).
Suppose $1\leq p,p_{1},p_{2}<\infty$, and $0\leq{\alpha}_{1},{\alpha}_{2}\leq 1$ obey ${\alpha}_{1}+{\alpha}_{2}=1$ and
$\frac{1}{p}=\frac{{\alpha}_{1}}{p_{1}}+\frac{{\alpha}_{2}}{p_{2}}$
Suppose that and $A=A_{1}+A_{2}$, then
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}\leq{\left\|Ef\right\|}^{{\alpha}_{1}}_{BL^{p_{1}}_{k,A_{1}}\left(U\right)}{\left\|Ef\right\|}^{{\alpha}_{2}}_{BL^{p_{2}}_{k,A_{2}}\left(U\right)}$
Lemma 2.19 (Standard $L^{2}$-estimate, [4], Section 5).
For $f\in L^{2}\left(2{\overline{B}}^{n-1}\backslash B^{n-1}\right)$ and non-negative integers $A$, there holds
${\left\|Ef\right\|}^{2}_{BL^{2}_{k,A}\left(B_{r}\right)}\lesssim r{\left\|f\right\|}^{2}_{2}$
Lemma 2.20 (Vanishing property, [4], Lemma 4.5).
Suppose $Ef$ is tangent to a transverse complete intersection $\boldsymbol{Z}$ in $B_{r}$. If ${\mathrm{deg}\boldsymbol{Z}\lesssim 1\ }$ and ${dim\boldsymbol{Z}\ }$ $\leq k-1$, then
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{r}\right)}\lesssim\mathrm{RapDec}\left(r\right){\left\|f\right\|}_{2}$
These inequalities will be used frequently in section 3.
Now suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\boldsymbol{\mathrm{Z}},B\left(0,r\right)\right]$, where $\boldsymbol{\mathrm{Z}}$ is a $m$-dimensional transverse complete intersection. Fix a ball $B\left(y,\rho\right)\subseteq B\left(0,r\right)$. To apply the induction on scale, we expect to find a decomposition
$\tilde{g}=\sum_{b\in\mathcal{B}}{{\tilde{g}}_{b}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$ (2.4)
where
${\tilde{g}}_{b}:=\sum_{\left(\zeta,w,L\right)\in\mathbb{T}\left[{\boldsymbol{\mathrm{Z}}}_{b}-y,B\left(0,\rho\right)\right]}{{\tilde{g}}^{L}_{\zeta,w}},\ \ {\boldsymbol{\mathrm{Z}}}_{b}:=\left\{z+b:z\in\boldsymbol{\mathrm{Z}}\right\}$
Roughly speaking, for each $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$, $N_{{\rho}^{1/2+{\delta}_{m}}}\left({\boldsymbol{\mathrm{Z}}}_{b}\right)\cap B\left(y,\rho\right)$ is a ‘slice’ of $N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{\mathrm{Z}}\right)\cap B\left(y,\rho\right)$. Let $g_{b}\left(\xi\right):=e^{-i{\psi}_{y}\left(\xi\right)}{\tilde{g}}_{b}\left(\xi\right)$, then
$Eg_{b}\left(x\right)\approx{\chi}_{N_{{\rho}^{1/2+{\delta}_{m}}}\left({\boldsymbol{\mathrm{Z}}}_{b}\right)}\left(x\right)Eg\left(x\right)$ (2.5)
If $\boldsymbol{\mathrm{Z}}$ is a plane, then the existence of such a decomposition follows from the Lemma 2.6. However, this is not easy for ${\mathrm{deg}\boldsymbol{\mathrm{Z}}\ }>1$, since the shapes of $\boldsymbol{\mathrm{Z}}$ is irregular. A rigorous version of (2.4) and (2.5) is given by the following statement. See [9, 4, 7] for details.
Lemma 2.21.
[…]
${\left\|Eg\right\|}^{p}_{BL^{p}_{k,A}\left(B\left(y,\rho\right)\right)}\lesssim{\mathrm{log}r\ }\sum_{b\in\mathcal{B}}{{\left\|Eg^{\prime}_{b}\right\|}^{p}_{BL^{p}_{k,A}\left(B_{\rho}\cap N_{{\rho}^{1/2+{\delta}_{m}}}\left(\boldsymbol{Z}+b\right)\right)}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$
and
$\ \ \sum_{b\in\mathcal{B}}{{\left\|g^{\prime}_{b}\right\|}^{2}_{2}}\lesssim{\left\|g^{\prime}\right\|}^{2}_{2}$
where
$g^{\prime}:=\sum_{\left(\theta,v,l\right)\in{\mathcal{T}}^{\prime}}{g^{l}_{\theta,v}}$
[…]
Reduction to Broad Estimates The main result of this paper is
Theorem 2.22.
Let $22\leq k\leq n-$ 2 and
$p>p_{n}\left(k\right):=2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}}$ (2.6)
Then there exists an integer $A$ such that
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}{\lesssim}_{p,\varepsilon}R^{\varepsilon}{\left\|f\right\|}^{2/p}_{2}{\left\|f\right\|}^{1-2/p}_{\infty}$ (2.7)
Theorem 2.22, the $k$-broad estimate, implies Theorem 1.5, the linear cone restriction estimate. In particular, it follows from
Theorem 2.23 ([4], Section 6.1).
For any $R>1$, the estimate (1.5) holds whenever the estimate (2.7) holds, for $p$ in the range
[…]
We choose a value of $k$ to optimizes the range of $p$ in (1.5). By applying the arguments in [13] to (2.6), one has
$2+\frac{12}{4n-5+2\left(k-1\right){\left(\frac{k}{n-1}\right)}^{1/2}}\leq p_{n}\left(k\right)\leq 2+\frac{12}{4n-5+2\left(k-1\right){\left(\frac{2n-1}{2k+1}\right)}^{1/2}\frac{k}{n-1}}$
By writing $k=vn+O\left(1\right)$, one has
$p_{n}\left(k\right)=2+\frac{6}{2+v^{3/2}}n^{-1}+O\left(n^{-2}\right)$
From (2.8),
$p\left(k,n\right)=2+\frac{4}{2-v}n^{-1}+O\left(n^{-2}\right)$ (2.9)
To obtain the best possible estimate in the theorem 1.5, $v$ should satisfy
$\frac{6}{2+v^{3/2}}=\frac{4}{2-v}$
Hence, we choose
$v^{\frac{1}{2}}={\left(\frac{3}{8}+\frac{1}{2\sqrt{2}}\right)}^{1/3}+{\left(\frac{3}{8}-\frac{1}{2\sqrt{2}}\right)}^{1/3}-\frac{1}{2}=0.67765\dots$
Plugging this back into (2.9) yields the linear restriction estimate (1.5) holds for
$p>2+\lambda n^{-1}+O\left(n^{-2}\right)$
where
$\lambda=\frac{4}{2-v}=2.59607\dots$
This gives theorem 1.5.
2.5 Transverse Equidistribution Estimates
By using a version of Heisenberg uncertainty principle, Guth [9] established the transverse equidistribution estimates for paraboloids. However, due to the thin thickness of tubes in the case of cone, $Eg$ may be concentrated on an extremely thin layer of $N_{R^{1/2+{\delta}_{m}}}\left(\boldsymbol{\mathrm{Z}}\right)$, and thus this property may fail to hold.
Figure 3: A scenario that suggests the transverse equidistribution fails.
For example, let $\boldsymbol{\mathrm{Z}}$ be a $\left(n-1\right)$-plane, and $B\left(y,r\right)\subseteq B\left(0,R\right)$ for some $y\in\boldsymbol{\mathrm{Z}}$ and $r\sim R^{2/3}$. Let $T^{l}_{\theta,v}$ be a $R$-tube with $T^{l}_{\theta,v}\subseteq N_{R^{1/2+{\delta}_{n-1}}}\left(\boldsymbol{\mathrm{Z}}\right)\cap B\left(0,R\right)$, $L\left(\theta\right)\parallel\boldsymbol{\mathrm{Z}}$, and $M\left(\theta\right)\bot\boldsymbol{\mathrm{Z}}$. Suppose $g$ is concentrated on $T^{l}_{\theta,v}$. Since $r^{1/2+{\delta}_{n-1}}\gg R^{\delta}$, there exist $b\in B\left(0,R^{1/2+{\delta}_{m}}\right)$ such that $T^{l}_{\theta,v}\subseteq N_{r^{1/2+{\delta}_{n-1}}}\left(\boldsymbol{\mathrm{Z}}\mathrm{+b}\right)\cap B\left(0,R\right)$. Therefore,
${\left\|Eg_{b}\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}={\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}+\mathrm{RapDec}\left(R\right){\left\|g\right\|}^{2}_{2}$
Since $r/R\sim R^{-1/3}$, one has
${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}\gg R^{O\left({\delta}_{n-1}\right)}{\left(\frac{r}{R}\right)}^{1/2}{\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$
If ${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$ doesn’t vanish, then the Lemma 2.4 gives
${\left\|g_{b}\right\|}^{2}_{2}\gg R^{O\left({\delta}_{n-1}\right)}{\left(\frac{r}{R}\right)}^{1/2}{\left\|g\right\|}^{2}_{2}$
which suggests the transverse equidistribution estimates fail.
[…]
${\mathbb{T}}_{B_{\rho}}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[r\right],T^{l}_{\theta,v}\cap B_{\rho}\neq\emptyset\right\}$
For a vector space $V$ spanned by ${\left\{v_{i}=\left(x_{i,1},x_{i,2},\dots,-x_{i,n}\right)\right\}}^{m}_{i=1}$, $V^{+}$ is defined by
[…]
Now we cover $B_{\rho}$ by finitely overlapping $r^{1/2+{\delta}_{m}}$-ball $B$. For some suitable constant $c>0$, if $cB\cap B_{\rho}\cap\boldsymbol{\mathrm{Z}}\boldsymbol{\mathrm{\neq}}\emptyset$, then we choose some $x\in cB\cap B_{\rho}\cap\boldsymbol{\mathrm{Z}}$, and use $V_{B}$ to denote $T_{x}\boldsymbol{\mathrm{Z}}$. Define
${\mathcal{B}}_{a}:=\left\{B:\mathrm{\ }\angle\left(V^{+}_{B},V^{\bot}_{B}\right)\gtrsim K^{-2}\right\}$
${\mathcal{B}}_{b}:=\left\{B:\mathrm{\ }\angle\left(V^{+}_{B},V^{\bot}_{B}\right)\lesssim K^{-2}\right\}$
${\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[\boldsymbol{\mathrm{Z}},B\left(0,r\right)\right]:T^{l}_{\theta,v}\cap B\neq\emptyset\right\}$
From the definition above, if $B\in{\mathcal{B}}_{b}$, then $\angle\left(M\left(\theta\right),V^{\bot}_{B}\right)$ may be small for some $\left(\theta,v,l\right)\in{\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}$, and thus the transverse equidistribution estimates may not hold for a function concentrated on ${\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}$. Fortunately, there are only a few $\theta$ such that $\left(\theta,v,l\right)\in{\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}$ for some $v,l$.
Lemma 2.24 ([4], lemma 5.8).
Suppose $B\in{\mathcal{B}}_{b}$̇et
$\mathbb{T}\left[{\theta}^{\prime}\right]:=\left\{\left(\theta,v,l\right):\theta={\theta}^{\prime}\right\},\ \ {\mathrm{\Theta}}_{B}:=\left\{\theta:\mathbb{T}\left[\theta\right]\cap{\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}\neq\emptyset\right\}$
Then ${\mathrm{\Theta}}_{B}$ is contained in the union of $O\left(1\right)$ many $K^{-2}$-sectors.
Now, we recall the definition of the essential part $g_{\mathrm{ess}}$ and the tail part $g_{\mathrm{tail}}$.
${\mathbb{T}}_{\mathrm{ess}}:=\left\{\left({\zeta}_{0},v_{0}\right):\exists\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left.y\right.\mathrm{\ s.t.\ }T^{l}_{\theta,v}\cap B\neq\emptyset\mathrm{\ }\mathrm{fo}\mathrm{r\ some\ }B\in{\mathcal{B}}_{a}\right\}$
${\mathbb{T}}_{\mathrm{tail}}:=\left\{\left({\zeta}_{0},v_{0}\right):\forall\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left.y\right.\mathrm{,\ }T^{l}_{\theta,v}\cap B=\emptyset\mathrm{\ }\mathrm{fo}\mathrm{r\ all\ }B\in{\mathcal{B}}_{a}\right\}$
$g_{\mathrm{ess}}:=\sum_{\left({\zeta}_{0},v_{0}\right)\in{\mathbb{T}}_{\mathrm{ess}}}{g_{{\zeta}_{0},v_{0}}},\ \ g_{\mathrm{tail}}:=\sum_{\left({\zeta}_{0},v_{0}\right)\in{\mathbb{T}}_{\mathrm{tail}}}{g_{{\zeta}_{0},v_{0}}}$ (2.10)
Suppose $A$ is large enough. By the definition of $k$-board norm, and lemma 2.24,
${\left\|Eg_{\mathrm{tail}}\right\|}^{p}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}\lesssim\sum_{B\in{\mathcal{B}}_{b}}{{\left\|Eg_{\mathrm{tail}}\right\|}^{p}_{BL^{p}_{k,A/2}\left(B\right)}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{p}_{2}\lesssim\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{p}_{2}$ (2.11)
By lemma 2.17 and (2.11),
${\left\|Eg\right\|}_{BL^{p}_{k,A}\left(B_{\rho}\right)}\lesssim{\left\|Eg_{\mathrm{ess}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}+{\left\|Eg_{\mathrm{tail}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}\lesssim{\left\|Eg_{\mathrm{ess}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}+\mathrm{RapDec}\left(R\right){\left\|g\right\|}_{2}$ (2.12)
[…]
Transverse Equidistribution Estimates for Essential Part As shown in [4], the transverse equidistribution estimates hold for the essential part $g_{\mathrm{ess}}$. The precise statement of the transverse equidistribution estimates for cone is:
Lemma 2.25 ([4], lemma 5.13).
Let $B_{\rho}$ be a $\rho$-ball with $\rho\gg r^{1/2+{\delta}_{m}}$ and $B_{\rho}\cap N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{Z}\right)\cap B\left(0,r\right)\neq\emptyset$. Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\boldsymbol{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B_{\rho}}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$  then
${\left\|g_{\mathrm{ess,}b\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g\right\|}^{2}_{2}$
where $g_{\mathrm{ess,}b\mathrm{\ }}:={\left(g_{\mathrm{ess\ }}\right)}_{b}$.
To apply the nested polynomial Wolff axioms, we need to generalize lemma 2.25.
Lemma 2.26.
Let $\rho\gg r^{1/2+{\delta}_{m}}$ and $B\left(y,\rho\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{Z}\right)\cap B\left(0,r\right)\neq\emptyset$. Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\boldsymbol{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B\left(y,\rho\right)}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$. For $\widetilde{\mathcal{T}}\subseteq\mathbb{T}\left[{\boldsymbol{\mathrm{Z}}}_{b}-y,B\left(0,\rho\right)\right]$, there holds
${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$
here $\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}$ denotes the set of $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ for which there exists some $\left(\zeta,w,L\right)\in\widetilde{\mathcal{T}}$ satisfying
${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq C{\rho}^{-1/2},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y\right),\ T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right)\leq Cr^{1/2+\delta}$
for some constant $C$, and ${\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}:={\widetilde{\left(g_{\mathrm{ess}}\right)}}_{\widetilde{\mathcal{T}}}$.
Proof.
Let $h=g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}$. By lemma 2.25,
${\left\|{\tilde{h}}_{\mathrm{ess},b}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}$ (2.13)
By lemma 2.1 and 2.6,
${\left\|{\tilde{h}}_{\mathrm{ess},b}\right\|}^{2}_{2}\gtrsim{\left\|{\widetilde{\left(h_{\mathrm{ess}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}={\left\|{\widetilde{\left(h_{\mathrm{ess,}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.14)
here $\mathrm{\uparrow}\widetilde{\mathcal{T}}$ denotes the set of $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ for which there exists some $\left(\zeta,w,L\right)\in\widetilde{\mathcal{T}}$ satisfying
${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq c{\rho}^{-1/2},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y\right),\ T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right)\leq cr^{1/2+\delta}$
for some large constant $c$. Also, lemma 2.6 gives
${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}={\left\|{\widetilde{\left(g_{\mathrm{ess,}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}={\left\|{\widetilde{\left(h_{\mathrm{ess,}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.15)
Combine (2.13), (2.14), and (2.15), one has
${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{ess,}\ \mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$
2.6 Polynomial Partitioning
[…]
${\left\|F\right\|}_{L^{1}\left(O_{1}\right)}={\left\|F\right\|}_{L^{1}\left(O_{2}\right)},\ \forall O_{1},O_{2}\in\mathrm{cell}\left(P\right)$ (2.16)
See [8] for details.
After a small perturbation and using Sard’s theorem [9], $P$ is guaranteed to be non-singular by relaxing (2.16). The precise statement of the result is as follows.
Theorem 2.27 ([9], theorem 5.5).
Suppose $F\in L^{1}$ and $d$ is a positive integer. Then, there exists a non-singular polynomial $P$ of degree $O\left(d\right)$ such that $\mathrm{cell}\left(P\right)\sim d^{n}$ and
${\left\|F\right\|}_{L^{1}\left(O\right)}\sim d^{-n}{\left\|F\right\|}_{L^{1}\left(B_{r}\right)},\ \forall O\in\mathrm{cell}\left(P\right)$
Suppose we have a polynomial $P$ described in the theorem 2.27. Let ${\delta}^{\prime}>0$ be a small number. Let $\mathcal{O}$ denote the set of connected components of $B_{r}\backslash N_{r^{1/2+{\delta}^{\prime}}}Z\left(P\right)$. By pigeonholing, we see that at least one of two cases hold:
(1).The cellular term dominates, i.e.,
$\sum_{O\in\mathcal{O}}{{\left\|F\right\|}_{L^{1}\left(O\right)}}\gtrsim{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
(2).The algebraic term dominates, i.e.,
${\left\|F\right\|}_{L^{1}\left(N_{r^{1/2+{\delta}^{\prime}}}Z\left(P\right)\right)}\gtrsim{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
[…]
Partitioning over lower dimensional sets Theorem 2.27 is not sufficient to prove theorem 2.22. We will use a more involved partitioning result, which partitions a general variety in ${\mathbb{R}}^{n}$.
Theorem 2.28 ([9], Section 8).
Fix $r\gg 1$, and $d\in{\mathbb{N}}^{+}$. Suppose $F\in L^{1}\left({\mathbb{R}}^{n}\right)$ is supported on $B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}$, where $\boldsymbol{\mathrm{Z}}$ is an $m$-dimensional transverse complete intersection of degree at most $d$. Then, at least one of the following cases holds:
Cellular case. There exists a polynomial $P:{\mathbb{R}}^{n}\to\mathbb{R}$ of degree $O\left(d\right)$ with the following properties:
(1).$\mathcal{O}\sim d^{m}$, and each $O\in\mathcal{O}$ is contained in a ball of radius $r/2$.
(2).For each $O\in\mathcal{O}$, ${\left\|F\right\|}_{L^{1}\left(O\right)}\sim d^{-m}{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
where $\mathcal{O}$ is a refinement of the set of connected components of $B_{r}\backslash N_{r^{1/2+{\delta}_{m}}}Z\left(P\right)$.
Algebraic case. The exists a $\left(m-1\right)$-dimensional transverse complete intersection $\boldsymbol{\mathrm{Y}}\subseteq\boldsymbol{\mathrm{Z}}$ of degree $O\left(d\right)$ such that
${\left\|F\right\|}_{L^{1}\left(B_{r}\right)}\lesssim{\left\|F\right\|}_{L^{1}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Y}}\right)}$
Now we apply the theorem 2.28 to the $k$-board norm estimates:
(1). If the cellular case holds, then it follows that
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}\right)}\lesssim d^{m}{\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(O\right)},\ \ \forall\ O\in\mathcal{O}$
where $\mathcal{O}$ is the collection of cells produced by theorem 2.28.
(2). If the algebraic case holds, then it follows that
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}\right)}\lesssim{\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Y}}\right)}$
where $\boldsymbol{\mathrm{Y}}$ the a $\left(m-1\right)$-dimensional transverse complete intersection produced by theorem 2.28.
From the fundamental theorem of algebra, one deduces that
Lemma 2.29.
For a $r$-tube $T^{l}_{\theta,v}$,
#$\#\left\{O\in{\mathcal{O}}^{\prime}:O\cap T^{l}_{\theta,v}\neq\emptyset\right\}\lesssim d$
3 Finding Polynomial Structure
In this section, we reformulate the induction arguments in [4] as a recursive process by combining the ideas from [12, 13]. The argument contains two different algorithms. In [alg 1] we apply the iterated polynomial partition until the tangent term dominates. This algorithm passes from an $m$-dimensional to an $\left(m-1\right)$-dimensional situation. This algorithm corresponds to the induction on the scale. [alg 2] is established by repeated application of [alg 1]. The iteration ends when we arrive in the minimal dimensional case. The lemma 2.16 is used to establish the bound. This algorithm corresponds to the induction on the dimension.
Without loss of generality, we assume that the error terms, $\mathrm{RapDec}\left(R\right){\left\|.\right\|}_{2}$, is always negligible in every iteration.
3.1 The first algorithm
This algorithm is processed by iterated polynomial partition. At each step of the iteration, one of the 3 cases will happen: cellular case, transverse case and tangent case. If the tangent case happens, or the scale is tiny enough, then the iteration will end.
Input. [alg 1] will take the following things as the input:
$\bullet$ A grain $\left(\boldsymbol{\mathrm{Z}},B\left(y,r\right)\right)$ of dimension $m$.
$\bullet$ A function $f\in L^{\infty}\left(2{\overline{B}}^{n-1}\backslash B^{n-1}\right)$ which is tangent to $\left(\boldsymbol{\mathrm{Z}},B\left(y,r\right)\right)$.
$\bullet$ An addmissible large integer $A\in\mathbb{N}$.
$\bullet$ In the initial step, we define
$O_{0}:=\left\{B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}\right\},\ \ f_{O_{0}}:=f,\ \ {\hslash}_{0}:=\emptyset,\ \ {\rho}_{0}:=r$
Output. At the $j$-th interation, [alg 1] outputs:
$\bullet$ History: A word ${\hslash}_{j}$ of length $j$ in the alphabet $\left\{\mathrm{a,c}\right\}$, which records how the cell $O_{j}\in{\mathcal{O}}_{j}$ were constructed by the iterated polynomial partitioning. Let #$\#\mathrm{a}\left(j\right)$ denote the number of ‘a’ in ${\hslash}_{j}$, and #$\#\mathrm{c}\left(j\right)$ denote the number of ‘c’ in ${\hslash}_{j}$.
$\bullet$ Spatial scale ${\rho}_{j}$: The ${\rho}_{j}$ is determined by the initial scale $r$ and ${\hslash}_{j}$. Define an auxiliary parameter ${\widetilde{\delta}}_{m-1}$ by
$\left(1-{\widetilde{\delta}}_{m-1}\right)\left(\frac{1}{2}+{\delta}_{m-1}\right)=\frac{1}{2}+{\delta}_{m}$
One deduces that
$\delta_{m-1}/2\leq{\widetilde{\delta}}_{m-1}\leq 2{\delta}_{m-1}$
$\bullet$ Cells: A family of subsets ${\mathcal{O}}_{j}$ of ${\mathbb{R}}^{n}$. Each cell $O_{j}\in{\mathcal{O}}_{j}$ is contained in some ${\rho}_{j}$-ball $B_{O_{j}}=B\left(y_{O_{j}},{\rho}_{j}\right)$.
$\bullet$ A collection of functions ${\left(f_{O_{j}}\right)}_{O_{j}\in{\mathcal{O}}_{j}}$. For each cell $O_{j}$ there exists a translate ${\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{:}=\boldsymbol{\mathrm{Z}}\boldsymbol{\mathrm{+}}x_{{\mathcal{O}}_{j}}$ such that $f_{O_{j}}$ is tangent to the grain $\left({\boldsymbol{\mathrm{Z}}}_{O_{j}},B_{O_{j}}\right)$
$\bullet$ A large integer $d$ which depends only on the admissible parameters and ${\mathrm{deg}\boldsymbol{\mathrm{Z}}\ }$.
$C^{\mathrm{I}}_{j,\delta}\left(d,r\right):=d^{\#\mathrm{c}\left(j\right)\delta}{{\mathrm{log}}^{\#\mathrm{a}\left(j\right)\left(1+\delta\right)}r\ },C^{\mathrm{II}}_{j,\delta}\left(d\right):=d^{\#\mathrm{c}\left(j\right)\delta+n\#\mathrm{a}\left(j\right)\left(1+\delta\right)}$
$C^{\mathrm{III}}_{j,\delta}\left(d,r\right):=d^{j\delta}r^{\overline{C}\#\mathrm{a}\left(j\right){\delta}_{m}},C^{\mathrm{IV}}_{j,\delta}\left(d,r\right):=C^{\mathrm{III}}_{j,\delta}\left(d,r\right),A_{j}:=3^{-\#\mathrm{a}\left(j\right)}A$
Property. At the $j$-th interation,
1. The mass of ${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}$ is concentrated over the $O_{j}\in{\mathcal{O}}_{j}$:
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\leq C^{\mathrm{I}}_{j,\delta}\left(d,r\right)\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$ (3.1)
2. The function $f_{O_{j}}$ satisfies
$\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|f_{O_{j}}\right\|}^{2}_{2}}\leq C^{\mathrm{II}}_{j,\delta}\left(d\right)d^{\#\mathrm{c}\left(j\right)}{\left\|f\right\|}^{2}_{2}$ (3.2)
3. Each $f_{O_{j}}$ satisfies for all $1\leq\rho\leq{\rho}_{j}$.
${\left\|f_{O_{j}}\right\|}^{2}_{2}\leq C^{\mathrm{III}}_{j,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}d^{-\#\mathrm{c}\left(j\right)\left(m-1\right)}{\left\|f\right\|}^{2}_{2}$ (3.3)
${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq C^{\mathrm{III}}_{j,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ },\ \forall\rho\in\left[1,{\rho}_{j}\right]$ (3.4)
4. For $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j}\right]$ and $O_{j}\in{\mathcal{O}}_{j}$,
${\left\|f_{O_{j},\ \mathcal{T}}\right\|}^{2}_{2}\leq C^{\mathrm{IV}}_{j,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\mathcal{T}}\right\|}^{2}_{2}$ (3.5)
here $\widetilde{f_{O_{j},\ \mathcal{T}}}:={\left(\widetilde{f_{O_{j}}}\right)}_{\mathcal{T}}$ and ${\uparrow}^{j}\mathcal{T}$ denotes the set of wave packets $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ satisfying
${\mathrm{dist}}_{H}\left(\theta,\zeta\right)\leq c_{j}{\rho}^{-1/2}_{j},\ \ {\mathrm{dist}}_{H}\left(T^{L}_{\zeta,w}\left(y_{O_{j}}\right),T^{l}_{\theta,v}\left(y\right)\cap B_{O_{j}}\right)\leq c_{j}r^{1/2+\delta}$
for some $\left(\zeta,w,L\right)\in\mathcal{T}$. Here
$c_{j}:=\sum^{j}_{i=1}{2^{-\left(j-1\right)/2}C}$
for some absolute constant $C$.
If $j=0$, these properties are trivial. For $j\geq 1$, we will verify these properties by induction on $j$.
Recursion steps. If ${\rho}_{j}\leq r^{{\widetilde{\delta}}_{m-1}}$, then the algorithm terminates. We set $J:=j$ and $\mathcal{O}:={\mathcal{O}}_{j}$. We mark this situation as [tiny]. Otherwise, for each $O_{j}\in{\mathcal{O}}_{j}$ we apply the polynomial partitioning, theorem 2.28, with degree $d$ to ${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}$. Let ${\mathcal{O}}_{j\mathrm{,cell}}\subseteq{\mathcal{O}}_{j}$ denote the set of all cells for which the cellular case holds and ${\mathcal{O}}_{j\mathrm{,alg}}:={\mathcal{O}}_{j}\backslash{\mathcal{O}}_{j\mathrm{,cell}}$. Then one of the following cases holds:
[…]
Cellular-dominant case Suppose
$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}\leq\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$ (3.6)
Define ${\hslash}_{j+1}$ by appending the $\mathrm{c}$ to ${\hslash}_{j}$. Then, the definition of the parameters gives
$\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right),\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right)+1,\ \ A_{j+1}=A_{j}$
$C^{\mathrm{I}}_{j+1,\delta}\left(d,r\right)=d^{\delta}C^{\mathrm{I}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{II}}_{j+1,\delta}\left(d\right)=d^{\delta}C^{\mathrm{II}}_{j,\delta}\left(d\right)$
$C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right)=d^{\delta}C^{\mathrm{III}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right)=d^{\delta}C^{\mathrm{IV}}_{j,\delta}\left(d,r\right)$
Also, let ${\rho}_{j+1}:={\rho}_{j}/2$. Since the cellular case happens, for each $O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}$, there exists a polynomial $P$ with degree $O\left(d\right)$ such that
1. $\mathcal{O}\left(O_{j}\right)\ \sim\ d^{m}$
2. For all $O\in\mathcal{O}$, $O$ is included in a ${\rho}_{j+1}$-ball $B_{O}:=B\left(y_{O},{\rho}_{j+1}\right)$.
3. ${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O\right)}\sim d^{-m}{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)},\ \ \forall O\in\mathcal{O}$
where $\mathcal{O}\left(O_{j}\right)$ is a refinement of the set of connected components of $O_{j}\backslash N_{{\rho}^{1/2+{\delta}_{m}}_{j}}Z\left(P\right)$. For $O\in\mathcal{O}\left(O_{j}\right)$, define
$f_{O}:=\sum_{\begin{array}[]{c}\left(\theta,v,l\right)\in\mathbb{T}\left[{\rho}_{j}\right]\\ T^{l}_{\theta,v}\cap O\neq\emptyset\end{array}}{{\left(f_{O_{j}}\right)}^{l}_{\theta,v}}$
By lemma 2.29, each $T^{l}_{\theta,v}$ intersects $O\left(d\right)$ cells, and thus the lemma 2.1 gives
$\sum_{O\in\mathcal{O}\left(O_{j}\right)}{{\left\|f_{O}\right\|}^{2}_{2}}\lesssim d{\left\|f_{O_{j}}\right\|}^{2}_{2}$
By pigeonholing, there exists a refinement ${\mathcal{O}}^{\prime}\left(O_{j}\right)\subseteq\mathcal{O}\left(O_{j}\right)$ with #$\#{\mathcal{O}}^{\prime}\ \sim\ d^{m}$ such that
${\left\|f_{O}\right\|}^{2}_{2}\lesssim d^{-m+1}{\left\|f_{O_{j}}\right\|}^{2}_{2},\ \ \forall O\in{\mathcal{O}}^{\prime}\left(O_{j}\right)$
Since ${\rho}_{j+1}\sim{\rho}_{j}$, there exists $\mathcal{B}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ with # $\#\mathcal{B}\ \sim\ 1$ such that
$N_{{\rho}^{1/2+{\delta}_{m}}_{j}}{\boldsymbol{\mathrm{Z}}}_{O_{j}}\subseteq\bigcup_{b\in\mathcal{B}}{N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{+}b\right)}$
For $O\in{\mathcal{O}}^{\prime}\left(O_{j}\right)$ and $b\in\mathcal{B}$, define
$O_{b}:=O\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{+}b\right),\ \ \widetilde{f_{O_{b}}}:=\sum_{\begin{array}[]{c}\left(\zeta,w,L\right)\in\mathbb{T}\left[{\rho}_{j+1}\right]\\ T^{L}_{\zeta,w}\left(y_{O}\right)\cap O_{b}\neq\emptyset\end{array}}{{\left(\widetilde{f_{O}}\right)}^{L}_{\zeta,w}}$
here $\widetilde{f_{O}}\left(\xi\right):=e^{i{\psi}_{y_{O}}\left(\xi\right)}f_{O}\left(\xi\right)$. By lemma 2.6, $f_{O_{b}}$ is tangent to the grain $\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{+}b,B_{O}\right)$. Let
${\mathcal{O}}_{j+1}:=\bigcup_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{\bigcup_{O\in{\mathcal{O}}^{\prime}\left(O_{j}\right)}{\left\{O_{b}:\ b\in\mathcal{B}\right\}}}$
Now we verify the properties. For property 1, 2, and 3, see [12] for details. Here we verify property 4. By lemma 2.1 and 2.6, for $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$,
${\left\|{\tilde{f}}_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$ (3.7)
where $\uparrow\mathcal{T}$ denotes set the $\left(\theta,v,l\right)\in\mathbb{T}\left[{\rho}_{j}\right]$ such that
${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq C{\rho}^{-\frac{1}{2}}_{j+1},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y_{O_{j+1}}\right),\ T^{l}_{\theta,v}\cap B\left(y_{O_{j+1}},{\rho}_{j+1}\right)\right)\leq C{\rho}^{1/2+\delta}_{j}$
for some $\left(\zeta,w,L\right)\in\mathcal{T}$. Here $C$ is the absolute constant that appears in the property 4. Combine (3.5) and (3.7),
${\left\|f_{O_{j+1},\ \mathcal{T}}\right\|}^{2}_{2}\lesssim d^{-\delta}C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
Since ${\rho}_{j+1}\ \sim\ {\rho}_{j}$, by choosing $d$ large enough, one has
${\left\|f_{O_{j+1},\ \mathcal{T}}\right\|}^{2}_{2}\leq C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
[…]
$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}\leq\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$
Define ${\hslash}_{j+1}$ by appending the $\mathrm{a}$ to ${\hslash}_{j}$. Then, the definition of the parameters gives
#$\#\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right)+1,\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right),\ \ A_{j+1}=A_{j}/3$
$C^{\mathrm{I}}_{j+1,\delta}\left(d,r\right)={{\mathrm{log}}^{1+\delta}r\ }C^{\mathrm{I}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{II}}_{j+1,\delta}\left(d\right)=d^{n\left(1+\delta\right)}C^{\mathrm{II}}_{j,\delta}\left(d\right)$
$C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right)=r^{\overline{C}{\delta}_{m}}d^{\delta}C^{\mathrm{III}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right)=r^{\overline{C}{\delta}_{m}}d^{\delta}C^{\mathrm{IV}}_{j,\delta}\left(d,r\right)$
Let ${\rho}_{j+1}:={\rho}^{1-{\widetilde{\delta}}_{m-1}}_{j}$. Since the algebraic case happens, for each $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$, there exists a $\left(m-1\right)$-dimensional transverse complete intersection $\boldsymbol{\mathrm{Y}}\subseteq{\boldsymbol{\mathrm{Z}}}_{O_{j}}$ of degree $C_{\mathrm{alg}}d$ such that
${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}\lesssim{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\boldsymbol{\mathrm{Y}}\right)}$
where $C_{\mathrm{alg}}$ is large constant. Cover $O_{j}$ by finitely overlapping ${\rho}_{j+1}$-balls $B$. Let ${\mathcal{B}}_{O_{j}}$ denotes the set of such balls. For each $B\in{\mathcal{B}}_{O_{j}}$, let
${\mathbb{T}}_{B}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[{\rho}_{j}\right]:T^{l}_{\theta,v}\cap B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\boldsymbol{\mathrm{Y}}\neq\emptyset\right\},f_{B}:={\left(f_{O_{j}}\right)}_{{\mathbb{T}}_{B}}$
Let ${\mathbb{T}}_{B\mathrm{,tang}}\subseteq{\mathbb{T}}_{B}$ be the set of ${\rho}_{j}$-scale wave packets $\left(\theta,v,l\right)$ such that
1. $T^{l}_{\theta,v}\cap 2B\subseteq N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\boldsymbol{\mathrm{Y}}\boldsymbol{=}N_{{\rho}^{1/2+{\delta}_{m-1}}_{j+1}}\boldsymbol{\mathrm{Y}}$
2. If $x\in T^{l}_{\theta,v}$ and $y\in\boldsymbol{\mathrm{Y}}\mathrm{\cap}2B$ satisfy $\left|y-x\right|\lesssim{\rho}^{1/2+{\delta}_{m}}_{j}=\ {\rho}^{1/2+{\delta}_{m-1}}_{j+1}$, then
$\angle\left(L\left(\theta\right),T_{y}\boldsymbol{\mathrm{Y}}\right)\lesssim{\rho}^{-1/2+{\delta}_{m-1}}_{j-1}$
Let ${\mathbb{T}}_{B\mathrm{,trans}}:={\mathbb{T}}_{B}\backslash{\mathbb{T}}_{B\mathrm{,tang}}$ and
$f_{B,\mathrm{tang}}:={\left(f_{O_{j}}\right)}_{{\mathbb{T}}_{B\mathrm{,tang}}},\ \ f_{B,\mathrm{trans}}:={\left(f_{O_{j}}\right)}_{{\mathbb{T}}_{B\mathrm{,trans}}}$
By lemma 2.18,
${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}\ }\left(O_{j}\right)}\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(B\right)}}\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(B\right)}}$
$\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{tang}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}+\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{trans}}\right\|}^{p}_{BL^{p}_{k,2A_{j+1}}\left(B\right)}}$ (3.8)
By pigeonholing, at least one of the following situations holds.
Tangential subcase: Let $C_{\mathrm{tang}}$ be a large constant. If
$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\leq\left(C_{\mathrm{tang}}/2\right)\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{tang}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}}$ (3.9)
then the algorithm terminates, and we set $J=j+1$. We mark this situation as [tang]. Let $\mathcal{S}$ be a collection of grains, which is defined by
$\mathcal{S}:=\left\{\left(\boldsymbol{\mathrm{Y}},B\right):B\in{\mathcal{B}}_{O_{j}}\mathrm{\ for\ some\ }O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}\right\}$
For $S=\left(\boldsymbol{\mathrm{Y}},B\right)\in\mathcal{S}$, define $f_{S}:=f_{B,\mathrm{tang}}$ and $B_{S}:=B$. Then lemma 2.6 implies that $f_{S}$ is tangent to $S$. Since (3.6) fails, (3.9) gives
$\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\leq C_{\mathrm{tang}}\sum_{S\in\mathcal{S}}{{\left\|Ef_{S}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B_{S}\right)}}$ (3.10)
By lemma 2.1 and corollary 2.3, for all $1\leq\rho\leq{\rho}_{j+1}$, there holds
${\mathop{\mathrm{max}}_{S\in\mathcal{S}}{\left\|f_{S}\right\|}^{2}_{2}\ }\leq C_{\mathrm{tang}}\mathop{\mathrm{max}}_{O_{j}\in{\mathcal{O}}_{j}}{\left\|f_{O_{j}}\right\|}^{2}_{2}$ (3.11)
${\mathop{\mathrm{max}}_{\begin{array}[]{c}S\in\mathcal{S}\\ \theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}\end{array}}{\left\|f_{S}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq C_{\mathrm{tang}}{\mathop{\mathrm{max}}_{\begin{array}[]{c}O_{j}\in{\mathcal{O}}_{j}\\ \theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}\end{array}}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (3.12)
Since #$\#{\mathcal{B}}_{O_{j}}\lesssim r^{n{\widetilde{\delta}}_{m-1}}$ for all $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$, the lemma 2.1 gives
$\sum_{S\in\mathcal{S}}{{\left\|f_{S}\right\|}^{2}_{2}}\leq C_{\mathrm{tang}}r^{n{\widetilde{\delta}}_{m-1}}\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|f_{O_{j}}\right\|}^{2}_{2}}$ (3.13)
Also, for $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$ and $S=\left(Y,B\right)\in\mathcal{S}$ with $B\in{\mathcal{B}}_{O_{j}}$, the lemma 2.1 and corollary 2.7 gives
${\left\|f_{S,\mathcal{T}}\right\|}^{2}_{2}\leq C_{\mathrm{tang}}{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$ (3.14)
where $\widetilde{f_{S,\mathcal{T}}}:={\left(\widetilde{f_{S}}\right)}_{\mathcal{T}}$
Transverse subcase: If (3.9) fails, then (3.8) gives
$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\lesssim\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{t}\mathrm{rans}}\right\|}^{p}_{BL^{p}_{k,2A_{j+1}}\left(B\right)}}}$
Let $f_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}:={\left(f_{B,\mathrm{t}\mathrm{rans}}\right)}_{\mathrm{ess}}$. From (2.12),
$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\lesssim\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}}$
Fix $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$ and $B\in{\mathcal{B}}_{O_{j}}$. By lemma 2.21, there exist ${\mathfrak{b}}_{B,O_{j}}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ and $g_{b}$ tangent to $\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b,B\right)$ for each $b\in{\mathfrak{b}}_{B,O_{j}}$ such that
${\left\|Ef_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}\lesssim{\mathrm{log}r\ }\sum_{b\in{\mathfrak{b}}_{B,O_{j}}}{{\left\|Eg_{b}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b\right)\right)}}$
$\sum_{b\in{\mathfrak{b}}_{B,O_{j}}}{{\left\|g_{b}\right\|}^{2}_{2}}\lesssim{\left\|f_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{2}_{2}$
Let
$\mathcal{O}\left(O_{j}\right):=\left\{B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b\right):b\in{\mathfrak{b}}_{B,O_{j}},\ B\in{\mathcal{B}}_{O_{j}}\right\}$
For $O=B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b\right)\in\mathcal{O}\left(O_{j}\right)$, define $f_{O}:=g_{b}$. Finally, let
${\mathcal{O}}_{j+1}:=\bigcup_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\mathcal{O}\left(O_{j}\right)}$
[…]
${\left\|f_{O_{j+1}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j}}\right\|}^{2}_{2}$ (3.15)
Combine (3.15) and (3.3),
${\left\|f_{O_{j+1}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)-\overline{C}{\delta}_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}d^{-\#\mathrm{c}\left(j+1\right)\left(m-1\right)}{\left\|f\right\|}^{2}_{2}$
The induction closes if $\overline{C}$ and $d$ is large enough. For a ${\rho}^{-1/2}$-sector $\theta$ with $1\leq\rho\leq{\rho}_{j+1}$, lemma 2.6, lemma 2.26, and corollary 2.2 gives
${\left\|f_{O_{j+1}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}\sum_{\begin{array}[]{c}{\theta}^{\prime}:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}\\ {\theta}^{\prime}\cap C\theta\neq\emptyset\end{array}}{{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left({\theta}^{\prime}\right)}}$
By pigeonholing, there exists a ${\rho}^{-1/2}$-sector ${\theta}^{\prime}$ such that ${\theta}^{\prime}\cap C\theta\neq\emptyset$ and
${\left\|f_{O_{j+1}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left({\theta}^{\prime}\right)}$ (3.16)
Combine (3.16) and (3.4),
${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq r^{O\left({\delta}_{m}\right)-\overline{C}{\delta}_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$
[…]
${\left\|{\tilde{f}}_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j}/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$ (3.17)
Combine (3.5) and (3.17)
${\left\|f_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)-\overline{C}{\delta}_{m}}d^{-\delta}{\left({\rho}_{j}/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
Since $\overline{C}$ and $d$ is large enough, one has
${\left\|f_{O_{j+1},\ \mathcal{T}}\right\|}^{2}_{2}\leq C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
Like the cellular case, ${\uparrow}^{j}\left(\uparrow\mathcal{T}\right)\subseteq{\uparrow}^{j+1}\mathcal{T}$, and thus the induction closes.
Remark 3.1.
A trivial observation is
#$\#\mathrm{c}\left(J\right)\lesssim{\mathrm{log}r\ },\ \ \#\mathrm{a}\left(J\right)\lesssim{\delta}^{-1}_{m-1}{\mathrm{log}{\delta}^{-1}_{m-1}\ }$ (3.18)
Hence, the accumulation of error is still negligible after the algorithm terminates. We omit the details.
Let $D:=d^{c\left(J\right)}$. If [tang] happens, then (3.1), (3.2), (3.3), (3.4), (3.5), (3.10), (3.11), (3.12), (3.13), and (3.14) gives
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox D^{\delta}\sum_{S\in\mathcal{S}}{{\left\|Ef_{S}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(B_{S}\right)}}$ (3.19)
$\sum_{S\in\mathcal{S}}{{\left\|f_{S}\right\|}^{2}_{2}}\lessapprox D^{1+\delta}{\left\|f\right\|}^{2}_{2}$ (3.20)
${\mathop{\mathrm{max}}_{S\in\mathcal{S}}{\left\|f_{S}\right\|}^{2}_{2}\ }\lessapprox D^{-\left(m-1\right)+\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.21)
${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{S}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox D^{\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (3.22)
${\left\|f_{S,\ \mathcal{T}}\right\|}^{2}_{2}\lessapprox D^{\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{J}\mathcal{T}}\right\|}^{2}_{2}$ (3.23)
If [tiny] happens, then (3.1), (3.2), (3.3), (3.4), and (3.5) gives
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox D^{\delta}\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(O\right)}}$ (3.24)
$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox D^{1+\delta}{\left\|f\right\|}^{2}_{2}$ (3.25)
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox D^{-\left(m-1\right)+\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.26)
3.2 The second algorithm
[alg 2] is established by repeated application of [alg 1] and terminates if [tiny] term dominates. Consider a family of Lebesgue exponents $p_{i}$ for $k\leq l\leq n$, which will be determined in Section 4, satisfying
$p_{k}\geq p_{k+1}\geq\dots\geq p_{n}=:p>2$
and define $0\leq{\alpha}_{i},{\beta}_{i}\leq 1$ in term of the $p_{i}$ by ${\alpha}_{n}={\beta}_{n}:=1$ and
${\alpha}_{i}:=\left(\frac{1}{2}-\frac{1}{p_{i+1}}\right){\left(\frac{1}{2}-\frac{1}{p_{i}}\right)}^{-1},\ \ {\beta}_{i}:=\left(\frac{1}{2}-\frac{1}{p}\right){\left(\frac{1}{2}-\frac{1}{p_{i}}\right)}^{-1}\ \mathrm{for}\ k\leq i\leq n-1$ (3.27)
Input: A function $f:2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$ with
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\gtrsim R^{\varepsilon}{\left\|f\right\|}_{2}$ (3.28)
Here $A\in\mathbb{N}$ is a large constant.
Output: The algorithm outputs the following objects.
$\bullet$ A collection $\mathcal{O}$ of cells such that each $O\in\mathcal{O}$ is included in a $R^{O\left({\delta}_{0}\right)}$-ball.
$\bullet$ For each $O\in\mathcal{O}$ a function $f_{O}$.
$\bullet$ Scales $\overrightarrow{r}=\left(r_{n},\dots,r_{m-1}\right)$ satisfying $R=r_{n}>r_{n-1}>\dots>r_{m-1}=1$.
$\bullet$ Integers $\overrightarrow{A}=\left(A_{n},\dots,A_{m-1}\right)$ satisfying $A=A_{n}>A_{n-1}>\dots>A_{m-1}$.
$\bullet$ A tuple of parameters $\overrightarrow{D}=\left(D_{n},\dots,D_{m-1}\right)$.
$\bullet$ For $m\leq l\leq n$ a family ${\overrightarrow{\mathcal{S}}}_{l}$ of level $\left(n-l+1\right)$ multigrain. Each ${\overrightarrow{S}}_{l}=\left(\left(Z_{l},B_{r_{l}}\right),\dots,\left(Z_{n},B_{r_{n}}\right)\right)\in{\overrightarrow{\mathcal{S}}}_{l}$ has multiscale ${\overrightarrow{r}}_{l}=\left(r_{l},\dots,r_{n}\right)$ and bounded degree. Also, for $l\leq s\leq t\leq n$:
$Z_{s}=Z^{\prime}_{s}+b\mathrm{,\ \ for\ some\ }b\mathrm{\ with\ }\left|b\right|\lesssim r^{1/2+{\delta}_{t}}_{t}\ \ \mathrm{and\ }Z^{\prime}_{s}\subseteq Z_{t}$
$\bullet$ For $m\leq l\leq n$ an assignment of a function $f_{{\overrightarrow{S}}_{l}}$ to each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}$. Each $f_{{\overrightarrow{S}}_{l}}$ is tangent to $S_{l}=\left(Z_{l},B_{r_{l}}\right)$, the first component of ${\overrightarrow{S}}_{l}$.
Property:
1.The inequality
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p_{m}}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}}\right)}^{{\beta}_{m}/p_{m}\ }$ (3.29)
holds for
$M\left(\overrightarrow{r},\overrightarrow{D}\right):={D^{\delta}_{m-1}\left(\prod^{n-1}_{i=m}{D_{i}}\right)}^{\left(n-m\right)\delta}\left(\prod^{n-1}_{i=m}{r^{\left({\beta}_{i+1}-{\beta}_{i}\right)/2}_{i}D^{\left({\beta}_{i+1}-{\beta}_{m}\right)/2}_{i}}\right)$ (3.30)
2.There holds
$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox\left(\prod^{n-1}_{i=m-1}{D^{1+\delta}_{i}}\right){\left\|f\right\|}^{2}_{2}$ (3.31)
3.For $m\leq l\leq n$,
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox r^{-\frac{n-l}{2}}_{l}\left(\prod^{l-1}_{i=m-1}{r^{-1/2}_{i}D^{-i+\delta}_{i}}\right){\mathop{\mathrm{max}}_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\ }$ (3.32)
[…]
${\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-1/2}_{j}D^{\delta}_{j-1}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}$ (3.33)
where $f^{\#}_{{\overrightarrow{S}}_{l}}:={\left(f_{{\overrightarrow{S}}_{n-1}}\right)}_{\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{l}\right]}$.
Initialization. We set
$r_{n}:=R,D_{n}:=1,A_{n}:=A$
${\overrightarrow{\mathcal{S}}}_{n}:=\left\{{\overrightarrow{S}}_{n}\right\},{\overrightarrow{S}}_{n}:=\left(S_{n}\right),S_{n}=\left({\mathbb{R}}^{n},\ B_{R}\right),f_{{\overrightarrow{S}}_{n}}:=f$
$\left(n+2-l\right)$-th step. Suppose the algorithm has run through $\left(n+1-l\right)$ steps and the following things have been determined.
${\overrightarrow{r}}_{l}=\left(r_{n},\dots,r_{l}\right),\ \ {\overrightarrow{A}}_{l}=\left(A_{n},\dots,A_{l}\right),\ \ {\overrightarrow{D}}_{l}=\left(D_{n},\dots,D_{l}\right),\ \ {\overrightarrow{\mathcal{S}}}_{l}$
For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}$ we apply [alg 1] to the $k$-board norm ${\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}$. Let #$\#\mathrm{a}\left[{\overrightarrow{S}}_{l}\right]$ and #$\#\mathrm{c}\left[{\overrightarrow{S}}_{l}\right]$ denote #$\#\mathrm{a}\left(J\right)$ and #$\#\mathrm{c}\left(J\right)$ output by [alg 1] respectively. Let $A\left[{\overrightarrow{S}}_{l}\right]$ and $r\left[{\overrightarrow{S}}_{l}\right]$ denote $A_{J}$ and $r_{J}$ output by [alg 1] respectively. Let
[…]
$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\leq 2\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.34)
By (3.18) and (3.34), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$ and $A_{l-1}\in\mathbb{N}$ such that $A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1}$ for all ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}$ and
$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\lessapprox\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.35)
For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}$, let $\mathcal{O}\left[{\overrightarrow{S}}_{l}\right]$ denote the final cells output by [alg 1]. For each $O\in\mathcal{O}\left[{\overrightarrow{S}}_{l}\right]$ let $f_{O}$ denote the function corresponding to $O$ output by [alg 1]. Finally, [alg 2] outputs the following additional things, and then terminates.
$m=l,\ \ r_{m-1}:=1,\ \ \mathcal{O}:=\bigcup_{{\overrightarrow{S}}_{m}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{m\mathrm{,tiny}}}{\mathcal{O}\left[{\overrightarrow{S}}_{m}\right]}$
[…]
Tangent-dominated case If (3.34) fails, then
$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\leq 2\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tang}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.36)
By (3.18) and (3.36), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$, $A_{l-1}\in\mathbb{N}$, and $r_{l-1}\in\left(0,R^{O\left({\delta}_{0}\right)}\right]$ such that
$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\lessapprox\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.37)
and
$A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1},\ \ r\left[{\overrightarrow{S}}_{l}\right]=r_{l-1},\ \ \forall{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}$
For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}$, let $S\left[{\overrightarrow{S}}_{l}\right]$ denote the grains output by [alg 1]. For each $S_{l\mathrm{-}\mathrm{1}}\in S\left[{\overrightarrow{S}}_{l}\right]$, let ${\overrightarrow{S}}_{l-1}:=\left(S_{l\mathrm{-}\mathrm{1}},{\overrightarrow{S}}_{l}\right)$ and $f_{{\overrightarrow{S}}_{l-1}}$ denote the function corresponding to $S_{l\mathrm{-}\mathrm{1}}$ output by [alg 1]. Finally, we define
${\overrightarrow{\mathcal{S}}}_{l\mathrm{-}\mathrm{1}}:=\bigcup_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}}{\left\{{\overrightarrow{S}}_{l-1}:S_{l\mathrm{-}\mathrm{1}}\in S\left[{\overrightarrow{S}}_{l}\right]\right\}}$
Remark 3.2.
By lemma 2.20, one has $m\geq k$. If $m<k$, then lemma 2.20 implies
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lesssim\mathrm{RapDec}\left(R\right){\left\|f\right\|}_{2}$
which contradicts (3.28).
[…]
Verify the properties Property 2 follows from (3.25) and the repeated application of (3.20). Property 3 follows from (LABEL:r3.26_) and the repeated application of (3.21). Property 1 follows from (3.19), (3.37), (3.24), (3.35), and lemma 2.18. See [12] for details. Here we verify property 4. Apply (3.23) repeatedly, one has
${\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox\prod^{n-2}_{i=l}{D^{\delta}_{i}{\left(\frac{r_{i+1}}{r_{i}}\right)}^{-\frac{n-i-1}{2}}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}=r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-1/2}_{j}D^{\delta}_{j-1}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}$
This gives property 4.
3.3 Establish the $k$-broad estimates
The task now is how to apply the properties guaranteed by the algorithms [alg 2]. Since each $O\in\mathcal{O}$ has diameter at most $R^{O\left({\delta}_{0}\right)}$, trivially one may bound
${\left\|Ef_{O}\right\|}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}\lessapprox{\left\|f_{O}\right\|}_{2}$ (3.38)
Combine (3.29) and (3.38),
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p_{m}}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}}\right)}^{{\beta}_{m}/p_{m}\ }$
$\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\right)}^{{\beta}_{m}/p_{m}\ }{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2{\beta}_{m}\left(\frac{1}{2}-\frac{1}{p_{m}}\right)}_{2}\ }$ (3.39)
Combine (3.30), (3.27), and (3.29),
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{i=m-1}{r^{\frac{{\beta}_{i+1}-{\beta}_{i}}{2}}_{i}D^{\frac{{\beta}_{i+1}}{2}-\left(\frac{1}{2}-\frac{1}{p}\right)+O\left(\delta\right)}_{i}}{\left\|f\right\|}^{\frac{2}{p}}_{2}{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{1-\frac{2}{p}}_{2}\ }$ (3.40)
By taking $l=n$ in (3.32),
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{i=m-1}{r^{-1/2}_{i}D^{-i+\delta}_{i}}\right){\left\|f\right\|}^{2}_{2}$ (3.41)
Combine (3.40) and (3.41),
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{i=m-1}{r^{\frac{{\beta}_{i+1}-{\beta}_{i}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{i}D^{\frac{{\beta}_{i+1}}{2}-\left(i+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)+O\left(\delta\right)}_{i}}{\left\|f\right\|}^{2}_{2}$ (3.42)
By choosing suitable parameters $p_{i}$ for $k\leq l\leq n$, one can guarantee
$\frac{{\beta}_{i+1}-{\beta}_{i}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)\leq 0,\ \ \ \ \ \ \ \ \frac{{\beta}_{i+1}}{2}-\left(i+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)<0,\ \ \forall m\leq i\leq n-1$
This gives the main result of Ou-Wang [4]. See [12] for details.
4 Proof of Theorem 2.22
In this section we adapt the method provided in [13] to prove theorem 2.22. Let $f:2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$ and $A$ is a large integer. Without loss of generality, we assume that $f$ satisfies (3.27). By applying [alg 2] to decompose the $k$-broad norm ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}$, we get the outputs described as Section 3.2.
First, we improve (3.42) by using lemma 2.16, the nested polynomial Wolff axioms.
Lemma 4.1.
For $m\leq l<n-1$,
${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox D^{\delta}_{n-1}\left(\prod^{n-2}_{i=l}{r^{-\frac{1}{2}}_{i}}\right){\left\|f\right\|}^{2}_{\infty}$
Proof.
Since $f^{\#}_{{\overrightarrow{S}}_{l}}$ is concentrated on the $r_{n-1}$-scale wave packets $\left(\theta,v,l\right)$ with $\theta\in\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right]$, lemma 2.16 implies
${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lesssim\#\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right]r^{-\frac{n-2}{2}}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox\left(\prod^{n-2}_{j=m}{r^{-\frac{1}{2}}_{j}}\right){\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (4.1)
By corollary 2.3,
${\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lesssim{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f_{{\overrightarrow{S}}_{n-1}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$
Put (3.22) into this bound,
${\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox D^{\delta}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq D^{\delta}_{n-1}{\left\|f\right\|}^{2}_{\infty}$ (4.2)
Then lemma 4.1 follows from (4.1) and (4.2). ∎
By combining (3.32), (3.33), and lemma 4.1, for $m\leq l<n-1$,
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox r^{-\frac{n-l}{2}}_{l}\left(\prod^{l-1}_{i=m-1}{r^{-\frac{1}{2}}_{i}D^{-i+\delta}_{i}}\right)r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-\frac{1}{2}}_{j}D^{\delta}_{j-1}}\left(\prod^{n-2}_{i=l}{r^{-\frac{1}{2}}_{i}}\right)D^{\delta}_{n-1}{\left\|f\right\|}^{2}_{\infty}$
$=\left(\prod^{n-1}_{j=m\ }{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{l-1}_{j=m-1}{D^{-j}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$ (4.3)
Let ${\gamma}_{m},{\gamma}_{m+1},\dots,{\gamma}_{n}$ be positive real numbers satisfying ${\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n}=1$, and
${\sigma}_{j}:=\sum^{j}_{i=m}{{\gamma}_{i}},\ \ {\tau}_{j}:=\left\{\begin{array}[]{c}{\sigma}_{j}\mathrm{,\ \ if\ }j\leq n-2\\ 0,\ \ \mathrm{if\ }j=n-1\end{array}\right.$
By taking the geometric mean of (3.41) and (4.3) with weights $\gamma_{i}$, we have
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{j=m}{r^{-\frac{1+{\tau}_{j}}{2}}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{-j\left(1-{\sigma}_{j}\right)+\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$
Put this bound into (3.40),
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{j=m}{r^{\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\tau}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}\prod^{n-1}_{j=m-1}{D^{\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)+O\left(\delta\right)}_{j}}{\left\|f\right\|}^{\frac{2}{p}}_{2}{\left\|f\right\|}^{1-\frac{2}{p}}_{\infty}$
It just remains to find suitable parameters $p_{i}$ for $m\leq i\leq n$ such that
$X_{j}:=\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\tau}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)\leq 0,\ \ \mathrm{for\ }m\leq j\leq n-1$
$Y_{j}:=\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)<0,\ \ \mathrm{for\ }m-1\leq j\leq n-1$
To find the critical values, we assume that $X_{j}=0$ for $m\leq j\leq n-1$ and $Y_{j}=0$ for $m-1\leq j\leq n-1$. By (3.27),
$\frac{1}{2}{\left(\frac{1}{2}-\frac{1}{p_{j+1}}\right)}^{-1}-\frac{1}{2}{\left(\frac{1}{2}-\frac{1}{p_{j}}\right)}^{-1}=\frac{1+{\tau}_{j}}{2},\ \ \mathrm{for\ }m\leq j\leq n-1$ (4.4)
$\frac{1}{2}{\left(\frac{1}{2}-\frac{1}{p_{j+1}}\right)}^{-1}=1+j\left(1-{\sigma}_{j}\right),\ \ \mathrm{for\ }m-1\leq j\leq n-1$ (4.5)
Putting (4.5) into (4.4),
$j\left(1-{\sigma}_{j}\right)-\left(j-1\right)\left(1-{\sigma}_{j-1}\right)=\frac{1+{\tau}_{j}}{2},\ \ \mathrm{for\ }m\leq j\leq n-1$ (4.6)
Thus, ${\gamma}_{m},{\gamma}_{m+1}\dots,{\gamma}_{n-1}$ can be determined recursively. Since the initial step and the recursive steps for $j\leq n-2$ are same as that in [13], we have
${\gamma}_{j}=\frac{m-1}{2}\frac{1}{j\left(j-1\right)}\prod^{j}_{i=m}{\frac{2i}{2i+1}},\ \ \mathrm{for\ }m\leq j\leq n-2$ (4.7)
Since the relationship (4.5) and ${\gamma}_{m},{\gamma}_{m+1}\dots,{\gamma}_{n-2}$ are same as that in [13], one has
$p_{j}=2+\frac{6}{2\left(j-1\right)+\left(m-1\right)\prod^{j-1}_{i=m}{\frac{2i}{2i+1}}}\ ,\ \ \mathrm{for\ }m\leq j\leq n-1$ (4.8)
The condition $X_{n-1}=0$ is equivalent to
${\left(\frac{1}{2}-\frac{1}{p_{n}}\right)}^{-1}-{\left(\frac{1}{2}-\frac{1}{p_{n-1}}\right)}^{-1}=1$ (4.9)
By putting (4.8) into (4.9), one has
$p_{n}=2+\frac{12}{4n-5+2\left(m-1\right)\prod^{n-2}_{i=m}{\frac{2i}{2i+1}}}$
By remark 3.2, $m\geq k$. Thus, if we take
$p>2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}}$
Then the broad estimate always (2.7) holds. It just remains to check ${\sigma}_{n-1}={\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n-1}\leq 1$. By (4.5),
${\sigma}_{n-2}={\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n-2}<\frac{m-1}{2}\sum_{j\geq m}{\frac{1}{j\left(j-1\right)}}=\frac{1}{2}$ (4.10)
Take $j=n-1$ in (4.6),
$\left(n-1\right)\left(1-{\sigma}_{n-1}\right)-\left(n-2\right)\left(1-{\sigma}_{n-2}\right)=1/2\ \Longrightarrow{\gamma}_{n-1}=\frac{1/2-{\sigma}_{n-2}}{n-1}$
By (4.10),
${\gamma}_{n-1}<\frac{1}{2\left(n-1\right)}$ (4.11)
Combine (4.10) and (4.11), we see that ${\sigma}_{n-1}<1$.

### v5

@@
 Then, for all $p,q>0$ satisfy
-$p>\frac{2(n-1)}{n-2},\frac{n-2}{q^{\prime}}\geq\frac{n}{p},\frac{1}{p}+\frac{1}{p^{\prime}}=\frac{1}{q}+\frac{1}{q^{\prime}}=1$ (1.2)
+$p>\frac{2(n-1)}{n-2},\frac{n-2}{q^{\prime}}\geq\frac{n}{p}$ (1.2)
 there holds
 $\|\hat{f}\|_{L^{q^{\prime}}(\mathcal{C};d\sigma)}\lesssim_{p,q}\|f\|_{p^{\prime}},\forall f:\mathbb{R}^{n}\to\mathbb{C}$
-where $d\sigma$ denotes the Riemannian measure on $\mathcal{C}$.
+where $d\sigma$ denotes the Riemannian measure on $\mathcal{C}$, and $p^{\prime},q^{\prime}$ are the Hölder conjugates of $p,q$.
 Let $Ef$ be the associated Fourier extension operator on $\mathcal{C}$, i.e.,
-[…]
-By a short duality argument, we see that conjecture 1.1 is equivalent to
+$Ef(x):=\int_{2\bar{B}^{n-1}\setminus B^{n-1}}e^{i(x_{1}\xi_{1}+...+x_{n-1}\xi_{n-1}|\xi|)}f(\xi)\,d\xi\$
+where $B^{n-1}$ is the unit ball centered at the origin. By a short duality argument, we see that conjecture 1.1 is equivalent to
 Conjecture 1.2.
-Let $n\geq 3$. For all $p,q>0$ satisfy (1.2), there holds
-$\|Ef\|_{p}\lesssim_{n,p,q}\|f\|_{q},\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$ (1.3)
+Let $n\geq 3$. For all $p,q>0$ satisfy (1.2),
+$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
 Conjecture 1.2 has so far been proven only in dimensions $n=3,4,5$ with contributions from Barcelo [1], Wolff [18], and Ou-Wang [15].
@@
 Theorem 1.3 ([15], Theorem 1).
-For $n\geq 3$, (1.3) holds whenever
-$q\geq p>\left\{\begin{array}[]{rcl}4&\mbox{if}&n=3\\ 2\frac{3n+1}{3n-3}&\mbox{if}&n>3\quad odd\\ 2\frac{3n}{3n-4}&\mbox{if}&n>3\quad even\end{array}\right.$ (1.4)
+Let $n\geq 3$. For all $p,q>2$ satisfy
+$q\geq p>\left\{\begin{array}[]{rcl}4&\ \mathrm{if}&n=3\\ 2\frac{3n+1}{3n-3}&\mathrm{if}&n>3\quad\mathrm{odd}\\ 2\frac{3n}{3n-4}&\mathrm{if}&n>3\quad\mathrm{even}\end{array}\right$.
+there holds
+$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
 Theorem 1.4 ([15], Theorem 2).
-For $n\geq 3$, (1.3) holds whenever the tuple $(p,q,k)$ is admissible in the sense that $2<q<p$, $2\leq k\leq n$ and
-$\left\{\begin{array}[]{rcl}p>2\frac{n+2}{n},q^{\prime}\leq\frac{n-2}{n}p&\mbox{if}&k=2\\ p>2\frac{n+k}{n+k-2},p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\mbox{if}&k\geq 3\end{array}\right.$
+Let $n\geq 3$. Suppose the tuple $(p,q,k)$ is admissible in the sense that $2<q<p$, $2\leq k\leq n$ and
+$\left\{\begin{array}[]{rcl}p>2\frac{n+2}{n},q^{\prime}\leq\frac{n-2}{n}p&\mathrm{if}&k=2\\ p>2\frac{n+k}{n+k-2},p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\mathrm{if}&k\geq 3\end{array}\right$.
+then
+$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
 The Fourier restriction problem has been the focus of extensive study for decades.
@@
 where $\lambda=2.596$… and $B_{R}$ is an arbitrary ball in $\mathbb{R}^{n}$ of radius $R$.
-By a real interpolation of restrict-type and a standard $\epsilon$-removal argument (see [6]), we see that
+By a real interpolation of restrict-type and a standard $\epsilon$-removal argument (see [17]), one has
 Corollary 1.6.
-For $n\geq 3$, (1.3) holds whenever $q\geq p$ and (1.6) holds.
+Let $n\geq 3$. For all $q\geq p>2$ satisfy (1.4), there holds
+$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
 Guth [5, 6] introduced polynomial partitioning to the study of Fourier restriction and established the improved results for paraboloids.
 Later, his results were subsequently refined by Wang[19], Wang-Wu[20], Hickman–Rogers [11], Hickman–Zahl [12], and Guo-Wang-Zhang [7].
-In [12, 13] the authors reformulated Guth’s induction-on-scale argument into a recursive algorithm, utilizing various Kakeya-type geometric results from [15, 16, 17], including the nested polynomial Wolff axiom, to bound the number of directions of the tubes those stay in varieties.
+In [11, 12] the authors reformulated Guth’s induction-on-scale argument into a recursive algorithm, utilizing various Kakeya-type geometric results from [13, HRZ, 24], including the nested polynomial Wolff axiom, to bound the number of directions of the tubes those stay in varieties.
 Beyond paraboloid restriction, polynomial partitioning has also been applied to many related problems in harmonic analysis, such as Bochner-Riesz conjecture (see [21, 8]), the Schrödinger maximal estimate (see [3]), and cone restriction (see [15]).
@@
 However, if we were to apply this strategy to the cone, the number of tube directions would not be well-bounded, as the number of sectors on a cone is significantly fewer than the number of caps on a paraboloid.
-Instead, we make a leaf backtrack to its $(n-1)$- dimensional ancestor.
+Instead, we make a leaf backtrack to its (n1)(n-1)- dimensional ancestor.
 We will provide a more detailed discussion on this topic in Section 3.
@@
 We inductively define $\delta$ and $\delta_{j}$ for $0\leq j\leq n$ as bellows:
-$\delta:=\epsilon^{10},\quad\delta_{j}:=\delta_{j-1}^{10}\quad\mbox{for}\quad 1\leq j\leq n,\quad\delta:=\delta_{n}^{10}$
+$\delta_{0}:=\epsilon^{10},\quad\delta_{j}:=\delta_{j-1}^{10}\quad\mbox{for}\quad 1\leq j\leq n,\quad\delta:=\delta_{n}^{10}$
 In addition, we say $A\lesssim_{L}B$, $B\gtrsim_{L}A$, or $A=O_{L}(B)$ if $A\leq C_{L}B$ for some constant $C_{L}$ depending only on $L$.
@@
 We say $A\ll B$ or $B\gg A$ if $CA\leq B$ for some choice of $C\geq 1$ that can be taken to be as large as desired provided it is admissible.We say $A\lessapprox B$ if $A\lesssim R^{O(\delta_{0})}B$.
-Let #$\#A$ denote the cardinality of a finite set $A$.
+Let #A\#A denote the cardinality of a finite set $A$.
 Let $N_{r}E$ denote the $r$-neighbourhood of a set $E$.
@@
 Let $T_{x}Z$ denote the tangent space of $Z$ at $x$.
-We say $A^{\prime}\subseteq A$ is a refinement of $A$ if #$\#A\lessapprox\#A^{\prime}$.
+We say $A^{\prime}\subseteq A$ is a refinement of $A$ if $A\lessapprox\#A^{\prime}$.
 We say $A$ = RapDec($r$) if $A\lesssim_{N}r^{-N}$.
@@
 2.2 Wave Packet Decomposition
-The wave packet decomposition is a crucial tool in our argument. Here we recall its definition and basic properties from section 2.2 and 5.2.1 of [4].
+The wave packet decomposition is a crucial tool in our argument. Here we recall its definition and basic properties from Sections 2.2 and 5.2.1 of [15].
 […]
@@
 $\|f_{\mathcal{T}}\|_{2}^{2}\sim\sum_{(\theta,v,l)\in\mathcal{T}}\|f_{\theta,v}^{l}\|_{2}^{2}$
-Like the case of paraboloid [12], there is also some local versions of the $L^{2}$-orthogonality for cone, which are a direct corollary of lemma 2.1:
+Like in the case of paraboloid [11], there are also some local versions of the $L^{2}$-orthogonality for the cone, which are direct corollaries of Lemma 2.1:
 Corollary 2.2.
@@
 ${\psi}_{y}\left(\xi\right):=y_{1}{\xi}_{1}+\dots+y_{n-1}{\xi}_{n-1}+y_{n}\left|\xi\right|$
-One can verify that $Eg\left(x\right)=E\tilde{g}\left(\tilde{x}\right)$ where $x=y+\tilde{x}$ for some $\tilde{x}\in B\left(0,\rho\right)$. Next, we break $\tilde{g}$ into scale $\rho$ wave packets
+[…]
 $\tilde{g}=\sum_{\left(\zeta,w,L\right)\in\mathbb{T}\left[\rho\right]}{{\tilde{g}}^{L}_{\zeta,w}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$
-Let $T^{L}_{\zeta,w}\left(y\right)$denote the $\rho$-tube corresponding to $\left(\zeta,w,L\right)$ in the $x$ coordinate (contained in $B\left(y,\rho\right)$).
+Let $T^{L}_{\zeta,w}\left(y\right)\$denote the $\rho$-tube corresponding to $\left(\zeta,w,L\right)$ in the $x$ coordinate (contained in $B\left(y,\rho\right)$).
 Define ${\widetilde{\mathbb{T}}}_{\theta,v,l}$ as
@@
 ${\mathrm{D}\mathrm{ist}}_{H}\left({\zeta}_{0},\zeta\right)\lesssim{\rho}^{-\frac{1}{2}},P^{L}_{\zeta,w}\subseteq B\left(v_{0},r^{\frac{1}{2}+2\delta}\right),\ \ \forall\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}},\quad g=\sum_{{\zeta}_{0},v_{0}}{g_{{\zeta}_{0},v_{0}}}+\mbox{RapDec}\left(r\right){\left\|g\right\|}_{2}$ (2.2)
-Since each $\left(\zeta,w,L\right)$ belongs to $\sim 1$ ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, the Lemma 2.1 implies
+Since each $\left(\zeta,w,L\right)\$belongs to $\sim 1$ ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, the Lemma 2.1 implies
 ${\left\|g\right\|}^{2}_{2}\sim\sum_{{\zeta}_{0},v_{0}}{{\left\|g_{{\zeta}_{0},v_{0}}\right\|}^{2}_{2}},\ \ {\left\|\tilde{g}\right\|}^{2}_{2}\sim\sum_{{\zeta}_{0},v_{0}}{{\left\|{\tilde{g}}_{{\zeta}_{0},v_{0}}\right\|}^{2}_{2}}$ (2.3)
-From the lemma 2.5 and 2.6, one has
+From Lemma 2.5 and 2.6, one has
 Corollary 2.7 ([15], Lemma 5.6).
@@
 $\nabla P_{1}\left(x\right)\wedge\dots\wedge\nabla P_{n-m}\left(x\right)\neq 0,\ \ \forall x\in Z\left(P_{1},\dots,P_{n-m}\right)$
-The degree of the transverse complete intersection is defined as ${\mathop{\mathrm{max}}_{1\leq j\leq n-m}{\mathrm{deg}P_{j}\ }\ }$.
+The degree of the transverse complete intersection is defined as $\mathop{\mathrm{max}}_{1\leq j\leq n-m}{\mathrm{deg}P_{j}\ }\$}.
 From the above definition, if $\mathrm{Z}:=Z\left(P_{1},\dots,P_{n-m}\right)$ is a transverse complete intersection, then $\mathrm{Z}$ is a smooth manifold of dimension $m$, in other words, $\mathrm{Z}$ has a tangent plane $T_{x}\mathrm{Z}$ at every point $x\in\mathrm{Z}$.
@@
 We use $\mathbb{T}\left[\boldsymbol{\mathrm{Z}},B\left(y,r\right)\right]$ to denote the $r$-tube tangent to $\boldsymbol{\mathrm{Z}}$ in $B\left(y,r\right)$.
-In secton 3 we need the following result, which can be roughly taken as a continuum version of the Bezout’s theorem, to controls the transverse intersections between a tube and an algebraic variety,
+In Section 3 we need the following result, which can be roughly taken as a continuum version of the Bezout’s theorem, to controls the transverse intersections between a tube and an algebraic variety,
 Lemma 2.10 ([6], Lemma 5.7).
@@
 ${\mathrm{dim}Z_{i}\ }=i,\quad\forall i$
-$N_{r^{1/2+{\delta}_{s}}_{s}}Z_{s}\subseteq\dots\subseteq N_{r^{1/2+{\delta}_{t}}_{t}}Z_{t}$
-$r_{s}\leq r_{s+1}\leq\dots\leq r_{t}$
-The parameter $\left(t-s+1\right)$ is referred to as the level of multigrain.
-The degree of $\overrightarrow{S}$ is defined as ${\mathop{\mathrm{max}}_{1\leq i\leq l}Z_{i}\ }$.
+$B(y_{i},r_{i})\cap B(y_{i+1},r_{i+1})\neq\emptyset,\quad\forall i$
+$N_{Cr^{1/2+{\delta}_{s}}_{s}}Z_{s}\subseteq\dots\subseteq N_{Cr^{1/2+{\delta}_{t}}_{t}}Z_{t}\mathrm{\ for\ some\ }C>0$
+$0<r_{s}\leq r_{s+1}\leq\dots\leq r_{t}$
+The parameter $\left(t-s+1\right)$ is called the multigrain level.
+The degree of $\overrightarrow{S}$ is defined as $\mathop{\mathrm{max}}_{1\leq i\leq l}Z_{i}\$}.
 The multiscale of $\overrightarrow{S}$ is the tuple $\overrightarrow{r}=\left(r_{s},r_{s+1},\dots,r_{t}\right)$.
@@
 $L\cap B\left(y_{i},2r_{i}\right)\subseteq N_{{\rho}_{j}}Z_{i}\cap B\left(y_{i},2r_{i}\right)$
-for any $L\in\mathcal{L}$, then what is the maximum possible value of $\#\mathcal{L}$? Here $\mathcal{L}$ is said to be $r^{-1/2}$-direction separated if
+for any $L\in\mathcal{L}$, then what is the maximum possible value of $\mathcal{L}$? Here $\mathcal{L}$ is said to be $r^{-1/2}$-direction separated if
 $\angle\left(L_{1},L_{2}\right)\gtrsim r^{-\frac{1}{2}},\ \ \mathrm{if}\ L_{1},L_{2}\in\mathcal{L}\boldsymbol{\ }\mathrm{and}\boldsymbol{\ }L_{1}\neq L_{2}$
@@
 Let $r\geq 1$. Let $Z\subset{\mathbb{R}}^{n}$ be an algebraic variety of dimension $m$ and degree at most $d$. If $\mathcal{T}\subseteq\mathbb{T}\left[r\right]$satisfies: $\forall T\in\mathcal{T}$,
-$T\cap B_{r}\cap N_{r^{1/2+{\delta}_{m}}}Z\neq\emptyset$
+$T\cap B_{r}\cap N_{r^{1/2+{\delta}_{m}}}Z\neq\emptyset\$
 $T\cap 2B_{r}\subseteq N_{r^{1/2+{\delta}_{m}}}Z$
 then
-#$\#\mathcal{T}{\lesssim}_{d}r^{O\left({\delta}_{m}\right)}r^{\frac{m-1}{2}}$
+$\mathcal{T}{\lesssim}_{d}r^{O\left({\delta}_{m}\right)}r^{\frac{m-1}{2}}$
 […]
@@
 For $i=m$, we set
-$\mathbb{T}\left[{\overrightarrow{S}}^{m}_{m}\right]:=\mathbb{T}\left[S_{m}\right]$
-For $m<i\leq n-1$, let $\mathbb{T}\left[{\overrightarrow{S}}^{i}_{m}\right]$ be the set of all $r_{i}$-tubes $T^{l}_{\theta,v}$ satisfying the following conditions:
-(a). $T^{l}_{\theta,v}\in\mathbb{T}\left[Z_{i},B\left(y_{i},r_{i}\right)\right]$
+$\mathbb{T}\left[{\overrightarrow{S}}^{m}_{m}\right]:=\mathbb{T}\left[S_{m}\right]\ \ \mathrm{\Theta}\left[{\overrightarrow{S}}^{m}_{m}\right]:=\left\{\theta:\exists T^{l}_{\theta,v}\in\mathbb{T}\left[{\overrightarrow{S}}^{m}_{m}\right]\ \mathrm{for\ some\ }v,l\right\}$
+For $m<i\leq n$, let $\mathbb{T}\left[{\overrightarrow{S}}^{i}_{m}\right]$ be the set of all $r_{i}$-tubes $T^{l}_{\theta,v}$ satisfying the following conditions:
+(a). $T^{l}_{\theta,v}\in\mathbb{T}\left[Z_{i},B\left(y_{i},r_{i}\right)\right]\$
 (b). There exists $\exists T^{L}_{\zeta,w}\in\mathbb{T}\left[Z_{i-1},B\left(y_{i-1},r_{i-1}\right)\right]$ such that
 $\quad{\mathrm{D}\mathrm{ist}}_{H}\left(w,\ \zeta\right)\lesssim r^{-1/2}_{i-1},\quad{\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w},\ T^{l}_{\theta,v}\cap B\left(y_{i},r_{i}\right)\right)\lesssim r^{1/2+\delta}_{i}$
+and
+[…]
 Finally, we set
-$\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\mathbb{T}\left[{\overrightarrow{S}}^{n-1}_{m}\right],\ \ \mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\left\{\theta:\exists T^{l}_{\theta,v}\in\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{m}\right]\ \mathrm{for\ some\ }v,l\right\}$
-One can observe the second condition in definition 2.15 is weaker than lemma 2.5.
+$\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\mathbb{T}\left[{\overrightarrow{S}}^{n-1}_{m}\right],\ \ \mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\ \mathrm{\Theta}\left[{\overrightarrow{S}}^{n-1}_{m}\right]$
+One can observe that the second condition in Definition 2.15 is weaker than Lemma 2.5.
 However, this is good enough for our arguments.
-In [13, 19, 25] a non-trivial extension of Theorem 2.14, which is a corollary to lemma 2.11 in [17], was used to establish a new result on the Fourier restriction, the Bochner-Riesz problem, and the Stein’s square functions respectively.
-By setting $r_{0}=r_{1}$ in this extension (see [13], Lemma 3.7), one has
-Lemma 2.16 (Nested polynomial Wolff axioms).
-Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$,
-${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d,\ \ \forall i\in\mathbb{N}\cap\left[m,n\right)$
+In [12, 8, 10] a nontrivial extension of Theorem 2.14, which is a corollary to Lemma 2.11 in [24], was used respectively to establish a new result on the Bochner-Riesz problem, Stein’s square function, and the Fourier restriction for paraboloid.
+The result is shown below:
+Lemma 2.16 ([12], Lemma 3.7, Nested polynomial Wolff axioms).
+Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$ such that for all $i\in\mathbb{N}\cap\left[m,n\right]$,
+${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d$
 then
-#$\#\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]{\lesssim}_{d}r^{O\left({\delta}_{0}\right)}_{n-1}\left(\prod^{n-2}_{j=m}{r^{-\frac{1}{2}}_{j}}\right)r^{\frac{n-2}{2}}_{n-1}$
-In [13, 19, 25] $S_{m}$ was passed back to $S_{n}:=\left({\mathbb{R}}^{n},B_{R}\right)$, and thus
-$\mathrm{\Theta}\left[{\overrightarrow{S}}_{m}\right]{\lessapprox}_{d}\left(\prod^{n-1}_{j=m}{r^{-1/2}_{j}}\right)R^{\left(n-1\right)/2}$
-Since $\mathrm{\Theta}\left[R\right]\sim R^{\left(n-1\right)/2}$ in the case of paraboloids, this bound gives a gain of $\prod^{n-1}_{j=m}{r^{-1/2}_{j}}$.
-However, in the case of cone this doesn’t guarantee a gain due to $\mathrm{\Theta}\left[R\right]\sim R^{\left(n-2\right)/2}$.
-This is the reason for passing $S_{m}$ back to $S_{n-1}$ rather than $S_{n}$ in our arguments.
+$\mathrm{\Theta}\left[{\overrightarrow{S}}^{n}_{m}\right]{\lesssim}_{d}r^{O\left({\delta}_{0}\right)}_{n}\left(\prod^{n-1}_{j=m}{r^{-\frac{1}{2}}_{j}}\right)r^{\frac{n-1}{2}}_{n}$
+A direct corollary of this lemma is
+Corollary 2.17.
+Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$ such that for all $i\in\mathbb{N}\cap\left[m,n\right]$,
+${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d$
+then
+$\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]{\lesssim}_{d}r^{O\left({\delta}_{0}\right)}_{n-1}\left(\prod^{n-2}_{j=m}{r^{-\frac{1}{2}}_{j}}\right)r^{\frac{n-2}{2}}_{n-1}$
+Proof.
+Apply Lemma 2.16 to ${\overrightarrow{S}}:=(S_{m},...,S_{n-1},({\mathbb{R}}^{n},B(y_{n-1},r_{n-1})))$. ∎
+In [12, 8, 10] $S_{m}$ is traced back to $S_{n}:=\left({\mathbb{R}}^{n},B_{R}\right)$, and thus Lemma 2.16 gives
+$\mathrm{\Theta}\left[{\overrightarrow{S}}^{n}_{m}\right]{\lessapprox}_{d}\left(\prod^{n-1}_{j=m}{r^{-1/2}_{j}}\right)R^{\left(n-1\right)/2}$
+Since the tubes have $\sim R^{\left(n-1\right)/2}$ directions in the case of paraboloids, this bound yields a gain of $\prod^{n-1}_{j=m}{r^{-1/2}_{j}}$.
+However, in the case of cone, this does not guarantee a gain due to $\mathrm{\Theta}\left[R\right]\sim R^{\left(n-2\right)/2}$.
+This is the why we need Corllary 2.17
 2.4 $k$-broad Norms
@@
 […]
-Basic Properties The parameter $A$ is used to make $k$-board norm behave like a regular $L^{p}$-norm. Here we recall some inequalities about .${\left\|.\right\|}_{BL^{p}_{k,A}}$.
+Basic Properties The parameter $A$ is used to make $k$-board norm behave like a regular $L^{p}$-norm. Here we recall some inequalities about ${\left\|.\right\|}_{BL^{p}_{k,A}}$.
 Lemma 2.18 (Triangle inequality, [15], Lemma 3.1).
@@
 Lemma 2.21 (Vanishing property, [15], Lemma 4.5).
-Suppose $Ef$ is tangent to a transverse complete intersection $\boldsymbol{Z}$ in $B_{r}$. If ${\mathrm{deg}\boldsymbol{Z}\lesssim 1\ }$ and ${dim\boldsymbol{Z}\ }$ $\leq k-1$, then
+Suppose $Ef$ is tangent to a transverse complete intersection $\boldsymbol{Z}$ in $B_{r}$. If $\mathrm{deg}\boldsymbol{Z}\lesssim 1\$} and $dim\boldsymbol{Z}\$} $\leq k-1$, then
 ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{r}\right)}\lesssim\mathrm{RapDec}\left(r\right){\left\|f\right\|}_{2}$
@@
 $Eg_{b}\left(x\right)\approx{\chi}_{N_{{\rho}^{1/2+{\delta}_{m}}}\left({\boldsymbol{\mathrm{Z}}}_{b}\right)}\left(x\right)Eg\left(x\right)$ (2.5)
-If $\boldsymbol{\mathrm{Z}}$ is a plane, then the existence of such a decomposition follows from the Lemma 2.6.
-However, this is not easy for ${\mathrm{deg}\boldsymbol{\mathrm{Z}}\ }>1$, since the shapes of $\boldsymbol{\mathrm{Z}}$ is irregular.
+If $\boldsymbol{\mathrm{Z}}$ is a plane, then such a decomposition exists due to Lemma 2.6.
+However, this question will be tricky if ${\mathrm{deg}\boldsymbol{\mathrm{Z}}\ }>1$, since the shapes of $\boldsymbol{\mathrm{Z}}$ is irregular.
 A rigorous version of (2.4) and (2.5) is given by the following statement.
@@
 and
-$\ \ \sum_{b\in\mathcal{B}}{{\left\|g^{\prime}_{b}\right\|}^{2}_{2}}\lesssim{\left\|g^{\prime}\right\|}^{2}_{2}$
+[…]
 where
@@
 Theorem 2.23.
-Let $22\leq k\leq n-$ 2 and
+Let $2\leq k\leq n-2$ and
 $p>p_{n}\left(k\right):=2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}}$ (2.6)
@@
 For any $R>1$, the estimate (1.3) holds whenever the estimate (2.7) holds, for $p$ in the range
-[…]
+$p\left(k,n\right)<p\leq\frac{2n}{n-2},p\left(k,n\right):=\left\{\begin{array}[]{c}2+\frac{2}{n-2},\ if\ 2\leq k\leq 3\\ 2+\frac{4}{2n-k-1},\ if\ k>3\end{array}\right$. (2.8)
 We choose a value of $k$ to optimizes the range of $p$ in (1.3). By applying the arguments in [12] to (2.6), one has
@@
 $p\left(k,n\right)=2+\frac{4}{2-v}n^{-1}+O\left(n^{-2}\right)$ (2.9)
-To obtain the best possible estimate in the theorem 1.5, $v$ should satisfy
+To obtain the best possible estimate in Theorem 1.5, $v$ should satisfy
 $\frac{6}{2+v^{3/2}}=\frac{4}{2-v}$
@@
 ${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}\gg R^{O\left({\delta}_{n-1}\right)}{\left(\frac{r}{R}\right)}^{1/2}{\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$
-If ${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$ doesn’t vanish, then the Lemma 2.4 gives
+If ${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$ doesn’t vanish, then Lemma 2.4 gives
 ${\left\|g_{b}\right\|}^{2}_{2}\gg R^{O\left({\delta}_{n-1}\right)}{\left(\frac{r}{R}\right)}^{1/2}{\left\|g\right\|}^{2}_{2}$
@@
 For a vector space $V$ spanned by ${\left\{v_{i}=\left(x_{i,1},x_{i,2},\dots,-x_{i,n}\right)\right\}}^{m}_{i=1}$, $V^{+}$ is defined by
-[…]
+$V^{+}:=\left\{\sum^{m}_{i=1}{a_{i}v^{+}_{i}}:a_{i}\in\mathbb{R}\mathrm{,}v^{+}_{i}=\left(x_{i,1},x_{i,2},\dots,x_{i,n}\right),\ \mathrm{f}\mathrm{or\ }1\leq i\leq m\right\}$
 Now we cover $B_{\rho}$ by finitely overlapping $r^{1/2+{\delta}_{m}}$-ball $B$.
@@
 Lemma 2.25 ([15], Lemma 5.8).
-Suppose $B\in{\mathcal{B}}_{b}$̇et
+Suppose $B\in{\mathcal{B}}_{b}$. Let
 $\mathbb{T}\left[{\theta}^{\prime}\right]:=\left\{\left(\theta,v,l\right):\theta={\theta}^{\prime}\right\},\ \ {\mathrm{\Theta}}_{B}:=\left\{\theta:\mathbb{T}\left[\theta\right]\cap{\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}\neq\emptyset\right\}$
@@
 $g_{\mathrm{ess}}:=\sum_{\left({\zeta}_{0},v_{0}\right)\in{\mathbb{T}}_{\mathrm{ess}}}{g_{{\zeta}_{0},v_{0}}},\ \ g_{\mathrm{tail}}:=\sum_{\left({\zeta}_{0},v_{0}\right)\in{\mathbb{T}}_{\mathrm{tail}}}{g_{{\zeta}_{0},v_{0}}}$ (2.10)
-Suppose $A$ is large enough. By the definition of $k$-board norm, and lemma 2.24,
+Suppose $A$ is large enough. By the definition of $k$-board norm and Lemma 2.25,
 ${\left\|Eg_{\mathrm{tail}}\right\|}^{p}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}\lesssim\sum_{B\in{\mathcal{B}}_{b}}{{\left\|Eg_{\mathrm{tail}}\right\|}^{p}_{BL^{p}_{k,A/2}\left(B\right)}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{p}_{2}\lesssim\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{p}_{2}$ (2.11)
@@
 ${\left\|F\right\|}_{L^{1}\left(O\right)}\sim d^{-n}{\left\|F\right\|}_{L^{1}\left(B_{r}\right)},\ \forall O\in\mathrm{cell}\left(P\right)$
-Suppose we have a polynomial $P$ described in the theorem 2.27.
+Suppose we have a polynomial $P$ described in Theorem 2.28.
 Let ${\delta}^{\prime}>0$ be a small number.
@@
 Cellular case. There exists a polynomial $P:{\mathbb{R}}^{n}\to\mathbb{R}$ of degree $O\left(d\right)$ with the following properties:
-(1).$\mathcal{O}\sim d^{m}$, and each $O\in\mathcal{O}$ is contained in a ball of radius $r/2$.
+[…]
 (2).For each $O\in\mathcal{O}$, ${\left\|F\right\|}_{L^{1}\left(O\right)}\sim d^{-m}{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
@@
 ${\left\|F\right\|}_{L^{1}\left(B_{r}\right)}\lesssim{\left\|F\right\|}_{L^{1}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Y}}\right)}$
-Now we apply the theorem 2.28 to the $k$-board norm estimates:
+Now we apply Theorem 2.29 to the $k$-board norm estimates:
 (1). If the cellular case holds, then it follows that
@@
 For a $r$-tube $T^{l}_{\theta,v}$,
-#$\#\left\{O\in{\mathcal{O}}^{\prime}:O\cap T^{l}_{\theta,v}\neq\emptyset\right\}\lesssim d$
+$\left\{O\in{\mathcal{O}}^{\prime}:O\cap T^{l}_{\theta,v}\neq\emptyset\right\}\lesssim d$
 3 Finding Polynomial Structure
@@
 The iteration ends when we arrive in the minimal dimensional case.
-The lemma 2.16 is used to establish the bound.
+Corollary 2.17, the nested polynomial Wolff axioms, is used to establish the bound.
 This algorithm corresponds to the induction on the dimension.
@@
 $\bullet$ History: A word ${\hslash}_{j}$ of length $j$ in the alphabet $\left\{\mathrm{a,c}\right\}$, which records how the cell $O_{j}\in{\mathcal{O}}_{j}$ were constructed by the iterated polynomial partitioning.
-Let #$\#\mathrm{a}\left(j\right)$ denote the number of ‘a’ in ${\hslash}_{j}$, and #$\#\mathrm{c}\left(j\right)$ denote the number of ‘c’ in ${\hslash}_{j}$.
+Let $\mathrm{a}\left(j\right)$ denote the number of ‘a’ in ${\hslash}_{j}$, and $\mathrm{c}\left(j\right)$ denote the number of ‘c’ in ${\hslash}_{j}$.
 $\bullet$ Spatial scale ${\rho}_{j}$: The ${\rho}_{j}$ is determined by the initial scale $r$ and ${\hslash}_{j}$. Define an auxiliary parameter ${\widetilde{\delta}}_{m-1}$ by
@@
 $\bullet$ Cells: A family of subsets ${\mathcal{O}}_{j}$ of ${\mathbb{R}}^{n}$.
-Each cell $O_{j}\in{\mathcal{O}}_{j}$ is contained in some ${\rho}_{j}$-ball $B_{O_{j}}=B\left(y_{O_{j}},{\rho}_{j}\right)$.
+Each cell $O_{j}\in{\mathcal{O}}_{j}$ is contained in some ${\rho}_{j}$-ball $B_{O_{j}}=B\left(y_{O_{j}},{\rho}_{j}\right)\$.
 $\bullet$ A collection of functions ${\left(f_{O_{j}}\right)}_{O_{j}\in{\mathcal{O}}_{j}}$.
 For each cell $O_{j}$ there exists a translate ${\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{:}=\boldsymbol{\mathrm{Z}}\boldsymbol{\mathrm{+}}x_{{\mathcal{O}}_{j}}$ such that $f_{O_{j}}$ is tangent to the grain $\left({\boldsymbol{\mathrm{Z}}}_{O_{j}},B_{O_{j}}\right)$
-$\bullet$ A large integer $d$ which depends only on the admissible parameters and ${\mathrm{deg}\boldsymbol{\mathrm{Z}}\ }$.
+$\bullet$ A large integer $d$ which depends only on the admissible parameters and $\mathrm{deg}\boldsymbol{\mathrm{Z}}\$}.
 $C^{\mathrm{I}}_{j,\delta}\left(d,r\right):=d^{\#\mathrm{c}\left(j\right)\delta}{{\mathrm{log}}^{\#\mathrm{a}\left(j\right)\left(1+\delta\right)}r\ },C^{\mathrm{II}}_{j,\delta}\left(d\right):=d^{\#\mathrm{c}\left(j\right)\delta+n\#\mathrm{a}\left(j\right)\left(1+\delta\right)}$
@@
 $\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}\leq\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$ (3.6)
-Define ${\hslash}_{j+1}$ by appending the $\mathrm{c}$ to ${\hslash}_{j}$. Then, the definition of the parameters gives
+Define ${\hslash}_{j+1}$ by appending the $\mathrm{c}$ to ${\hslash}_{j}$. Now
 $\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right),\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right)+1,\ \ A_{j+1}=A_{j}$
@@
 1. $\mathcal{O}\left(O_{j}\right)\ \sim\ d^{m}$
-2. For all $O\in\mathcal{O}$, $O$ is included in a ${\rho}_{j+1}$-ball $B_{O}:=B\left(y_{O},{\rho}_{j+1}\right)$.
+2. For all $\in\mathcal{O}, OO$ is included in a ${\rho}_{j+1}$-ball $B_{O}:=B\left(y_{O},{\rho}_{j+1}\right)$.
 3. ${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O\right)}\sim d^{-m}{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)},\ \ \forall O\in\mathcal{O}$
@@
 $f_{O}:=\sum_{\begin{array}[]{c}\left(\theta,v,l\right)\in\mathbb{T}\left[{\rho}_{j}\right]\\ T^{l}_{\theta,v}\cap O\neq\emptyset\end{array}}{{\left(f_{O_{j}}\right)}^{l}_{\theta,v}}$
-By lemma 2.29, each $T^{l}_{\theta,v}$ intersects $O\left(d\right)$ cells, and thus the lemma 2.1 gives
+By Lemma 2.30, each $T^{l}_{\theta,v}$ intersects $O\left(d\right)$ cells, and thus Lemma 2.1 gives
 $\sum_{O\in\mathcal{O}\left(O_{j}\right)}{{\left\|f_{O}\right\|}^{2}_{2}}\lesssim d{\left\|f_{O_{j}}\right\|}^{2}_{2}$
-By pigeonholing, there exists a refinement ${\mathcal{O}}^{\prime}\left(O_{j}\right)\subseteq\mathcal{O}\left(O_{j}\right)$ with #$\#{\mathcal{O}}^{\prime}\ \sim\ d^{m}$ such that
+By pigeonholing, there exists a refinement ${\mathcal{O}}^{\prime}\left(O_{j}\right)\subseteq\mathcal{O}\left(O_{j}\right)$ with ${\mathcal{O}}^{\prime}\ \sim\ d^{m}$ such that
 ${\left\|f_{O}\right\|}^{2}_{2}\lesssim d^{-m+1}{\left\|f_{O_{j}}\right\|}^{2}_{2},\ \ \forall O\in{\mathcal{O}}^{\prime}\left(O_{j}\right)$
-Since ${\rho}_{j+1}\sim{\rho}_{j}$, there exists $\mathcal{B}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ with # $\#\mathcal{B}\ \sim\ 1$ such that
+Since ${\rho}_{j+1}\sim{\rho}_{j}$, there exists $\mathcal{B}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ with # $\mathcal{B}\ \sim\ 1$ such that
 $N_{{\rho}^{1/2+{\delta}_{m}}_{j}}{\boldsymbol{\mathrm{Z}}}_{O_{j}}\subseteq\bigcup_{b\in\mathcal{B}}{N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{+}b\right)}$
@@
 ${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq C{\rho}^{-\frac{1}{2}}_{j+1},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y_{O_{j+1}}\right),\ T^{l}_{\theta,v}\cap B\left(y_{O_{j+1}},{\rho}_{j+1}\right)\right)\leq C{\rho}^{1/2+\delta}_{j}$
-for some $\left(\zeta,w,L\right)\in\mathcal{T}$. Here $C$ is the absolute constant that appears in the property 4. Combine (3.5) and (3.7),
+for some $\left(\zeta,w,L\right)\in\mathcal{T}$. Here $C$ is the absolute constant that appears in Property 4. Combine (3.5) and (3.7),
 ${\left\|f_{O_{j+1},\ \mathcal{T}}\right\|}^{2}_{2}\lesssim d^{-\delta}C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
@@
 $\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}\leq\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$
-Define ${\hslash}_{j+1}$ by appending the $\mathrm{a}$ to ${\hslash}_{j}$. Then, the definition of the parameters gives
-#$\#\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right)+1,\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right),\ \ A_{j+1}=A_{j}/3$
+Define ${\hslash}_{j+1}$ by appending the $\mathrm{a}$ to ${\hslash}_{j}$. Now
+$\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right)+1,\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right),\ \ A_{j+1}=A_{j}/3$
 $C^{\mathrm{I}}_{j+1,\delta}\left(d,r\right)={{\mathrm{log}}^{1+\delta}r\ }C^{\mathrm{I}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{II}}_{j+1,\delta}\left(d\right)=d^{n\left(1+\delta\right)}C^{\mathrm{II}}_{j,\delta}\left(d\right)$
@@
 ${\mathop{\mathrm{max}}_{\begin{array}[]{c}S\in\mathcal{S}\\ \theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}\end{array}}{\left\|f_{S}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq C_{\mathrm{tang}}{\mathop{\mathrm{max}}_{\begin{array}[]{c}O_{j}\in{\mathcal{O}}_{j}\\ \theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}\end{array}}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (3.12)
-Since #$\#{\mathcal{B}}_{O_{j}}\lesssim r^{n{\widetilde{\delta}}_{m-1}}$ for all $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$, the lemma 2.1 gives
+Since ${\mathcal{B}}_{O_{j}}\lesssim r^{n{\widetilde{\delta}}_{m-1}}$ for all $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$, Lemma 2.1 gives
 $\sum_{S\in\mathcal{S}}{{\left\|f_{S}\right\|}^{2}_{2}}\leq C_{\mathrm{tang}}r^{n{\widetilde{\delta}}_{m-1}}\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|f_{O_{j}}\right\|}^{2}_{2}}$ (3.13)
-Also, for $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$ and $S=\left(Y,B\right)\in\mathcal{S}$ with $B\in{\mathcal{B}}_{O_{j}}$, the lemma 2.1 and corollary 2.7 gives
+Also, for $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$ and $S=\left(Y,B\right)\in\mathcal{S}$ with $B\in{\mathcal{B}}_{O_{j}}$, Lemma 2.1 and Corollary 2.7 gives
 ${\left\|f_{S,\mathcal{T}}\right\|}^{2}_{2}\leq C_{\mathrm{tang}}{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$ (3.14)
@@
 A trivial observation is
-#$\#\mathrm{c}\left(J\right)\lesssim{\mathrm{log}r\ },\ \ \#\mathrm{a}\left(J\right)\lesssim{\delta}^{-1}_{m-1}{\mathrm{log}{\delta}^{-1}_{m-1}\ }$ (3.18)
+$\mathrm{c}\left(J\right)\lesssim{\mathrm{log}r\ },\ \ \#\mathrm{a}\left(J\right)\lesssim{\delta}^{-1}_{m-1}{\mathrm{log}{\delta}^{-1}_{m-1}\ }$ (3.18)
 Hence, the accumulation of error is still negligible after the algorithm terminates. We omit the details.
@@
 $\bullet$ For $m\leq l\leq n$ a family ${\overrightarrow{\mathcal{S}}}_{l}$ of level $\left(n-l+1\right)$ multigrain.
-Each ${\overrightarrow{S}}_{l}=\left(\left(Z_{l},B_{r_{l}}\right),\dots,\left(Z_{n},B_{r_{n}}\right)\right)\in{\overrightarrow{\mathcal{S}}}_{l}$ has multiscale ${\overrightarrow{r}}_{l}=\left(r_{l},\dots,r_{n}\right)$ and bounded degree.
+Each ${\overrightarrow{S}}_{l}=\left(\left(Z_{l},B_{r_{l}}\right),\dots,\left(Z_{n},B_{r_{n}}\right)\right)\in{\overrightarrow{\mathcal{S}}}_{l}$ has a multiscale ${\overrightarrow{r}}_{l}=\left(r_{l},\dots,r_{n}\right)$ and a bounded degree.
 Also, for $l\leq s\leq t\leq n$:
@@
 For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}$ we apply [alg 1] to the $k$-board norm ${\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}$.
-Let #$\#\mathrm{a}\left[{\overrightarrow{S}}_{l}\right]$ and #$\#\mathrm{c}\left[{\overrightarrow{S}}_{l}\right]$ denote #$\#\mathrm{a}\left(J\right)$ and #$\#\mathrm{c}\left(J\right)$ output by [alg 1] respectively.
+Let $\mathrm{a}\left[{\overrightarrow{S}}_{l}\right]$ and $\mathrm{c}\left[{\overrightarrow{S}}_{l}\right]$ denote $\mathrm{a}\left(J\right)$ and $\mathrm{c}\left(J\right)$ output by [alg 1] respectively.
 Let $A\left[{\overrightarrow{S}}_{l}\right]$ and $r\left[{\overrightarrow{S}}_{l}\right]$ denote $A_{J}$ and $r_{J}$ output by [alg 1] respectively.
 Let
-[…]
-[…]
+${\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}:=\left\{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}\mathrm{:}\mathrm{[alg\ 1]\ terminates\ due\ to\ [t}\mathrm{iny}\mathrm{]}\right\}$
+${\overrightarrow{\mathcal{S}}}_{l\mathrm{,tang}}:=\left\{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}\mathrm{:}\mathrm{[alg\ 1]\ terminates\ due\ to\ [t}\mathrm{ang}\mathrm{]}\right\}$
 […]
@@
 Remark 3.2.
-By lemma 2.20, one has $m\geq k$. If $m<k$, then lemma 2.20 implies
+If $m<k$, then Lemma 2.21 implies
 ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lesssim\mathrm{RapDec}\left(R\right){\left\|f\right\|}_{2}$
-which contradicts (3.28).
+which contradicts (3.28). Thus, $m\geq k$.
 […]
 Verify the properties Property 2 follows from (3.25) and the repeated application of (3.20).
-Property 3 follows from (LABEL:r3.26_) and the repeated application of (3.21).
+Property 3 follows from (3.26) and the repeated application of (3.21).
 Property 1 follows from (3.19), (3.37), (3.24), (3.35), and Lemma 2.19.
@@
 $\frac{{\beta}_{i+1}-{\beta}_{i}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)\leq 0,\ \ \ \ \ \ \ \ \frac{{\beta}_{i+1}}{2}-\left(i+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)<0,\ \ \forall m\leq i\leq n-1$
-This gives the main result of Ou-Wang [4]. See [12] for details.
+This gives Theorem 1.3. See [11] for details.
 4 Proof of Theorem 2.23
@@
 Without loss of generality, we assume that $f$ satisfies (3.27).
-By applying [alg 2] to decompose the $k$-broad norm ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}$, we get the outputs described as Section 3.2.
-First, we improve (3.42) by using lemma 2.16, the nested polynomial Wolff axioms.
+Applying [alg 2] to decompose the $k$-broad norm ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}$, we get the outputs described in Section 3.2.
+First, we improve (3.42) by Corollary 2.17, the nested polynomial Wolff axioms.
 Lemma 4.1.
@@
 Proof.
-Since $f^{\#}_{{\overrightarrow{S}}_{l}}$ is concentrated on the $r_{n-1}$-scale wave packets $\left(\theta,v,l\right)$ with $\theta\in\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right]$, lemma 2.16 implies
+Since $f^{\#}_{{\overrightarrow{S}}_{l}}$ is concentrated on the $r_{n-1}$-scale wave packets $\left(\theta,v,l\right)$ with $\theta\in\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right$], Corollary 2.17 implies
 ${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lesssim\#\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right]r^{-\frac{n-2}{2}}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox\left(\prod^{n-2}_{j=m}{r^{-\frac{1}{2}}_{j}}\right){\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (4.1)
@@
 ${\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox D^{\delta}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq D^{\delta}_{n-1}{\left\|f\right\|}^{2}_{\infty}$ (4.2)
-Then lemma 4.1 follows from (4.1) and (4.2). ∎
-By combining (3.32), (3.33), and lemma 4.1, for $m\leq l<n-1$,
+Now Lemma (4.1) follows from (4.1) and (4.2). ∎
+Combining (3.32), (3.33), and Lemma 4.1, we see that for $m\leq l<n-1$,
 ${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox r^{-\frac{n-l}{2}}_{l}\left(\prod^{l-1}_{i=m-1}{r^{-\frac{1}{2}}_{i}D^{-i+\delta}_{i}}\right)r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-\frac{1}{2}}_{j}D^{\delta}_{j-1}}\left(\prod^{n-2}_{i=l}{r^{-\frac{1}{2}}_{i}}\right)D^{\delta}_{n-1}{\left\|f\right\|}^{2}_{\infty}$
-$=\left(\prod^{n-1}_{j=m\ }{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{l-1}_{j=m-1}{D^{-j}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$ (4.3)
+$\left(\prod^{n-1}_{j=m\ }{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{l-1}_{j=m-1}{D^{-j}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$ (4.3)
 Let ${\gamma}_{m},{\gamma}_{m+1},\dots,{\gamma}_{n}$ be positive real numbers satisfying ${\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n}=1$, and
 ${\sigma}_{j}:=\sum^{j}_{i=m}{{\gamma}_{i}},\ \ {\tau}_{j}:=\left\{\begin{array}[]{c}{\sigma}_{j}\mathrm{,\ \ if\ }j\leq n-2\\ 0,\ \ \mathrm{if\ }j=n-1\end{array}\right$.
-By taking the geometric mean of (3.41) and (4.3) with weights $\gamma_{i}$, we have
+Take the geometric mean of (3.41) and (4.3) with weights $\gamma_{i}$,
 ${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{j=m}{r^{-\frac{1+{\tau}_{j}}{2}}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{-j\left(1-{\sigma}_{j}\right)+\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$
@@
 ${\sigma}_{n-2}={\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n-2}<\frac{m-1}{2}\sum_{j\geq m}{\frac{1}{j\left(j-1\right)}}=\frac{1}{2}$ (4.10)
-Take $j=n-1$ in (4.6),
+Take j=n1j=n-1 in (4.6),
 $\left(n-1\right)\left(1-{\sigma}_{n-1}\right)-\left(n-2\right)\left(1-{\sigma}_{n-2}\right)=1/2\ \Longrightarrow{\gamma}_{n-1}=\frac{1/2-{\sigma}_{n-2}}{n-1}$

### v7

@@
-The Cone Restriction: An old Approach Revisited
-Abstract
-We consider Ou-Wang’s approach to cone restriction via polynomial partitioning.
-By restructuring their induction arguments into a recursive algorithm and applying the nested polynomial Wolff axioms, we refine the bounds on cone restriction estimate in higher dimensions.
-1 Introduction
-In this paper, we present a small improvement on cone restriction in higher dimensions.
-The Fourier restriction is one of the most significant open challenges in harmonic analysis, addresses the restriction of the Fourier transform to a compact manifold.
-Stein [16] conjectured that the restriction of Fourier transforms to a well-curved surface, such as sphere, paraboloids, or the cone studied in this paper, is a bounded $L^{p^{\prime}}\to L^{q^{\prime}}$ operator for some $p^{\prime}$ and $q^{\prime}$.
-The precise statement of cone restriction conjecture can be formulated as follows.
+On Cone Restriction Estimates in Higher Dimensions
+Abstract.
+We revisit the Ou-Wang’s approach to the cone restriction problem via polynomial partitioning.
+By recasting their inductive scheme as a recursive algorithm and incorporating the nested polynomial Wolff axioms, we obtain improved bounds for cone restriction estimates in higher dimensions.
+1. Introduction
+1.1. Statement of results
+In this paper, we obtain a modest improvement for the cone restriction problem in higher dimensions.
+The Fourier restriction problem is one of the central open questions in harmonic analysis.
+It concerns the possibility of restricting the Fourier transform of an $L^{p}$ function to a curved hypersurface.
+Stein [16] conjectured that, for a well-curved hypersurface such as the sphere, the paraboloid, or the cone considered in this paper, the Fourier transform admits a bounded restriction operator from $L^{p^{\prime}}$ to $L^{q^{\prime}}$ for an appropriate range of exponents $p^{\prime}$ and $q^{\prime}$.
+We now state the cone restriction conjecture precisely.
 Conjecture 1.1.
@@
 $\mathcal{C}:=\{(\xi_{1},\xi_{2},...,\xi_{n}):\xi_{1}^{2}+\xi_{1}^{3}+...+\xi_{n-1}^{2}=\xi_{n}^{2}\}$ (1.1)
-Then, for all $p,q>0$ satisfy
+Then for all $p,q$ satisfying
 $p>\frac{2(n-1)}{n-2},\frac{n-2}{q^{\prime}}\geq\frac{n}{p}$ (1.2)
-there holds
-$\|\hat{f}\|_{L^{q^{\prime}}(\mathcal{C};d\sigma)}\lesssim_{p,q}\|f\|_{p^{\prime}},\forall f:\mathbb{R}^{n}\to\mathbb{C}$
+we have
+$\|\hat{f}\|_{L^{q^{\prime}}(\mathcal{C};d\sigma)}\lesssim_{p,q}\|f\|_{L^{p^{\prime}}(\mathbb{R}^{n})},\forall f:\mathbb{R}^{n}\to\mathbb{C}$
 where $d\sigma$ denotes the Riemannian measure on $\mathcal{C}$, and $p^{\prime},q^{\prime}$ are the Hölder conjugates of $p,q$.
 Let $Ef$ be the associated Fourier extension operator on $\mathcal{C}$, i.e.,
-$Ef(x):=\int_{2\bar{B}^{n-1}\setminus B^{n-1}}e^{i(x_{1}\xi_{1}+...+x_{n-1}\xi_{n-1}|\xi|)}f(\xi)\,d\xi\$
-where $B^{n-1}$ is the unit ball centered at the origin. By a short duality argument, we see that conjecture 1.1 is equivalent to
+$Ef(x):=\int_{2\bar{B}^{n-1}\setminus B^{n-1}}e^{i(x_{1}\xi_{1}+...+x_{n-1}\xi_{n-1}+x_{n}|\xi|)}f(\xi)\,d\xi\$
+where $B^{n-1}\subset\mathbb{R}^{n-1}$ is the unit ball centered at the origin. By a short duality argument, we see that Conjecture 1.1 is equivalent to
 Conjecture 1.2.
-Let $n\geq 3$. For all $p,q>0$ satisfy (1.2),
-$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
-Conjecture 1.2 has so far been proven only in dimensions $n=3,4,5$ with contributions from Barcelo [1], Wolff [18], and Ou-Wang [15].
-Prior to this work, the best-known range for higher dimensions was established by Ou-Wang [15], with results as follows.
+Let $n\geq 3$. For all $p,q$ satisfying (1.2), the estimates
+$\|Ef\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{p,q}\|f\|_{L^{q}(\mathbb{R}^{n-1})}$ (1.3)
+holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
+Conjecture 1.2 has been confirmed only in the cases $n=3,4,5$ , through the work of Barceló [1], Wolff [18], and Ou–Wang [15].
+In higher dimensions, the best previously known range was obtained by Ou–Wang [15], whose results can be stated as follows.
 Theorem 1.3 ([15], Theorem 2).
-Let $n\geq 3$. For all $p,q>2$ satisfy
-$q\geq p>\left\{\begin{array}[]{rcl}4&\ \mathrm{if}&n=3\\ 2\frac{3n+1}{3n-3}&\mathrm{if}&n>3\quad\mathrm{odd}\\ 2\frac{3n}{3n-4}&\mathrm{if}&n>3\quad\mathrm{even}\end{array}\right$.
-there holds
-$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
-Theorem 1.4 ([15], Theorem 2).
-Let $n\geq 3$. Suppose the tuple $(p,q,k)$ is admissible in the sense that $2<q<p$, $2\leq k\leq n$ and
+Let $n\geq 3$. Suppose the tuple $(p,q,k)$ is admissible in the sense that $2\leq q\leq p$, $2\leq k\leq n$ and
 $\left\{\begin{array}[]{rcl}p>2\frac{n+2}{n},q^{\prime}\leq\frac{n-2}{n}p&\mathrm{if}&k=2\\ p>2\frac{n+k}{n+k-2},p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\mathrm{if}&k\geq 3\end{array}\right$.
-then
-$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
+then the estimate (1.3) holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
+This theorem yields the following corollary.
+Corollary 1.4 ([15], Theorem 1).
+Let $n\geq 3$. For all $p$ satisfying
+$p>\left\{\begin{array}[]{rcl}4&\ \mathrm{if}&n=3\\ 2\frac{3n+1}{3n-3}&\mathrm{if}&n>3\quad\mathrm{odd}\\ 2\frac{3n}{3n-4}&\mathrm{if}&n>3\quad\mathrm{even}\end{array}\right$. (1.4)
+the estimate
+$\|Ef\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{p}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$ (1.5)
+holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
+We now state our main result. Given integers $n\geq 3$ and $k$ with $2\leq k\leq n-2$, define $p(n,k)$ and $q(n,k)$ by
+$p(n,k):=2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}}$
+$\frac{1}{q(n,k)}=\frac{1}{p(n,k)}+\frac{1}{4n-8}\big(\frac{6}{p(n,k)}-1\big)$
+For $q\geq 2$, we define $p_{n,k}(q)$ by
+[…]
+where $\alpha_{n,k}(q)$ is defined by
+$\frac{1}{q}=\frac{\alpha_{n,k}(q)}{2}+\frac{1-\alpha_{n,k}(q)}{q(n,k)}$
+Theorem 1.5.
+Let $n\geq 3$. Suppose the tuple $(p,q,k)$ is admissible in the sense that $2\leq q\leq p$, $2\leq k\leq n-2$ and
+[…]
+then the estimate (1.3) holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
+Arguing as in [12], we obtain
+Corollary 1.6.
+For $n\geq 3$ and all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$, the estimate (1.5) holds whenever
+$p>2+\lambda n^{-1}+O(n^{-2}),\quad\lambda=2.596$… (1.6)
+Observe that (1.4) implies that the estimate (1.5) holds for all $f:\mathbb{R}^{n-1}\to\mathbb{C}$ whenever
+$p>2+\frac{8}{3n}+O(n^{-2}).$
+Consequently, Corollary 1.6 yields an improved $L^{p}\to L^{p}$ bound for the cone restriction conjecture in higher dimensions.
+Let $B_{R}\subset\mathbb{R}^{n}$ denote a ball of radius $R$.
+By a real interpolation of restricted-type, together with a standard $\epsilon$-removal argument (see [17, 4, 21]) and a bilinear interpolation (see [15]), Theorem 1.3 reduces to
+Proposition 1.7.
+Let $n\geq 3$. Suppose the tuple $(p,q,k)$ satisfies $2\leq q\leq p$, $2\leq k\leq n-2$, and
+$\begin{cases}p\geq p_{n,2}(q),q^{\prime}\leq\frac{n-2}{n}p&\text{if $k=2$},\\[4.0pt] p\geq p_{n,k}(q),p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\text{if $k\geq 3$}.\\[4.0pt] \end{cases},$
+then the estimate
+$\|Ef\|_{L^{p}(B_{R})}\lesssim_{p,q,\varepsilon}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{q}}\|f\|_{L^{\infty}(\mathbb{R}^{n-1})}^{1-\frac{2}{q}}$ (1.7)
+holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
 The Fourier restriction problem has been the focus of extensive study for decades.
@@
 However, sharp restriction estimates are known only for very few surfaces and dimensions.
-For instance, the restriction conjecture for the paraboloid and sphere remains unsolved for $n\geq 3$ It is known that the restriction conjecture for paraboloids or spheres in $\mathbb{R}^{n}$ implies the corresponding conjecture for cones in $\mathbb{R}^{n+1}$ (see [14]), which highlights the potential applications of our results.
-In this work, we show that:
-Theorem 1.5.
-For $n\geq 3$,
-$\|Ef\|_{L^{p}(B_{R})}\lesssim_{p,\epsilon}R^{\epsilon}\|f\|_{2}^{2/p}\|f\|_{\infty}^{1-2/p}$ (1.3)
-whenever
-$p>2+\lambda n^{-1}+O(n^{-2})$ (1.4)
-where $\lambda=2.596$… and $B_{R}$ is an arbitrary ball in $\mathbb{R}^{n}$ of radius $R$.
-By a real interpolation of restrict-type and a standard $\epsilon$-removal argument (see [17]), one has
-Corollary 1.6.
-Let $n\geq 3$. For all $q\geq p>2$ satisfy (1.4), there holds
-$\|Ef\|_{p}\lesssim_{p,q}\|f\|_{q},\quad\forall f:2\bar{B}^{n-1}\setminus B^{n-1}\to\mathbb{C}$
-Guth [5, 6] introduced polynomial partitioning to the study of Fourier restriction and established the improved results for paraboloids.
+For instance, the restriction conjecture for the paraboloid and sphere remains unsolved for $n\geq 3$.
+It is known that the restriction conjecture for paraboloids or spheres in $\mathbb{R}^{n}$ implies the corresponding conjecture for cones in $\mathbb{R}^{n+1}$ (see [14]), which highlights the potential applications of our results.
+Guth [5, 6] introduced polynomial partitioning to the study of Fourier restriction and established an improved results for paraboloids.
 Later, his results were subsequently refined by Wang[19], Wang-Wu[20], Hickman–Rogers [11], Hickman–Zahl [12], and Guo-Wang-Zhang [7].
-In [11, 12] the authors reformulated Guth’s induction-on-scale argument into a recursive algorithm, utilizing various Kakeya-type geometric results from [13, HRZ, 24], including the nested polynomial Wolff axiom, to bound the number of directions of the tubes those stay in varieties.
+In [11, 12] the authors reformulated Guth’s induction-on-scale argument into a recursive algorithm, utilizing various Kakeya-type geometric results from [13, 25], including the nested polynomial Wolff axiom, to bound the number of directions of the tubes those stay in varieties.
 Beyond paraboloid restriction, polynomial partitioning has also been applied to many related problems in harmonic analysis, such as Bochner-Riesz conjecture (see [22, 8]), the Schrödinger maximal estimate (see [3]), and cone restriction (see [15]).
@@
 Instead, we make a leaf backtrack to its (n1)(n-1)- dimensional ancestor.
-We will provide a more detailed discussion on this topic in Section 3.
+1.2. Structure of the article
 This article is not self-contained and heavily references the work of Guth [6], Ou-Wang [15], Hickman-Rogers [11], and Hickman-Zahl [12].
@@
 Finally, Section 4 discusses the numerology of Theorem 1.7.
-[…]
-Acknowledgements The author would like to thank Xiaochun Li for suggesting the problem and for numerous insightful discussions.
-The author is also grateful to Larry Guth and Shunhua Shizuo for their valuable guidance in learning decoupling theory.
-2 Preliminaries
-2.1 Notational Conventions
-In this paper, we work with Schwartz function $f$ or $g$ supported in $2\bar{B}^{n-1}\setminus B^{n-1}$.
+1.3. Notational Conventions
+In this paper, we work with Schwartz functions $f$ and $g$ supported in $2\overline{B}^{\,n-1}\setminus B^{n-1}$.
 We write $x\in\mathbb{R}^{n}$ as $x=(x^{\prime},x_{n})$.
-The parameters $n,k,\epsilon$ are fixed.
-All implicit constants are allowed to depend on $n,k,\epsilon$.
-We inductively define $\delta$ and $\delta_{j}$ for $0\leq j\leq n$ as bellows:
-$\delta_{0}:=\epsilon^{10},\quad\delta_{j}:=\delta_{j-1}^{10}\quad\mbox{for}\quad 1\leq j\leq n,\quad\delta:=\delta_{n}^{10}$
-In addition, we say $A\lesssim_{L}B$, $B\gtrsim_{L}A$, or $A=O_{L}(B)$ if $A\leq C_{L}B$ for some constant $C_{L}$ depending only on $L$.
-We say $A\sim_{L}B$ if both $A\lesssim_{L}B$ and $A\gtrsim_{L}B$ hold.
-We say $A\ll B$ or $B\gg A$ if $CA\leq B$ for some choice of $C\geq 1$ that can be taken to be as large as desired provided it is admissible.We say $A\lessapprox B$ if $A\lesssim R^{O(\delta_{0})}B$.
+The parameters $n$, $k$, and $\epsilon$ are fixed, and all implicit constants are allowed to depend on them.
+We inductively define $\delta$ and $\delta_{j}$ for $0\leq j\leq n$ by
+$\delta_{0}:=\epsilon^{10},\qquad\delta_{j}:=\delta_{j-1}^{10}\quad(1\leq j\leq n),\qquad\delta:=\delta_{n}^{10}.$
+We write $A\lesssim_{L}B$, $B\gtrsim_{L}A$, or $A=O_{L}(B)$ if $A\leq C_{L}B$ for some constant $C_{L}$ depending only on $L$.
+We write $A\sim_{L}B$ if both $A\lesssim_{L}B$ and $A\gtrsim_{L}B$ hold.
+We write $A\ll B$ or $B\gg A$ if $CA\leq B$ for some admissible constant $C\geq 1$ that may be taken arbitrarily large.
+We write $A\lessapprox B$ if $A\lesssim R^{O(\delta_{0})}B$.
 Let #A\#A denote the cardinality of a finite set $A$.
 Let $N_{r}E$ denote the $r$-neighbourhood of a set $E$.
-Let $B_{r}$ denote an arbitrary ball in $\mathbb{R}^{n}$ of radius $r$.
+Let $B_{r}\subset\mathbb{R}^{n}$ denote an arbitrary ball of radius $r$.
 Let $T_{x}Z$ denote the tangent space of $Z$ at $x$.
-We say $A^{\prime}\subseteq A$ is a refinement of $A$ if $A\lessapprox\#A^{\prime}$.
-We say $A$ = RapDec($r$) if $A\lesssim_{N}r^{-N}$.
-We say $f$ is essentially supported in $\Omega$ if $f(x)\lesssim$ RapDec($R$) $\|f\|_{2}$.
-We say $Ef$ is essentially supported in $\Omega$ if $Ef(x)\lesssim$ RapDec($R$) $\|f\|_{2}$.
-2.2 Wave Packet Decomposition
-The wave packet decomposition is a crucial tool in our argument. Here we recall its definition and basic properties from Sections 2.2 and 5.2.1 of [15].
-[…]
-Figure 1: A sector of a truncated cone.
-Given a function $f$ with supp $f\subseteq 2\bar{B}^{n-1}\setminus B^{n-1}$, one writes $f=\sum_{\theta\in\Theta[r]}f_{\theta}$ where $f_{\theta}=f\psi_{\theta}$ for $\psi_{\theta}$ a smooth partition of unity subordinate to the cover $\theta$.
-For each $\theta$, we break up $f_{\theta}$ as follows.
-First, we decompose $\mathbb{R}^{n-1}$ into finitely overlapping cubes of side length  $r^{1/2+\delta}$, centered at $v\in r^{1/2+\delta}Z^{n-1}$.
-Now, we can write $\hat{f_{\theta}}=\sum_{v}\hat{f_{\theta}}\eta_{v}$ , where $\eta_{v}$ is a smooth partition of unity subordinate to this cover.
-Second, for each $v$, we decompose $\hat{f_{\theta}}\eta_{v}$ into even finer pieces according to $\theta$.
-Cover the cube centered at $v$ by parallel tine plates $P_{\theta,v}^{l}$ with radius  $r^{1/2+\delta}$, thickness  $r^{\delta/2}$, and normal direction $\xi_{\theta}$.
-Here $\xi_{\theta}$ is the point on the central line of a sector $\theta$ with $|\xi_{\theta}|=1$.
-Hence, one can write
-$f=\sum_{\theta,v,l\in\mathbb{T}[r]}(\hat{f_{\theta}}\eta_{v}\eta_{\theta,v}^{l})^{\vee}$
-where $\eta_{\theta,v}^{l}$ is a smooth partition of unity subordinate to this cover, and $\mathbb{T}[r]$ denotes the set of wave packets of scale $r$.
-Since $(\hat{f_{\theta}}\eta_{v}\eta_{\theta,v}^{l})^{\vee}$ is essentially supported in $\theta$ (Strictly speaking, this should be $C\theta$ for some constant $C$.
-However, it is convenient to omit the constant,as it does not impact the main arguments.
-Therefore, we will continue to do so in the following discussion.), we can choose bump functions $\psi_{\theta}^{\prime}$ so that $\psi_{\theta}^{\prime}=1$ on a $r^{-1/2}$ neighborhood of supp $\psi_{\theta}$, and decays rapidly outside this region.
-Finally, we define
-$f_{\theta,v}^{l}:=\psi_{\theta}^{\prime}(\hat{f_{\theta}}\eta_{v}\eta_{\theta,v}^{l})^{\vee}$
-and then
-$f=\sum_{(\theta,v,l)\in\mathbb{T}[r]}f_{\theta,v}^{l}+\mbox{RapDec}(R)\|f\|_{2}$ (2.1)
+We say that $A^{\prime}\subseteq A$ is a refinement of $A$ if $A\lessapprox\#A^{\prime}$.
+We write $A=\mathrm{RapDec}(r)$ if $A\lesssim_{N}r^{-N}$.
+We say that a function $F$ is essentially supported in $\Omega$ if
+$F(x)\lesssim\mathrm{RapDec}(R)\,\|f\|_{2}\quad\text{for all }x\notin\Omega.$
+1.4. Acknowledgements
+The author would like to thank Xiaochun Li for suggesting the problem and for numerous insightful discussions.
+The author is also grateful to Larry Guth and Shunhua Shizuo for their valuable guidance in learning decoupling theory.
+1.5. Funding
+This research received no external funding.
+2. Preliminaries
+2.1. Wave Packet Decomposition
+Wave packet decomposition is a central tool in our argument. We briefly recall its definition and basic properties from Sections 2.2 and 5.2.1 of [15].
+Definition.
+Fix $R\gg 1$ and let $R^{\delta_{0}}\leq r\leq R$.
+Cover $2\overline{B}^{\,n-1}\setminus B^{n-1}$ by finitely overlapping $r^{-1/2}$-sectors $\theta\in\Theta[r]$.
+A sector $\theta\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ is called an $r^{-1/2}$-sector if it has unit length in the radial direction and angular width $r^{-1/2}$.
+We denote by $\Theta[r]$ the collection of all such finitely overlapping $r^{-1/2}$-sectors.
+Given a Schwartz function $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$, we write
+$f=\sum_{\theta\in\Theta[r]}f_{\theta},\qquad f_{\theta}:=f\,\psi_{\theta},$
+where $\{\psi_{\theta}\}$ is a smooth partition of unity subordinate to the cover $\{\theta\}$.
+For each $\theta$, we further decompose $f_{\theta}$ as follows.
+First, decompose $\mathbb{R}^{n-1}$ into finitely overlapping cubes of side length $r^{1/2+\delta}$ centered at $v\in r^{1/2+\delta}\mathbb{Z}^{n-1}$.
+Then
+$\widehat{f_{\theta}}=\sum_{v}\widehat{f_{\theta}}\,\eta_{v},$
+where $\{\eta_{v}\}$ is a smooth partition of unity subordinate to this cover.
+Next, for each $v$, we decompose $\widehat{f_{\theta}}\eta_{v}$ into finer pieces adapted to $\theta$.
+Cover the cube centered at $v$ by parallel thin plates $P_{\theta,v}^{l}$ of radius $r^{1/2+\delta}$, thickness $r^{\delta/2}$, and normal direction $\xi_{\theta}$, where $\xi_{\theta}$ denotes the unit vector pointing in the central direction of the sector $\theta$.
+Hence,
+$f=\sum_{(\theta,v,l)\in\mathbb{T}[r]}\big(\widehat{f_{\theta}}\,\eta_{v}\,\eta_{\theta,v}^{l}\big)^{\vee},$
+where $\{\eta_{\theta,v}^{l}\}$ is a smooth partition of unity subordinate to this cover, and $\mathbb{T}[r]$ denotes the collection of wave packets at scale $r$.
+Since $\big(\widehat{f_{\theta}}\,\eta_{v}\,\eta_{\theta,v}^{l}\big)^{\vee}$ is essentially supported in $\theta$ (more precisely in $C\theta$ for some constant $C$), we choose bump functions $\psi_{\theta}^{\prime}$ such that $\psi_{\theta}^{\prime}=1$ on an $r^{-1/2}$–neighbourhood of $\operatorname{supp}\psi_{\theta}$ and decays rapidly outside this region.
+We then define
+$f_{\theta,v}^{l}:=\psi_{\theta}^{\prime}\big(\widehat{f_{\theta}}\,\eta_{v}\,\eta_{\theta,v}^{l}\big)^{\vee}.$
+Consequently,
+$f=\sum_{(\theta,v,l)\in\mathbb{T}[r]}f_{\theta,v}^{l}\;+\;\mathrm{RapDec}(R)\,\|f\|_{2}.$ (2.1)
 For any $\mathcal{T}\subseteq\mathbb{T}[r]$, define
-$f_{\mathcal{T}}:=\sum_{(\theta,v,l)\in\mathcal{T}}f_{\theta,v}^{l}$
-We say $f$ is concentrated on wave packets from $\mathcal{T}\subseteq\mathbb{T}[r]$ if
-$\|f_{\mathbb{T}[r]\setminus\mathcal{T}}\|_{\infty}\lesssim\mbox{RapRec}(R)\|f\|_{2}$
-[…]
-$L^{2}$-OrthogonalityBy Plancherel theorem, the functions $f_{\theta,v}^{l}$ are almost orthogonal in $L^{2}$.
+$f_{\mathcal{T}}:=\sum_{(\theta,v,l)\in\mathcal{T}}f_{\theta,v}^{l}.$
+We say that $f$ is concentrated on the wave packets from $\mathcal{T}\subseteq\mathbb{T}[r]$ if
+$\|f_{\mathbb{T}[r]\setminus\mathcal{T}}\|_{\infty}\lesssim\mathrm{RapDec}(R)\,\|f\|_{2}.$
+$L^{2}$-Orthogonality. By Plancherel theorem, the functions $f_{\theta,v}^{l}$ are almost orthogonal in $L^{2}$.
 Lemma 2.1 ([15], Section 2.2).
-Let $r\geq 1$, then for any set $\mathcal{T}\subseteq\mathbb{T}[r]$,
+Suppose $r\geq 1$. For any $\mathcal{T}\subseteq\mathbb{T}[r]$, there holds
 $\|f_{\mathcal{T}}\|_{2}^{2}\sim\sum_{(\theta,v,l)\in\mathcal{T}}\|f_{\theta,v}^{l}\|_{2}^{2}$
-Like in the case of paraboloid [11], there are also some local versions of the $L^{2}$-orthogonality for the cone, which are direct corollaries of Lemma 2.1:
+As in the paraboloid case [11], there are corresponding local versions of the $L^{2}$-orthogonality for the cone, which follow directly from Lemma 2.1:
 Corollary 2.2.
-Let $1\leq\rho\leq r$, then for any set $\mathcal{T}\subseteq\mathbb{T}[r]$ and $\rho^{-1/2}$-sector $\theta^{*}$,
+Let $1\leq\rho\leq r$. For any set $\mathcal{T}\subseteq\mathbb{T}[r]$ and $\rho^{-1/2}$-sector $\theta^{*}$, we have
 $\|\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{f^{l}_{\theta,v}}\|^{2}_{L^{2}\left({\theta}^{*}\right)}\lesssim\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{{\left\|f^{l}_{\theta,v}\right\|}^{2}_{L^{2}\left(C{\theta}^{*}\right)}}$
@@
 Corollary 2.3.
-Let $1\leq\rho\leq r$, then for any set $\mathcal{T}\subseteq\mathbb{T}[r]$,
+Let $1\leq\rho\leq r$. For any set $\mathcal{T}\subseteq\mathbb{T}[r]$, we have
 ${\mathop{\mathrm{max}}_{{\theta}^{*}\mathrm{:}{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{f^{l}_{\theta,v}}\right\|}^{2}_{L^{2}\left({\theta}^{*}\right)}\ }\lesssim{\mathop{\mathrm{max}}_{{\theta}^{*}\mathrm{:}{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{{\left\|f^{l}_{\theta,v}\right\|}^{2}_{L^{2}\left(C{\theta}^{*}\right)}}\ }$
@@
 for some constant $C$.
-[…]
-Spatial concentration For a wave packet $\left(\theta,v,l\right)$, the restriction of $Ef^{l}_{\theta,v}$ to the ball $B\left(0,r\right)$ is essentially supported on a tube. More precisely, let
-$T^{l}_{\theta,v}:=N_{r^{\delta}}\left(P^{l}_{\theta,v}+rL\left(\theta\right)\right),\ \ rL\left(\theta\right):=\left\{tL\left(\theta\right):-r\leq t\leq r\right\}$
-be a $r$-tube with long direction $L\left(\theta\right)$ and mini direction $M\left(\theta\right)$, here
-$L\left(\theta\right):=\left(-{\xi}_{\theta},1\right),\ \ M\left(\theta\right):=\left({\xi}_{\theta},1\right)$
-Figure 2: The spatial structure of a tube.
-It is well-known that $Ef^{l}_{\theta,v}$ is essentially supported on $T^{l}_{\theta,v}$. Also, there is a standard $L^{2}$-estimate for $Ef$, see [6, 15] for details.
+Spatial Concentration. For a wave packet $(\theta,v,l)$, the restriction of $Ef_{\theta,v}^{l}$ to the ball $B(0,r)$ is essentially supported in a tube. More precisely, define
+$T_{\theta,v}^{l}:=N_{r^{\delta}}\!\left(P_{\theta,v}^{l}+rL(\theta)\right),\qquad rL(\theta):=\{\,t\,L(\theta):-r\leq t\leq r\,\},$
+which is an $r$-tube with long direction $L(\theta)$ and short direction $M(\theta)$, where
+$L(\theta):=(-\xi_{\theta},\,1),\qquad M(\theta):=(\xi_{\theta},\,1).$
+It is well known that $Ef_{\theta,v}^{l}$ is essentially supported in $T_{\theta,v}^{l}$.
+2.2. A Standard $L^{2}$ Estimate.
+Moreover, there is a standard $L^{2}$ estimate for EfEf; see [6, 15] for details.
 Lemma 2.4.
-For a function $f:2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$, we have
+Suppose $f$ is a Schwartz function and supported in $2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$, then
 ${\left\|Ef\right\|}_{L^{2}\left(B_{R}\right)}\lesssim R^{1/2}{\left\|f\right\|}_{2}$
-Also, if $R^{1/2+\delta}\leq r\leq R$ and $f$ is concentrated on wave packets $\left(\theta,v,l\right)$ with $T^{l}_{\theta,v}\cap B_{r}\neq\emptyset$, then
+Also, if $R^{1/2+\delta}\leq r\leq R$ and $f$ is concentrated on $R$-scale wave packets $\left(\theta,v,l\right)$ with $T^{l}_{\theta,v}\cap B_{r}\neq\emptyset$, then
 ${\left\|Ef\right\|}_{L^{2}\left(B_{10r}\right)}\sim r^{1/2}{\left\|f\right\|}_{2}$
-[…]
-Comparing wave packet decompositions at different scales Let $r^{1/2+2\delta}\leq\rho\leq r$ and $y\in B\left(0,r\right)$.
-Suppose $g$ is concentrated on the $r$-scale wave packet $\left(\theta,v,l\right)$ such that $T_{\theta,v,l}\cap B\left(y,\rho\right)\neq\emptyset$.
-To apply induction on scale, we need decompose a function into wave packet at different scales.
-We break up $g$ into wave packets over the ball $B\left(y,\rho\right)$ at scale $\rho$ as follows.
-First, we apply a transform to recenter $B\left(y,\rho\right)$ at the origin, i.e., $\tilde{g}\left(\xi\right):=e^{i{\psi}_{y}\left(\xi\right)}g\left(\xi\right)$, here
-${\psi}_{y}\left(\xi\right):=y_{1}{\xi}_{1}+\dots+y_{n-1}{\xi}_{n-1}+y_{n}\left|\xi\right|$
+2.3. Comparing Wave Packet Decompositions at Different Scales.
+Let $r^{1/2+2\delta}\leq\rho\leq r$ and let $y\in B(0,r)$.
+Suppose $g$ is concentrated on an $r$-scale wave packet $(\theta,v,l)$ with $T_{\theta,v,l}\cap B(y,\rho)\neq\emptyset$.
+To perform induction on scales, we decompose $g$ into wave packets at the smaller scale $\rho$ inside the ball $B(y,\rho)$.
+We begin by recentering $B(y,\rho)$ at the origin via the modulation
+$\tilde{g}(\xi):=e^{i\psi_{y}(\xi)}\,g(\xi),$
+where
+${\psi}_{y}\left(\xi\right):=y_{1}{\xi}_{1}+\dots+y_{n-1}{\xi}_{n-1}+y_{n}|\xi|$
 […]
@@
 Define ${\widetilde{\mathbb{T}}}_{\theta,v,l}$ as
-${\widetilde{\mathbb{T}}}_{\theta,v,l}:=\left\{\left(\zeta,w,L\right):\ \mathrm{D}\mathrm{ist}\left(\theta,\zeta\right)\lesssim{\rho}^{-1/2},\mathrm{D}\mathrm{ist}\left(P^{L}_{\zeta,w},P^{l}_{\theta,v}+P^{0}_{\theta}-{\partial}_{\xi}{\psi}_{y}\left({\xi}_{\theta}\right)\lesssim r^{\delta}\right)\right\}$
-here
+${\widetilde{\mathbb{T}}}_{\theta,v,l}:=\left\{\left(\zeta,w,L\right):\ \mathrm{Dist}\left(\theta,\zeta\right)\lesssim{\rho}^{-1/2},\mathrm{Dist}\left(P^{L}_{\zeta,w},P^{l}_{\theta,v}+P^{0}_{\theta}-{\partial}_{\xi}{\psi}_{y}\left({\xi}_{\theta}\right)\right)\lesssim r^{\delta}\right\}$
+where
 $\mathrm{D}\mathrm{ist}\left(A,B\right):={\mathop{\mathrm{max}}_{a\in A}{\mathop{\mathrm{min}}_{b\in B}\mathrm{dist}\left(a,b\right)\ }\ }\ \mathrm{for}\ A,B\subseteq{\mathbb{R}}^{n}$
-and $P^{0}_{\theta}$ is the tine plates centered at origin, with radius $\sim r^{1/2+\delta}$, thickness $\sim r^{\delta/2}$, and normal direction ${\xi}_{\theta}$.
+and $P^{0}_{\theta}$ denotes the thin plate centered at the origin, with radius $\sim r^{1/2+\delta}$, thickness $\sim r^{\delta/2}$, and normal direction $\xi_{\theta}$.
 Lemma 2.5 ([15], Lemma 5.4).
-For any $\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{\theta,v,l}$, there holds
+For any $\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{\theta,v,l}$, we have
 $\angle\left(L\left(\theta\right),L\left(\zeta\right)\right)\lesssim{\rho}^{-1/2},\angle\left(M\left(\theta\right),M\left(\zeta\right)\right)\lesssim{\rho}^{-1/2}$
 $\mathrm{D}\mathrm{ist}\left(T^{L}_{\zeta,w}\left(y\right),\ \left[T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right]+P^{0}_{\theta}\right)\lesssim r^{\delta}$
-here $\angle\left(v_{1},v_{2}\right)$denotes the angle between two non-zero vectors $v_{1},v_{2}\in{\mathbb{R}}^{n}$.
-In [15] the relationship between the two wave packet decompositions was studied in detail.
-For example, if $g$ is concentrated on wave packets from $\mathcal{T}\subseteq\mathbb{T}\left[r\right]$, then which scale $\rho$ wave packets will $\tilde{g}$ be concentrated on?
-We state the results from [15] below.
+where $\angle\left(v_{1},v_{2}\right)$denotes the angle between two non-zero vectors $v_{1},v_{2}\in{\mathbb{R}}^{n}$.
+The relationship between these two wave packet decompositions was studied in detail in [15].
+In particular, if $g$ is concentrated on wave packets from $\mathcal{T}\subseteq\mathbb{T}[r]$, one may ask which wave packets at scale $\rho$ the function $\tilde{g}$ is concentrated on.
+We recall the relevant results from [15] below.
 Lemma 2.6 ([15], Lemma 5.3).
@@
 $\widetilde{\mathcal{T}}=\bigcup_{\left(\theta,v,l\right)\in\mathcal{T}}{{\widetilde{\mathbb{T}}}_{\theta,v,l}}$
-[…]
+Roughly speaking, Lemma 2.6 asserts that $(g_{\theta,v}^{l})^{\widetilde{}}$ is concentrated on scale $\rho$ wave packets that are parallel to $T_{\theta,v}^{l}$ and contained in $c\,T_{\theta,v}^{l}$ for some constant $c$.
+2.4. Medium Tubes.
+Next, we group $\mathbb{T}[r]$ and $\mathbb{T}[\rho]$ into subcollections.
+For a $\rho^{-1/2}$-sector $\zeta_{0}$ and $v_{0}\in r^{1/2+2\delta}\mathbb{Z}^{n-1}\cap B(0,\rho)$, let $\mathbb{T}_{\zeta_{0},v_{0}}(y)$ denote the collection of $r$-tubes $T_{\theta,v}^{l}$ such that
 ${\mathrm{D}\mathrm{ist}}_{H}\left({\zeta}_{0},\theta\right)\lesssim{\rho}^{-1/2},\emptyset\neq T^{l}_{\theta,v}\cap B\left(y,\rho\right)\subseteq B\left(v_{0},r^{1/2+2\delta}\right)+\rho L\left({\zeta}_{0}\right)+\left\{y\right\}$
-here ${\mathrm{D}\mathrm{ist}}_{H}$ denotes the Hausdorff distance.
-For a $r$-tube $T^{l}_{\theta,v}$, if there are multiple $\left({\zeta}_{0},v_{0}\right)$ satisfying the above conditions, then we assign $T^{l}_{\theta,v}$ to only one ${\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)$.
-Also, we also define
+here $\mathrm{Dist}_{H}$ denotes the Hausdorff distance.
+If an $r$-tube $T_{\theta,v}^{l}$ satisfies the above conditions for multiple pairs $(\zeta_{0},v_{0})$, we assign $T_{\theta,v}^{l}$ to only one $\mathbb{T}_{\zeta_{0},v_{0}}(y)$.
+We also define
 $g_{{\zeta}_{0},v_{0}}:=\sum_{\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)}{g^{l}_{\theta,v}}$
 ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}:=\bigcup_{\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)}{{\widetilde{\mathbb{T}}}_{\theta,v,l}},\ \ {\tilde{g}}_{{\zeta}_{0},v_{0}}:=\sum_{\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}}{{\tilde{g}}^{L}_{\zeta,w}}$
-One immediately concludes that
-${\mathrm{D}\mathrm{ist}}_{H}\left({\zeta}_{0},\zeta\right)\lesssim{\rho}^{-\frac{1}{2}},P^{L}_{\zeta,w}\subseteq B\left(v_{0},r^{\frac{1}{2}+2\delta}\right),\ \ \forall\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}},\quad g=\sum_{{\zeta}_{0},v_{0}}{g_{{\zeta}_{0},v_{0}}}+\mbox{RapDec}\left(r\right){\left\|g\right\|}_{2}$ (2.2)
-Since each $\left(\zeta,w,L\right)\$belongs to $\sim 1$ ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, the Lemma 2.1 implies
+We immediately conclude that for all $(\zeta,w,L)\in{\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$
+${\mathrm{Dist}}_{H}({\zeta}_{0},\zeta)\lesssim{\rho}^{-\frac{1}{2}},P^{L}_{\zeta,w}\subseteq B(v_{0},r^{\frac{1}{2}+2\delta})$
+and
+$g=\sum_{{\zeta}_{0},v_{0}}{g_{{\zeta}_{0},v_{0}}}+\mbox{RapDec}\left(r\right){\left\|g\right\|}_{2}$
+Since each $\left(\zeta,w,L\right)\$belongs to $\sim 1$ ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, Lemma 2.1 implies
 ${\left\|g\right\|}^{2}_{2}\sim\sum_{{\zeta}_{0},v_{0}}{{\left\|g_{{\zeta}_{0},v_{0}}\right\|}^{2}_{2}},\ \ {\left\|\tilde{g}\right\|}^{2}_{2}\sim\sum_{{\zeta}_{0},v_{0}}{{\left\|{\tilde{g}}_{{\zeta}_{0},v_{0}}\right\|}^{2}_{2}}$ (2.2)
-From Lemma 2.5 and 2.6, one has
+It follows from Lemmas 2.5 and 2.6 that
 Corollary 2.7 ([15], Lemma 5.6).
@@
 $\bigcup_{\begin{array}[]{c}{\mathrm{D}\mathrm{ist}}_{H}\left({\zeta}_{0},\zeta\right)\lesssim{\rho}^{-1/2}\\ {\mathrm{D}\mathrm{ist}}_{H}\left(v_{0},v\right)\lesssim r^{1/2+2\delta}\end{array}}{{\mathbb{T}}_{\zeta,v}\left(y\right)}$
-[…]
-Tangential and transverse wave packets Here we recall some definitions and properties about the tangential and transverse wave packets from [6, 15].
+2.5. Tangential and Transverse Wave Packets.
+We recall several definitions and properties of tangential and transverse wave packets from [6, 15].
 Definition 2.8 (Transverse complete intersection, [6], Section 5.1).
@@
 The degree of the transverse complete intersection is defined as $\mathop{\mathrm{max}}_{1\leq j\leq n-m}{\mathrm{deg}P_{j}\ }\$}.
-From the above definition, if $\mathrm{Z}:=Z\left(P_{1},\dots,P_{n-m}\right)$ is a transverse complete intersection, then $\mathrm{Z}$ is a smooth manifold of dimension $m$, in other words, $\mathrm{Z}$ has a tangent plane $T_{x}\mathrm{Z}$ at every point $x\in\mathrm{Z}$.
+From the above definition, if $\mathrm{Z}:=Z(P_{1},\dots,P_{n-m})$ is a transverse complete intersection, then $\mathrm{Z}$ is a smooth manifold of dimension $m$.
+In particular, $\mathrm{Z}$ admits a tangent space $T_{x}\mathrm{Z}$ at every point $x\in\mathrm{Z}$.
 Definition 2.9 (Tangential and transverse tubes, [6], Section 5.2).
-Let $\boldsymbol{\mathrm{Z}}$ be a transverse complete intersection of dimension $m$.
-For a fixed ball $B\left(y,r\right)$, a $r$-tube $T^{l}_{\theta,v}$ with $T^{l}_{\theta,v}\cap B\left(y,r\right)\cap\boldsymbol{\mathrm{Z}}\neq\emptyset$ is said to be tangent to $\mathrm{Z}$ in $B\left(y,r\right)$ if $T^{l}_{\theta,v}\cap B\left(y,2r\right)\subseteq N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{\mathrm{Z}}\right)$ and
-$\angle\left(L\left(\theta\right),T_{x}\boldsymbol{\mathrm{Z}}\right)\lesssim r^{-1/2+{\delta}_{m}},\ \ \forall x\in\mathrm{Z}\cap B\left(y,r\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(T^{l}_{\theta,v})\right)$
-Otherwise, we say that $T^{l}_{\theta,v}$ is transverse to $\boldsymbol{\mathrm{Z}}$ in $B\left(y,r\right)$.
-We use $\mathbb{T}\left[\boldsymbol{\mathrm{Z}},B\left(y,r\right)\right]$ to denote the $r$-tube tangent to $\boldsymbol{\mathrm{Z}}$ in $B\left(y,r\right)$.
-In Section 3 we need the following result, which can be roughly taken as a continuum version of the Bezout’s theorem, to controls the transverse intersections between a tube and an algebraic variety,
+Let $\bm{\mathrm{Z}}$ be a transverse complete intersection of dimension $m$.
+For a fixed ball $B\left(y,r\right)$, a $r$-tube $T^{l}_{\theta,v}$ with $T^{l}_{\theta,v}\cap B\left(y,r\right)\cap\bm{\mathrm{Z}}\neq\emptyset$ is said to be tangent to $\mathrm{Z}$ in $B\left(y,r\right)$ if $T^{l}_{\theta,v}\cap B\left(y,2r\right)\subseteq N_{r^{1/2+{\delta}_{m}}}\left(\bm{\mathrm{Z}}\right)$ and
+$\angle\left(L\left(\theta\right),T_{x}\bm{\mathrm{Z}}\right)\lesssim r^{-1/2+{\delta}_{m}},\ \ \forall x\in\mathrm{Z}\cap B\left(y,r\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(T^{l}_{\theta,v})\right)$
+Otherwise, we say that $T^{l}_{\theta,v}$ is transverse to $\bm{\mathrm{Z}}$ in $B\left(y,r\right)$.
+We use $\mathbb{T}\left[\bm{\mathrm{Z}},B\left(y,r\right)\right]$ to denote the set of $r$-tubes tangent to $\bm{\mathrm{Z}}$ in $B\left(y,r\right)$.
+In Section 3, we require the following result, which can be viewed as a continuum analog of Bézout’s theorem, to control the transverse intersections between a tube and an algebraic variety.
 Lemma 2.10 ([6], Lemma 5.7).
@@
 For any $\alpha>0$, define
-$Z_{>\alpha}:=\left\{x\in Z:\angle\left(l,T_{x}\boldsymbol{\mathrm{Z}}\right)>\alpha\right\}$
+$Z_{>\alpha}:=\left\{x\in Z:\angle\left(l,T_{x}\bm{\mathrm{Z}}\right)>\alpha\right\}$
 Then $Z_{>\alpha}\cap T$ is contained in a union of $\lesssim D^{n}$ balls of radius $\lesssim r{\alpha}^{-1}$.
-2.3 Some Kakeya-type geometric results
-First, we introduce some definitions about grains and multigrain from Section 3 of [12].
+2.6. Reduction to Broad Estimates
+In [6], Guth decomposed $\|Ef\|_{L^{p}(B_{R})}$ into a broad part and a narrow part.
+The narrow part is essentially supported in a lower-dimensional subspace of $\mathbb{R}^{n}$.
+Using decoupling [2] together with induction on scales, he reduced the linear restriction estimates to estimating the broad part.
+Ou–Wang [15] followed the same strategy.
+Fix a large constant $K$.
+Decompose $2\overline{B}^{\,n-1}\setminus B^{n-1}$ in frequency space into $K^{-1}$-sectors $\tau$.
+Given a Schwartz function $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$, write
+$f=\sum_{\tau}f_{\tau},\qquad f_{\tau}:=f\,\psi_{\tau},$
+where $\{\psi_{\tau}\}_{\tau}$ is a smooth partition of unity subordinate to the cover $\{\tau\}$.
+Cover $B_{R}$ by finitely overlapping $B_{K^{2}}$.
+Fix a parameter $1<A\lesssim K^{\epsilon}$.
+For each $B_{K^{2}}$, define $\mu_{Ef}(B_{K^{2}})$ by
+$\mu_{Ef}(B_{K^{2}}):={\mathop{\mathrm{min}}_{V_{1},\dots,\ V_{A}\ \left(k-1\right)\mathrm{-}\mathrm{subspace\ of}\ {\mathbb{R}}^{n}}\left(\mathop{\mathrm{max}}_{\tau:\angle\left(L\left(\tau\right),V_{a}\right)\gtrsim K^{-2},\forall 1\leq a\leq A}\int_{B_{K^{2}}}{{\left|Ef_{\tau}\right|}^{p}}\right)\ }$
+where
+$\angle\left(L\left(\tau\right),V_{a}\right):={\mathop{\mathrm{min}}_{v\in L\left(\tau\right)}\angle\left(v,V_{a}\right)\ }$
+For $U\subseteq{\mathbb{R}}^{n}$ the $k$-broad norm over $U$ is defined by
+${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}:={\left(\sum_{B_{K^{2}}}{\frac{\left|B_{K^{2}}\cap U\right|}{\left|B_{K^{2}}\right|}}{\mu}_{Ef}\left(B_{K^{2}}\right)\right)}^{\frac{1}{p}}$
+Here we recall some basic properties about broad norm.
+Lemma 2.11 (Triangle inequality, [15], Lemma 3.1).
+Suppose that $1\leq p<\infty$, $f=g+h$, and $A=A_{1}+A_{2}$, where $A,A_{1},A_{2}$ are non-negative integers. Then
+${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}\lesssim{\left\|Eg\right\|}_{BL^{p}_{k,A_{1}}\left(U\right)}+{\left\|Eh\right\|}_{BL^{p}_{k,A_{2}}\left(U\right)}$
+Lemma 2.12 (Hölder’s inequality, [15], Lemma 3.2).
+Suppose $1\leq p,p_{1},p_{2}<\infty$, and $0\leq{\alpha}_{1},{\alpha}_{2}\leq 1$ obey ${\alpha}_{1}+{\alpha}_{2}=1$ and
+$\frac{1}{p}=\frac{{\alpha}_{1}}{p_{1}}+\frac{{\alpha}_{2}}{p_{2}}$
+Suppose that and $A=A_{1}+A_{2}$, then
+${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}\leq{\left\|Ef\right\|}^{{\alpha}_{1}}_{BL^{p_{1}}_{k,A_{1}}\left(U\right)}{\left\|Ef\right\|}^{{\alpha}_{2}}_{BL^{p_{2}}_{k,A_{2}}\left(U\right)}$
+Lemma 2.13 (Standard $L^{2}$-estimate, [15], Section 5).
+For $f\in L^{2}\left(2{\overline{B}}^{n-1}\backslash B^{n-1}\right)$ and non-negative integers $A$, there holds
+${\left\|Ef\right\|}^{2}_{BL^{2}_{k,A}\left(B_{r}\right)}\lesssim r{\left\|f\right\|}^{2}_{2}$
+Lemma 2.14 (Vanishing property, [15], Lemma 4.5).
+Suppose $Ef$ is tangent to a transverse complete intersection $\bm{Z}$ in $B_{r}$. If $\mathrm{deg}\bm{Z}\lesssim 1\$} and $dim\bm{Z}\$} $\leq k-1$, then
+${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{r}\right)}\lesssim\mathrm{RapDec}\left(r\right){\left\|f\right\|}_{2}$
+Broad estimates may imply linear estimates. More precisely,
+Theorem 2.15 ([15], Section 6.2).
+Let $n\geq 3$ and $2\leq q\leq p$. Suppose for some integer $A>0$ the estimate
+$\|Ef\|_{BL_{k,A}^{p}(B_{R})}\lesssim_{p,q,\varepsilon}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{q}}\|f\|_{L^{\infty}(\mathbb{R}^{n-1})}^{1-\frac{2}{q}}$ (2.3)
+holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$. If there exists an integer $k$ with $2\leq k\leq n$ and
+$\left\{\begin{array}[]{rcl}q^{\prime}\leq\frac{n-2}{n}p&\mathrm{if}&k=2\\ p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\mathrm{if}&k\geq 3\end{array}\right$.
+then the estimate (1.7) holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
+Hence, together with Lemma 2.12, Proposition 1.7 reduces to
+Proposition 2.16.
+Suppose $n\geq 3$ and $2\leq k\leq n-2$.
+For $p=p(n,k)$, $q=q(n,k)$ and some integer $A\lesssim_{\varepsilon,n}1$, the estimate (2.3) holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
+2.7. Transverse Equidistribution Estimates
+Using a version of the Heisenberg uncertainty principle, Guth [6] established transverse equidistribution estimates for the paraboloid.
+However, in the cone setting the tubes are much thinner, and $Eg$ may be concentrated on an extremely thin layer of $N_{R^{1/2+\delta_{m}}}(\mathbf{Z})$, so this property may fail to hold.
+To overcome this geometric obstruction, [15] decomposes $g$ into an essential part $g_{\mathrm{ess}}$ and a tail part $g_{\mathrm{tail}}$.
+Roughly speaking, $g_{\mathrm{ess}}$ is concentrated on wave packets $(\theta,v,l)$ for which $M(\theta)$ is not perpendicular to the tangent space of $\mathbf{Z}$ at some point $z\in\mathbf{Z}\cap N_{R^{1/2+\delta_{m}}}(T_{\theta,v}^{l})$; consequently, transverse equidistribution estimates hold for $g_{\mathrm{ess}}$.
+The tail term $g_{\mathrm{tail}}$ can be made negligible by choosing $A$ sufficiently large.
+Let $\bm{\mathrm{Z}}$ be an $m$-dimensional transverse complete intersection with degree $\leq d$.
+Let $B_{\rho}$ be a $\rho$-ball with $\rho\gg r^{1/2+{\delta}_{m}}$ and $B_{\rho}\cap N_{r^{1/2+{\delta}_{m}}}\left(\bm{\mathrm{Z}}\right)\cap B\left(0,r\right)\neq\emptyset$.
+Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\bm{\mathrm{Z}},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B_{\rho}}$, where
+${\mathbb{T}}_{B_{\rho}}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[r\right],T^{l}_{\theta,v}\cap B_{\rho}\neq\emptyset\right\}$
+For a vector space $V$ spanned by ${\left\{v_{i}=\left(x_{i,1},x_{i,2},\dots,-x_{i,n}\right)\right\}}^{m}_{i=1}$, we define $V^{+}$ by
+$V^{+}:=\left\{\sum^{m}_{i=1}{a_{i}v^{+}_{i}}:a_{i}\in\mathbb{R}\mathrm{,}v^{+}_{i}=\left(x_{i,1},x_{i,2},\dots,x_{i,n}\right),\ \mathrm{f}\mathrm{or\ }1\leq i\leq m\right\}$
+We cover $B_{\rho}$ by finitely overlapping balls $B$ of radius $r^{1/2+\delta_{m}}$.
+For a suitable constant $c>0$, if $cB\cap B_{\rho}\cap\mathbf{Z}\neq\emptyset$, we choose a point $x\in cB\cap B_{\rho}\cap\mathbf{Z}$ and denote $V_{B}:=T_{x}\mathbf{Z}$.
+Let
+${\mathcal{B}}_{a}:=\left\{B:\mathrm{\ }\angle\left(V^{+}_{B},V^{\bot}_{B}\right)\gtrsim K^{-2}\right\}$
+${\mathcal{B}}_{b}:=\left\{B:\mathrm{\ }\angle\left(V^{+}_{B},V^{\bot}_{B}\right)\lesssim K^{-2}\right\}$
+${\mathbb{T}}_{\bm{\mathrm{Z}},B}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[\bm{\mathrm{Z}},B\left(0,r\right)\right]:T^{l}_{\theta,v}\cap B\neq\emptyset\right\}$
+From the above definition, if $B\in\mathcal{B}_{b}$, then $\angle\!\big(M(\theta),V_{B}^{\perp}\big)$ may be small for some $(\theta,v,l)\in\mathbb{T}_{\mathbf{Z},B}$, and consequently the transverse equidistribution estimates may fail for functions concentrated on $\mathbb{T}_{\mathbf{Z},B}$.
+Fortunately, there are only few sectors $\theta$ for which $(\theta,v,l)\in\mathbb{T}_{\mathbf{Z},B}$ for some $v,l$.
+Lemma 2.17 ([15], Lemma 5.8).
+Suppose $B\in{\mathcal{B}}_{b}$. Let
+$\mathbb{T}\left[{\theta}^{\prime}\right]:=\left\{\left(\theta,v,l\right):\theta={\theta}^{\prime}\right\},\ \ {\mathrm{\Theta}}_{B}:=\left\{\theta:\mathbb{T}\left[\theta\right]\cap{\mathbb{T}}_{\bm{\mathrm{Z}},B}\neq\emptyset\right\}$
+Then ${\mathrm{\Theta}}_{B}$ is contained in the union of $O\left(1\right)$ many $K^{-2}$-sectors.
+Now, we recall the definition of the essential part $g_{\mathrm{ess}}$ and the tail part $g_{\mathrm{tail}}$.
+${\mathbb{T}}_{\mathrm{ess}}:=\left\{\left({\zeta}_{0},v_{0}\right):\exists\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left.y\right.\mathrm{\ s.t.\ }T^{l}_{\theta,v}\cap B\neq\emptyset\mathrm{\ }\mathrm{fo}\mathrm{r\ some\ }B\in{\mathcal{B}}_{a}\right\}$
+${\mathbb{T}}_{\mathrm{tail}}:=\left\{\left({\zeta}_{0},v_{0}\right):\forall\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left.y\right.\mathrm{,\ }T^{l}_{\theta,v}\cap B=\emptyset\mathrm{\ }\mathrm{fo}\mathrm{r\ all\ }B\in{\mathcal{B}}_{a}\right\}$
+$g_{\mathrm{ess}}:=\sum_{\left({\zeta}_{0},v_{0}\right)\in{\mathbb{T}}_{\mathrm{ess}}}{g_{{\zeta}_{0},v_{0}}},\ \ g_{\mathrm{tail}}:=\sum_{\left({\zeta}_{0},v_{0}\right)\in{\mathbb{T}}_{\mathrm{tail}}}{g_{{\zeta}_{0},v_{0}}}$ (2.4)
+Suppose $A$ is chosen to be sufficiently large. By the definition of $k$-broad norm and Lemma 2.17,
+${\left\|Eg_{\mathrm{tail}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}=\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$
+Combining this with Lemma 2.11, we have
+${\left\|Eg\right\|}_{BL^{p}_{k,A}\left(B_{\rho}\right)}\lesssim{\left\|Eg_{\mathrm{ess}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}+\mathrm{RapDec}\left(R\right){\left\|g\right\|}_{2}$ (2.5)
+As shown in [15], the transverse equidistribution estimates hold for the essential part $g_{\mathrm{ess}}$.
+The precise statement of the transverse equidistribution estimates for the cone is as follows:
+Lemma 2.18 ([15], Lemma 5.13).
+Let $B_{\rho}$ be a $\rho$-ball with $\rho\gg r^{1/2+{\delta}_{m}}$ and $B_{\rho}\cap N_{r^{1/2+{\delta}_{m}}}\left(\bm{Z}\right)\cap B\left(0,r\right)\neq\emptyset$.
+Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\bm{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B_{\rho}}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$  then
+${\left\|g_{\mathrm{ess,}b\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g\right\|}^{2}_{2}$
+where $g_{\mathrm{ess,}b\mathrm{\ }}:={\left(g_{\mathrm{ess\ }}\right)}_{b}$.
+To apply the nested polynomial Wolff axioms, we need to generalize Lemma 2.18.
+Lemma 2.19.
+Let $\rho\gg r^{1/2+{\delta}_{m}}$ and $B\left(y,\rho\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(\bm{Z}\right)\cap B\left(0,r\right)\neq\emptyset$.
+Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\bm{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B\left(y,\rho\right)}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$.
+For $\widetilde{\mathcal{T}}\subseteq\mathbb{T}\left[{\bm{\mathrm{Z}}}_{b}-y,B\left(0,\rho\right)\right]$, there holds
+${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$
+where $\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}$ denotes the set of $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ for which there exists some $\left(\zeta,w,L\right)\in\widetilde{\mathcal{T}}$ satisfying
+${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq C{\rho}^{-1/2},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y\right),\ T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right)\leq Cr^{1/2+\delta}$
+for some constant $C$, and ${\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}:={{\left(g_{\mathrm{ess}}\right)}}_{\widetilde{\mathcal{T}}}^{\widetilde{}}$.
+Proof.
+Let $h=g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}$. By Lemma 2.18,
+${\left\|{\tilde{h}}_{\mathrm{ess},b}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}$ (2.6)
+By Lemma 2.1 and 2.6,
+${\left\|{\tilde{h}}_{\mathrm{ess},b}\right\|}^{2}_{2}\gtrsim{\left\|{\left(h_{\mathrm{ess,\uparrow\tilde{\mathcal{T}}}}\right)}^{\widetilde{}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.7)
+here $\mathrm{\uparrow}\widetilde{\mathcal{T}}$ denotes the set of $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ for which there exists some $\left(\zeta,w,L\right)\in\widetilde{\mathcal{T}}$ satisfying
+${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq c{\rho}^{-1/2},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y\right),\ T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right)\leq cr^{1/2+\delta}$
+for some large constant $c$. Also, Lemma 2.6 gives
+${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}={\left\|{\left(h_{\mathrm{ess,\uparrow\tilde{\mathcal{T}}}}\right)}^{\widetilde{}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.8)
+Combining (2.6), (2.7), and (2.8), we obtain
+${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{ess,}\ \mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$
+2.8. Polynomial Partitioning
+Polynomial partitioning plays a central role in broad estimates.
+In this subsection, we review the theory of polynomial partitioning.
+This method is the primary tool used by Guth [5, 6] and Ou–Wang [15] to improve restriction estimates, and it has also led to progress on several related problems, including the Bochner–Riesz conjecture [22] and the Schrödinger maximal estimate [3].
+For a polynomial $P:\mathbb{R}^{n}\to\mathbb{R}$, let $\mathrm{cell}(P)$ denote the collection of connected components of $B_{r}\setminus Z(P)$, which partitions the space into finitely many disjoint regions.
+Each $O\in\mathrm{cell}(P)$ is called a cell cut out by $P$.
+Theorem 2.20 ([6], Section 8).
+Fix $r\gg 1$, and $d\in{\mathbb{N}}^{+}$.
+Suppose $F\in L^{1}\left({\mathbb{R}}^{n}\right)$ is supported on $B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\bm{\mathrm{Z}}$, where $\bm{\mathrm{Z}}$ is an $m$-dimensional transverse complete intersection of degree at most $d$.
+Then, at least one of the following cases holds:
+Cellular case. There exists a polynomial $P:{\mathbb{R}}^{n}\to\mathbb{R}$ of degree $O\left(d\right)$ with the following properties:
+[…]
+(2).For each $O\in\mathcal{O}$, ${\left\|F\right\|}_{L^{1}\left(O\right)}\sim d^{-m}{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
+where $\mathcal{O}$ is a refinement of the set of connected components of $B_{r}\backslash N_{r^{1/2+{\delta}_{m}}}Z\left(P\right)$.
+Algebraic case. There exists a $\left(m-1\right)$-dimensional transverse complete intersection $\bm{\mathrm{Y}}\subseteq\bm{\mathrm{Z}}$ of degree $O\left(d\right)$ such that
+${\left\|F\right\|}_{L^{1}\left(B_{r}\right)}\lesssim{\left\|F\right\|}_{L^{1}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\bm{\mathrm{Y}}\right)}$
+We now apply polynomial partitioning to the broad estimates. If the cellular case holds, then
+$\|Ef\|^{p}_{BL^{p}_{k,A}\!\left(B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Z}\right)}\lesssim d^{m}\|Ef\|^{p}_{BL^{p}_{k,A}(O)}\quad\text{for all }O\in\mathcal{O},$
+where $\mathcal{O}$ denotes the collection of cells produced by Theorem 2.20. If the algebraic case holds, then
+$\|Ef\|^{p}_{BL^{p}_{k,A}\!\left(B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Z}\right)}\lesssim\|Ef\|^{p}_{BL^{p}_{k,A}\!\left(B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Y}\right)},$
+where $\mathbf{Y}$ is an (m1)(m-1)-dimensional transverse complete intersection produced by Theorem 2.20.
+By the fundamental theorem of algebra, we deduce that
+Lemma 2.21.
+For any $r$-tube $T^{l}_{\theta,v}$,
+$\left\{O\in{\mathcal{O}}^{\prime}:O\cap T^{l}_{\theta,v}\neq\emptyset\right\}\lesssim d$
+2.9. Nested Polynomial Wolff Axiom.
+To refine the work of Ou–Wang [15], we require the nested polynomial Wolff axioms. We begin by recalling the definitions of grains and multigrains from Section 3 of [12].
 Definition 2.22 ([12], Definition 3.3).
-A grain is defined to be a pair $\left(Z,B\left(y,r\right)\right)$ where $Z$ is a transverse complete intersection and $B_{r}$ is a ball of some radius $r>0$.
+A grain is defined to be a pair $\left(Z,B\left(y,r\right)\right)$ where $Z$ is a transverse complete intersection and $B_{r}$ is a ball of radius $r>0$.
 The dimension of a grain $\left(Z,B\left(y,r\right)\right)$ is that of the variety $Z$, while its degree is that of $Z$ and its scale is $r$.
@@
 Let $S=\left(Z,B\left(y,r\right)\right)$ be a grain of dimension $m$.
-A function $f$ with $\mathrm{supp}\ f\subseteq 2{\overline{B}}^{n-1}\backslash B^{n-1}$ is said to be tangent to $S$ if it is concentrated on scale $r$ wave packets belong to the collection:
+A function $f$ with $\mathrm{supp}\ f\subseteq 2{\overline{B}}^{n-1}\backslash B^{n-1}$ is said to be tangent to $S$ if it is concentrated on scale $r$ wave packets belonging to the collection:
 $\mathbb{T}\left[S\right]:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]:T^{l}_{\theta,v}\left(y\right)\ \mathrm{is}\ \mathrm{tangent\ to}\ Z\ \mathrm{in}\ B\left(y,r\right)\right\}$
@@
 $0<r_{s}\leq r_{s+1}\leq\dots\leq r_{t}$
-The parameter $\left(t-s+1\right)$ is called the multigrain level.
-The degree of $\overrightarrow{S}$ is defined as $\mathop{\mathrm{max}}_{1\leq i\leq l}Z_{i}\$}.
-The multiscale of $\overrightarrow{S}$ is the tuple $\overrightarrow{r}=\left(r_{s},r_{s+1},\dots,r_{t}\right)$.
+The parameter (ts+1)(t-s+1) is called the multigrain level.
+The degree of $\overrightarrow{S}$ is defined as $\max_{1\leq i\leq l}Z_{i}$.
+The multiscale of $\overrightarrow{S}$ is the tuple $\overrightarrow{r}=(r_{s},r_{s+1},\dots,r_{t})$.
 Finally, we write
 ${\overrightarrow{S}}_{1}=\left(S_{s},S_{s+1},\dots,S_{t}\right)\preccurlyeq{\overrightarrow{S}}_{2}=\left(S_{s^{\prime}},S_{s^{\prime}+1},\dots,S_{t}\right),\ \ \mathrm{if\ }s^{\prime}\geq s$
-The restriction conjecture is associated with a Kakeya-type problem concerning the continuum incidence theory of tubes.
-Specifically, given a multigrain ${\overrightarrow{S}}_{s,t}=\left(S_{s},\dots,S_{t}\right)$ with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$, if $\mathcal{L}$ is a set of $r^{-1/2}_{t}$-direction separated rays that satisfies
-$L\cap B\left(y_{i},r_{i}\right)\cap N_{{\rho}_{j}}Z_{i}\neq\emptyset$
-$L\cap B\left(y_{i},2r_{i}\right)\subseteq N_{{\rho}_{j}}Z_{i}\cap B\left(y_{i},2r_{i}\right)$
-for any $L\in\mathcal{L}$, then what is the maximum possible value of $\mathcal{L}$? Here $\mathcal{L}$ is said to be $r^{-1/2}$-direction separated if
-$\angle\left(L_{1},L_{2}\right)\gtrsim r^{-\frac{1}{2}},\ \ \mathrm{if}\ L_{1},L_{2}\in\mathcal{L}\boldsymbol{\ }\mathrm{and}\boldsymbol{\ }L_{1}\neq L_{2}$
-[…]
-Theorem 2.14 (Polynomial Wolff Axiom).
-Let $r\geq 1$. Let $Z\subset{\mathbb{R}}^{n}$ be an algebraic variety of dimension $m$ and degree at most $d$. If $\mathcal{T}\subseteq\mathbb{T}\left[r\right]$satisfies: $\forall T\in\mathcal{T}$,
-$T\cap B_{r}\cap N_{r^{1/2+{\delta}_{m}}}Z\neq\emptyset\$
-$T\cap 2B_{r}\subseteq N_{r^{1/2+{\delta}_{m}}}Z$
-then
-$\mathcal{T}{\lesssim}_{d}r^{O\left({\delta}_{m}\right)}r^{\frac{m-1}{2}}$
-[…]
+The restriction conjecture is closely related to a Kakeya-type problem concerning the continuum incidence theory of tubes.
+More precisely, let $\overrightarrow{S}_{s,t}=(S_{s},\dots,S_{t})$ be a multigrain with $S_{i}=(Z_{i},B(y_{i},r_{i}))$.
+Suppose $\mathcal{L}$ is a collection of $r_{t}^{-1/2}$-direction-separated rays $L$ satisfying
+$L\cap B(y_{i},r_{i})\cap N_{\rho_{j}}Z_{i}\neq\emptyset$
+and
+$L\cap B(y_{i},2r_{i})\subseteq N_{\rho_{j}}Z_{i}\cap B(y_{i},2r_{i})$
+then what is the maximal possible cardinality of $\mathcal{L}$?
+Understanding this incidence problem is crucial for obtaining effective estimates for $Ef$.
+It is well known that the restriction conjecture implies the Kakeya conjecture, the latter providing a heuristic bound on the number of incidences between direction-separated tubes.
+In [12, 8, 10], a multiscale structure, referred to as nested tubes, is introduced to describe the spatial relationships between tubes at different scales.
 Definition 2.25 (Nested Tubes).
@@
 $\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\mathbb{T}\left[{\overrightarrow{S}}^{n-1}_{m}\right],\ \ \mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\ \mathrm{\Theta}\left[{\overrightarrow{S}}^{n-1}_{m}\right]$
-One can observe that the second condition in Definition 2.15 is weaker than Lemma 2.5.
-However, this is good enough for our arguments.
-In [12, 8, 10] a nontrivial extension of Theorem 2.14, which is a corollary to Lemma 2.11 in [24], was used respectively to establish a new result on the Bochner-Riesz problem, Stein’s square function, and the Fourier restriction for paraboloid.
-The result is shown below:
+In [12, 8, 10], a result on nested tubes as a corollary of Lemma 2.11 in [25]—was used to obtain new results on the Bochner–Riesz problem, Stein’s square function, and Fourier restriction for the paraboloid.
+We state this result below.
 Lemma 2.26 ([12], Lemma 3.7, Nested polynomial Wolff axioms).
@@
 $\mathrm{\Theta}\left[{\overrightarrow{S}}^{n}_{m}\right]{\lesssim}_{d}r^{O\left({\delta}_{0}\right)}_{n}\left(\prod^{n-1}_{j=m}{r^{-\frac{1}{2}}_{j}}\right)r^{\frac{n-1}{2}}_{n}$
-A direct corollary of this lemma is
+Consequently,
 Corollary 2.27.
@@
 Apply Lemma 2.26 to ${\overrightarrow{S}}:=(S_{m},...,S_{n-1},({\mathbb{R}}^{n},B(y_{n-1},r_{n-1})))$. ∎
-In [12, 8, 10] $S_{m}$ is traced back to $S_{n}:=\left({\mathbb{R}}^{n},B_{R}\right)$, and thus Lemma 2.16 gives
-$\mathrm{\Theta}\left[{\overrightarrow{S}}^{n}_{m}\right]{\lessapprox}_{d}\left(\prod^{n-1}_{j=m}{r^{-1/2}_{j}}\right)R^{\left(n-1\right)/2}$
-Since the tubes have $\sim R^{\left(n-1\right)/2}$ directions in the case of paraboloids, this bound yields a gain of $\prod^{n-1}_{j=m}{r^{-1/2}_{j}}$.
-However, in the case of cone, this does not guarantee a gain due to $\mathrm{\Theta}\left[R\right]\sim R^{\left(n-2\right)/2}$.
-This is the why we need Corllary 2.17
-2.4 $k$-broad Norms
-[…]
-$L\left(\tau\right):=\left\{L\left(\theta\right):\theta\subset\tau\right\}$
-Cover $B_{R}$ by finitely overlapping $B_{K^{2}}$. Fix a parameter $1<A\lesssim K^{\epsilon}$. For each $B_{K^{2}}$, define
-${\mu}_{Ef}\left(B_{K^{2}}\right):={\mathop{\mathrm{min}}_{V_{1},\dots,\ V_{A}\ \left(k-1\right)\mathrm{-}\mathrm{subspace\ of}\ {\mathbb{R}}^{n}}\left(\mathop{\mathrm{max}}_{\tau:\angle\left(L\left(\tau\right),V_{a}\right)\gtrsim K^{-2},\forall 1\leq a\leq A}\int_{B_{K^{2}}}{{\left|Ef_{\tau}\right|}^{p}}\right)\ }$
-here
-$\angle\left(L\left(\tau\right),V_{a}\right):={\mathop{\mathrm{min}}_{v\in G\left(\tau\right)}\angle\left(v,V_{a}\right)\ }$
-Compared with the paraboloid case [6], the angel condition is modified to $\angle\left(G\left(\tau\right),V_{a}\right)\gtrsim K^{-2}$, due to some geometric feature of the cone.
-See [15] for details.
-For $U\subseteq{\mathbb{R}}^{n}$ the $k$-board norm over $U$ is defined as follows
-${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}:={\left(\sum_{B_{K^{2}}}{\frac{\left|B_{K^{2}}\cap U\right|}{\left|B_{K^{2}}\right|}}{\mu}_{Ef}\left(B_{K^{2}}\right)\right)}^{\frac{1}{p}}$
-[…]
-Basic Properties The parameter $A$ is used to make $k$-board norm behave like a regular $L^{p}$-norm. Here we recall some inequalities about ${\left\|.\right\|}_{BL^{p}_{k,A}}$.
-Lemma 2.18 (Triangle inequality, [15], Lemma 3.1).
-Suppose that $1\leq p<\infty$, $f=g+h$, and $A=A_{1}+A_{2}$, where $A,A_{1},A_{2}$ are non-negative integers. Then
-${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}\lesssim{\left\|Eg\right\|}_{BL^{p}_{k,A_{1}}\left(U\right)}+{\left\|Eh\right\|}_{BL^{p}_{k,A_{2}}\left(U\right)}$
-Lemma 2.19 (Hölder’s inequality, [15], Lemma 3.2).
-Suppose $1\leq p,p_{1},p_{2}<\infty$, and $0\leq{\alpha}_{1},{\alpha}_{2}\leq 1$ obey ${\alpha}_{1}+{\alpha}_{2}=1$ and
-$\frac{1}{p}=\frac{{\alpha}_{1}}{p_{1}}+\frac{{\alpha}_{2}}{p_{2}}$
-Suppose that and $A=A_{1}+A_{2}$, then
-${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}\leq{\left\|Ef\right\|}^{{\alpha}_{1}}_{BL^{p_{1}}_{k,A_{1}}\left(U\right)}{\left\|Ef\right\|}^{{\alpha}_{2}}_{BL^{p_{2}}_{k,A_{2}}\left(U\right)}$
-Lemma 2.20 (Standard $L^{2}$-estimate, [15], Section 5).
-For $f\in L^{2}\left(2{\overline{B}}^{n-1}\backslash B^{n-1}\right)$ and non-negative integers $A$, there holds
-${\left\|Ef\right\|}^{2}_{BL^{2}_{k,A}\left(B_{r}\right)}\lesssim r{\left\|f\right\|}^{2}_{2}$
-Lemma 2.21 (Vanishing property, [15], Lemma 4.5).
-Suppose $Ef$ is tangent to a transverse complete intersection $\boldsymbol{Z}$ in $B_{r}$. If $\mathrm{deg}\boldsymbol{Z}\lesssim 1\$} and $dim\boldsymbol{Z}\$} $\leq k-1$, then
-${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{r}\right)}\lesssim\mathrm{RapDec}\left(r\right){\left\|f\right\|}_{2}$
-These inequalities will be used frequently in Section 3.
-Now suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\boldsymbol{\mathrm{Z}},B\left(0,r\right)\right]$, where $\boldsymbol{\mathrm{Z}}$ is a $m$-dimensional transverse complete intersection.
-Fix a ball $B\left(y,\rho\right)\subseteq B\left(0,r\right)$.
-To apply the induction on scale, we expect to find a decomposition
-$\tilde{g}=\sum_{b\in\mathcal{B}}{{\tilde{g}}_{b}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$ (2.4)
-where
-${\tilde{g}}_{b}:=\sum_{\left(\zeta,w,L\right)\in\mathbb{T}\left[{\boldsymbol{\mathrm{Z}}}_{b}-y,B\left(0,\rho\right)\right]}{{\tilde{g}}^{L}_{\zeta,w}},\ \ {\boldsymbol{\mathrm{Z}}}_{b}:=\left\{z+b:z\in\boldsymbol{\mathrm{Z}}\right\}$
-Roughly speaking, for each $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$, $N_{{\rho}^{1/2+{\delta}_{m}}}\left({\boldsymbol{\mathrm{Z}}}_{b}\right)\cap B\left(y,\rho\right)$ is a ‘slice’ of $N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{\mathrm{Z}}\right)\cap B\left(y,\rho\right)$.
-Let $g_{b}\left(\xi\right):=e^{-i{\psi}_{y}\left(\xi\right)}{\tilde{g}}_{b}\left(\xi\right)$, then
-$Eg_{b}\left(x\right)\approx{\chi}_{N_{{\rho}^{1/2+{\delta}_{m}}}\left({\boldsymbol{\mathrm{Z}}}_{b}\right)}\left(x\right)Eg\left(x\right)$ (2.5)
-If $\boldsymbol{\mathrm{Z}}$ is a plane, then such a decomposition exists due to Lemma 2.6.
-However, this question will be tricky if ${\mathrm{deg}\boldsymbol{\mathrm{Z}}\ }>1$, since the shapes of $\boldsymbol{\mathrm{Z}}$ is irregular.
-A rigorous version of (2.4) and (2.5) is given by the following statement.
-See [6, 15, 4] for details.
-Lemma 2.22.
-[…]
-${\left\|Eg\right\|}^{p}_{BL^{p}_{k,A}\left(B\left(y,\rho\right)\right)}\lesssim{\mathrm{log}r\ }\sum_{b\in\mathcal{B}}{{\left\|Eg^{\prime}_{b}\right\|}^{p}_{BL^{p}_{k,A}\left(B_{\rho}\cap N_{{\rho}^{1/2+{\delta}_{m}}}\left(\boldsymbol{Z}+b\right)\right)}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$
-and
-[…]
-where
-$g^{\prime}:=\sum_{\left(\theta,v,l\right)\in{\mathcal{T}}^{\prime}}{g^{l}_{\theta,v}}$
-[…]
-Reduction to Broad Estimates The main result of this paper is
-Theorem 2.23.
-Let $2\leq k\leq n-2$ and
-$p>p_{n}\left(k\right):=2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}}$ (2.6)
-Then there exists an integer $A$ such that
-${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}{\lesssim}_{p,\varepsilon}R^{\varepsilon}{\left\|f\right\|}^{2/p}_{2}{\left\|f\right\|}^{1-2/p}_{\infty}$ (2.7)
-Theorem 2.23, the $k$-broad estimate, implies Theorem 1.5, the linear cone restriction estimate. In particular, it follows from
-Theorem 2.24 ([15], Section 6.1).
-For any $R>1$, the estimate (1.3) holds whenever the estimate (2.7) holds, for $p$ in the range
-$p\left(k,n\right)<p\leq\frac{2n}{n-2},p\left(k,n\right):=\left\{\begin{array}[]{c}2+\frac{2}{n-2},\ if\ 2\leq k\leq 3\\ 2+\frac{4}{2n-k-1},\ if\ k>3\end{array}\right$. (2.8)
-We choose a value of $k$ to optimizes the range of $p$ in (1.3). By applying the arguments in [12] to (2.6), one has
-$2+\frac{12}{4n-5+2\left(k-1\right){\left(\frac{k}{n-1}\right)}^{1/2}}\leq p_{n}\left(k\right)\leq 2+\frac{12}{4n-5+2\left(k-1\right){\left(\frac{2n-1}{2k+1}\right)}^{1/2}\frac{k}{n-1}}$
-By writing $k=vn+O\left(1\right)$, one has
-$p_{n}\left(k\right)=2+\frac{6}{2+v^{3/2}}n^{-1}+O\left(n^{-2}\right)$
-From (2.8),
-$p\left(k,n\right)=2+\frac{4}{2-v}n^{-1}+O\left(n^{-2}\right)$ (2.9)
-To obtain the best possible estimate in Theorem 1.5, $v$ should satisfy
-$\frac{6}{2+v^{3/2}}=\frac{4}{2-v}$
-Hence, we choose
-$v^{\frac{1}{2}}={\left(\frac{3}{8}+\frac{1}{2\sqrt{2}}\right)}^{1/3}+{\left(\frac{3}{8}-\frac{1}{2\sqrt{2}}\right)}^{1/3}-\frac{1}{2}=0.67765\dots$
-Plugging this back into (2.9) yields the linear restriction estimate (1.3) holds for
-$p>2+\lambda n^{-1}+O\left(n^{-2}\right)$
-where
-$\lambda=\frac{4}{2-v}=2.59607\dots$
-This gives Theorem 1.5.
-2.5 Transverse Equidistribution Estimates
-By using a version of Heisenberg uncertainty principle, Guth [6] established the transverse equidistribution estimates for paraboloids.
-However, due to the thin thickness of tubes in the case of cone, $Eg$ may be concentrated on an extremely thin layer of $N_{R^{1/2+{\delta}_{m}}}\left(\boldsymbol{\mathrm{Z}}\right)$, and thus this property may fail to hold.
-Figure 3: A scenario that suggests the transverse equidistribution fails.
-For example, let $\boldsymbol{\mathrm{Z}}$ be a $\left(n-1\right)$-plane, and $B\left(y,r\right)\subseteq B\left(0,R\right)$ for some $y\in\boldsymbol{\mathrm{Z}}$ and $r\sim R^{2/3}$.
-Let $T^{l}_{\theta,v}$ be a $R$-tube with $T^{l}_{\theta,v}\subseteq N_{R^{1/2+{\delta}_{n-1}}}\left(\boldsymbol{\mathrm{Z}}\right)\cap B\left(0,R\right)$, $L\left(\theta\right)\parallel\boldsymbol{\mathrm{Z}}$, and $M\left(\theta\right)\bot\boldsymbol{\mathrm{Z}}$.
-Suppose $g$ is concentrated on $T^{l}_{\theta,v}$.
-Since $r^{1/2+{\delta}_{n-1}}\gg R^{\delta}$, there exist $b\in B\left(0,R^{1/2+{\delta}_{m}}\right)$ such that $T^{l}_{\theta,v}\subseteq N_{r^{1/2+{\delta}_{n-1}}}\left(\boldsymbol{\mathrm{Z}}\mathrm{+b}\right)\cap B\left(0,R\right)$.
-Therefore,
-${\left\|Eg_{b}\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}={\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}+\mathrm{RapDec}\left(R\right){\left\|g\right\|}^{2}_{2}$
-Since $r/R\sim R^{-1/3}$, one has
-${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}\gg R^{O\left({\delta}_{n-1}\right)}{\left(\frac{r}{R}\right)}^{1/2}{\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$
-If ${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$ doesn’t vanish, then Lemma 2.4 gives
-${\left\|g_{b}\right\|}^{2}_{2}\gg R^{O\left({\delta}_{n-1}\right)}{\left(\frac{r}{R}\right)}^{1/2}{\left\|g\right\|}^{2}_{2}$
-which suggests the transverse equidistribution estimates fail.
-[…]
-${\mathbb{T}}_{B_{\rho}}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[r\right],T^{l}_{\theta,v}\cap B_{\rho}\neq\emptyset\right\}$
-For a vector space $V$ spanned by ${\left\{v_{i}=\left(x_{i,1},x_{i,2},\dots,-x_{i,n}\right)\right\}}^{m}_{i=1}$, $V^{+}$ is defined by
-$V^{+}:=\left\{\sum^{m}_{i=1}{a_{i}v^{+}_{i}}:a_{i}\in\mathbb{R}\mathrm{,}v^{+}_{i}=\left(x_{i,1},x_{i,2},\dots,x_{i,n}\right),\ \mathrm{f}\mathrm{or\ }1\leq i\leq m\right\}$
-Now we cover $B_{\rho}$ by finitely overlapping $r^{1/2+{\delta}_{m}}$-ball $B$.
-For some suitable constant $c>0$, if $cB\cap B_{\rho}\cap\boldsymbol{\mathrm{Z}}\boldsymbol{\mathrm{\neq}}\emptyset$, then we choose some $x\in cB\cap B_{\rho}\cap\boldsymbol{\mathrm{Z}}$, and use $V_{B}$ to denote $T_{x}\boldsymbol{\mathrm{Z}}$.
-Define
-${\mathcal{B}}_{a}:=\left\{B:\mathrm{\ }\angle\left(V^{+}_{B},V^{\bot}_{B}\right)\gtrsim K^{-2}\right\}$
-${\mathcal{B}}_{b}:=\left\{B:\mathrm{\ }\angle\left(V^{+}_{B},V^{\bot}_{B}\right)\lesssim K^{-2}\right\}$
-${\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[\boldsymbol{\mathrm{Z}},B\left(0,r\right)\right]:T^{l}_{\theta,v}\cap B\neq\emptyset\right\}$
-From the definition above, if $B\in{\mathcal{B}}_{b}$, then $\angle\left(M\left(\theta\right),V^{\bot}_{B}\right)$ may be small for some $\left(\theta,v,l\right)\in{\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}$, and thus the transverse equidistribution estimates may not hold for a function concentrated on ${\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}$.
-Fortunately, there are only a few $\theta$ such that $\left(\theta,v,l\right)\in{\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}$ for some $v,l$.
-Lemma 2.25 ([15], Lemma 5.8).
-Suppose $B\in{\mathcal{B}}_{b}$. Let
-$\mathbb{T}\left[{\theta}^{\prime}\right]:=\left\{\left(\theta,v,l\right):\theta={\theta}^{\prime}\right\},\ \ {\mathrm{\Theta}}_{B}:=\left\{\theta:\mathbb{T}\left[\theta\right]\cap{\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}\neq\emptyset\right\}$
-Then ${\mathrm{\Theta}}_{B}$ is contained in the union of $O\left(1\right)$ many $K^{-2}$-sectors.
-Now, we recall the definition of the essential part $g_{\mathrm{ess}}$ and the tail part $g_{\mathrm{tail}}$.
-${\mathbb{T}}_{\mathrm{ess}}:=\left\{\left({\zeta}_{0},v_{0}\right):\exists\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left.y\right.\mathrm{\ s.t.\ }T^{l}_{\theta,v}\cap B\neq\emptyset\mathrm{\ }\mathrm{fo}\mathrm{r\ some\ }B\in{\mathcal{B}}_{a}\right\}$
-${\mathbb{T}}_{\mathrm{tail}}:=\left\{\left({\zeta}_{0},v_{0}\right):\forall\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left.y\right.\mathrm{,\ }T^{l}_{\theta,v}\cap B=\emptyset\mathrm{\ }\mathrm{fo}\mathrm{r\ all\ }B\in{\mathcal{B}}_{a}\right\}$
-$g_{\mathrm{ess}}:=\sum_{\left({\zeta}_{0},v_{0}\right)\in{\mathbb{T}}_{\mathrm{ess}}}{g_{{\zeta}_{0},v_{0}}},\ \ g_{\mathrm{tail}}:=\sum_{\left({\zeta}_{0},v_{0}\right)\in{\mathbb{T}}_{\mathrm{tail}}}{g_{{\zeta}_{0},v_{0}}}$ (2.10)
-Suppose $A$ is large enough. By the definition of $k$-board norm and Lemma 2.25,
-${\left\|Eg_{\mathrm{tail}}\right\|}^{p}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}\lesssim\sum_{B\in{\mathcal{B}}_{b}}{{\left\|Eg_{\mathrm{tail}}\right\|}^{p}_{BL^{p}_{k,A/2}\left(B\right)}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{p}_{2}\lesssim\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{p}_{2}$ (2.11)
-By Lemma 2.18 and (2.11),
-${\left\|Eg\right\|}_{BL^{p}_{k,A}\left(B_{\rho}\right)}\lesssim{\left\|Eg_{\mathrm{ess}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}+{\left\|Eg_{\mathrm{tail}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}\lesssim{\left\|Eg_{\mathrm{ess}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}+\mathrm{RapDec}\left(R\right){\left\|g\right\|}_{2}$ (2.12)
-[…]
-Transverse Equidistribution Estimates for Essential Part As shown in [15], the transverse equidistribution estimates hold for the essential part $g_{\mathrm{ess}}$.
-The precise statement of the transverse equidistribution estimates for cone is:
-Lemma 2.26 ([15], Lemma 5.13).
-Let $B_{\rho}$ be a $\rho$-ball with $\rho\gg r^{1/2+{\delta}_{m}}$ and $B_{\rho}\cap N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{Z}\right)\cap B\left(0,r\right)\neq\emptyset$.
-Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\boldsymbol{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B_{\rho}}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$  then
-${\left\|g_{\mathrm{ess,}b\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g\right\|}^{2}_{2}$
-where $g_{\mathrm{ess,}b\mathrm{\ }}:={\left(g_{\mathrm{ess\ }}\right)}_{b}$.
-To apply the nested polynomial Wolff axioms, we need to generalize Lemma 2.26.
-Lemma 2.27.
-Let $\rho\gg r^{1/2+{\delta}_{m}}$ and $B\left(y,\rho\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{Z}\right)\cap B\left(0,r\right)\neq\emptyset$.
-Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\boldsymbol{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B\left(y,\rho\right)}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$.
-For $\widetilde{\mathcal{T}}\subseteq\mathbb{T}\left[{\boldsymbol{\mathrm{Z}}}_{b}-y,B\left(0,\rho\right)\right]$, there holds
-${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$
-here $\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}$ denotes the set of $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ for which there exists some $\left(\zeta,w,L\right)\in\widetilde{\mathcal{T}}$ satisfying
-${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq C{\rho}^{-1/2},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y\right),\ T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right)\leq Cr^{1/2+\delta}$
-for some constant $C$, and ${\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}:={\widetilde{\left(g_{\mathrm{ess}}\right)}}_{\widetilde{\mathcal{T}}}$.
-Proof.
-Let $h=g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}$. By Lemma 2.26,
-${\left\|{\tilde{h}}_{\mathrm{ess},b}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}$ (2.13)
-By Lemma 2.1 and 2.6,
-${\left\|{\tilde{h}}_{\mathrm{ess},b}\right\|}^{2}_{2}\gtrsim{\left\|{\widetilde{\left(h_{\mathrm{ess}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}={\left\|{\widetilde{\left(h_{\mathrm{ess,}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.14)
-here $\mathrm{\uparrow}\widetilde{\mathcal{T}}$ denotes the set of $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ for which there exists some $\left(\zeta,w,L\right)\in\widetilde{\mathcal{T}}$ satisfying
-${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq c{\rho}^{-1/2},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y\right),\ T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right)\leq cr^{1/2+\delta}$
-for some large constant $c$. Also, Lemma 2.6 gives
-${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}={\left\|{\widetilde{\left(g_{\mathrm{ess,}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}={\left\|{\widetilde{\left(h_{\mathrm{ess,}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.15)
-Combine (2.13), (2.14), and (2.15), one has
-${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{ess,}\ \mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$
-2.6 Polynomial Partitioning
-[…]
-${\left\|F\right\|}_{L^{1}\left(O_{1}\right)}={\left\|F\right\|}_{L^{1}\left(O_{2}\right)},\ \forall O_{1},O_{2}\in\mathrm{cell}\left(P\right)$ (2.16)
-See [5] for details.
-After a small perturbation and using Sard’s theorem [6], $P$ is guaranteed to be non-singular by relaxing (2.16). The precise statement of the result is as follows.
-Theorem 2.28 ([6], Theorem 5.5).
-Suppose $F\in L^{1}$ and $d$ is a positive integer. Then, there exists a non-singular polynomial $P$ of degree $O\left(d\right)$ such that $\mathrm{cell}\left(P\right)\sim d^{n}$ and
-${\left\|F\right\|}_{L^{1}\left(O\right)}\sim d^{-n}{\left\|F\right\|}_{L^{1}\left(B_{r}\right)},\ \forall O\in\mathrm{cell}\left(P\right)$
-Suppose we have a polynomial $P$ described in Theorem 2.28.
-Let ${\delta}^{\prime}>0$ be a small number.
-Let $\mathcal{O}$ denote the set of connected components of $B_{r}\backslash N_{r^{1/2+{\delta}^{\prime}}}Z\left(P\right)$.
-By pigeonholing, we see that at least one of two cases hold:
-(1).The cellular term dominates, i.e.,
-$\sum_{O\in\mathcal{O}}{{\left\|F\right\|}_{L^{1}\left(O\right)}}\gtrsim{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
-(2).The algebraic term dominates, i.e.,
-${\left\|F\right\|}_{L^{1}\left(N_{r^{1/2+{\delta}^{\prime}}}Z\left(P\right)\right)}\gtrsim{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
-[…]
-Partitioning over lower dimensional sets Theorem 2.28 is not sufficient to prove Theorem 2.23.
-We will use a more involved partitioning result, which partitions a general variety in ${\mathbb{R}}^{n}$.
-Theorem 2.29 ([6], Section 8).
-Fix $r\gg 1$, and $d\in{\mathbb{N}}^{+}$.
-Suppose $F\in L^{1}\left({\mathbb{R}}^{n}\right)$ is supported on $B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}$, where $\boldsymbol{\mathrm{Z}}$ is an $m$-dimensional transverse complete intersection of degree at most $d$.
-Then, at least one of the following cases holds:
-Cellular case. There exists a polynomial $P:{\mathbb{R}}^{n}\to\mathbb{R}$ of degree $O\left(d\right)$ with the following properties:
-[…]
-(2).For each $O\in\mathcal{O}$, ${\left\|F\right\|}_{L^{1}\left(O\right)}\sim d^{-m}{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
-where $\mathcal{O}$ is a refinement of the set of connected components of $B_{r}\backslash N_{r^{1/2+{\delta}_{m}}}Z\left(P\right)$.
-Algebraic case. The exists a $\left(m-1\right)$-dimensional transverse complete intersection $\boldsymbol{\mathrm{Y}}\subseteq\boldsymbol{\mathrm{Z}}$ of degree $O\left(d\right)$ such that
-${\left\|F\right\|}_{L^{1}\left(B_{r}\right)}\lesssim{\left\|F\right\|}_{L^{1}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Y}}\right)}$
-Now we apply Theorem 2.29 to the $k$-board norm estimates:
-(1). If the cellular case holds, then it follows that
-${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}\right)}\lesssim d^{m}{\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(O\right)},\ \ \forall\ O\in\mathcal{O}$
-where $\mathcal{O}$ is the collection of cells produced by Theorem 2.29.
-(2). If the algebraic case holds, then it follows that
-${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}\right)}\lesssim{\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Y}}\right)}$
-where $\boldsymbol{\mathrm{Y}}$ the a $\left(m-1\right)$-dimensional transverse complete intersection produced by Theorem 2.29.
-From the fundamental theorem of algebra, one deduces that
-Lemma 2.30.
-For a $r$-tube $T^{l}_{\theta,v}$,
-$\left\{O\in{\mathcal{O}}^{\prime}:O\cap T^{l}_{\theta,v}\neq\emptyset\right\}\lesssim d$
-3 Finding Polynomial Structure
-In this section, we reformulate the induction arguments in [15] as a recursive process by combining the ideas from [11, 12].
-The argument contains two different algorithms.
-In [alg 1] we apply the iterated polynomial partition until the tangent term dominates.
-This algorithm passes from an $m$-dimensional to an $\left(m-1\right)$-dimensional situation.
-This algorithm corresponds to the induction on the scale.
-[alg 2] is established by repeated application of [alg 1].
-The iteration ends when we arrive in the minimal dimensional case.
-Corollary 2.17, the nested polynomial Wolff axioms, is used to establish the bound.
-This algorithm corresponds to the induction on the dimension.
-Without loss of generality, we assume that the error terms, $\mathrm{RapDec}\left(R\right){\left\|.\right\|}_{2}$, is always negligible in every iteration.
-3.1 The first algorithm
-This algorithm is processed by iterated polynomial partition.
-At each step of the iteration, one of the 3 cases will happen: cellular case, transverse case and tangent case.
-If the tangent case happens, or the scale is tiny enough, then the iteration will end.
-Input. [alg 1] will take the following things as the input:
-$\bullet$ A grain $\left(\boldsymbol{\mathrm{Z}},B\left(y,r\right)\right)$ of dimension $m$.
-$\bullet$ A function $f\in L^{\infty}\left(2{\overline{B}}^{n-1}\backslash B^{n-1}\right)$ which is tangent to $\left(\boldsymbol{\mathrm{Z}},B\left(y,r\right)\right)$.
-$\bullet$ An addmissible large integer $A\in\mathbb{N}$.
-$\bullet$ In the initial step, we define
-$O_{0}:=\left\{B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}\right\},\ \ f_{O_{0}}:=f,\ \ {\hslash}_{0}:=\emptyset,\ \ {\rho}_{0}:=r$
-Output. At the $j$-th interation, [alg 1] outputs:
-$\bullet$ History: A word ${\hslash}_{j}$ of length $j$ in the alphabet $\left\{\mathrm{a,c}\right\}$, which records how the cell $O_{j}\in{\mathcal{O}}_{j}$ were constructed by the iterated polynomial partitioning.
-Let $\mathrm{a}\left(j\right)$ denote the number of ‘a’ in ${\hslash}_{j}$, and $\mathrm{c}\left(j\right)$ denote the number of ‘c’ in ${\hslash}_{j}$.
-$\bullet$ Spatial scale ${\rho}_{j}$: The ${\rho}_{j}$ is determined by the initial scale $r$ and ${\hslash}_{j}$. Define an auxiliary parameter ${\widetilde{\delta}}_{m-1}$ by
-$\left(1-{\widetilde{\delta}}_{m-1}\right)\left(\frac{1}{2}+{\delta}_{m-1}\right)=\frac{1}{2}+{\delta}_{m}$
-One deduces that
-$\delta_{m-1}/2\leq{\widetilde{\delta}}_{m-1}\leq 2{\delta}_{m-1}$
-$\bullet$ Cells: A family of subsets ${\mathcal{O}}_{j}$ of ${\mathbb{R}}^{n}$.
-Each cell $O_{j}\in{\mathcal{O}}_{j}$ is contained in some ${\rho}_{j}$-ball $B_{O_{j}}=B\left(y_{O_{j}},{\rho}_{j}\right)\$.
-$\bullet$ A collection of functions ${\left(f_{O_{j}}\right)}_{O_{j}\in{\mathcal{O}}_{j}}$.
-For each cell $O_{j}$ there exists a translate ${\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{:}=\boldsymbol{\mathrm{Z}}\boldsymbol{\mathrm{+}}x_{{\mathcal{O}}_{j}}$ such that $f_{O_{j}}$ is tangent to the grain $\left({\boldsymbol{\mathrm{Z}}}_{O_{j}},B_{O_{j}}\right)$
-$\bullet$ A large integer $d$ which depends only on the admissible parameters and $\mathrm{deg}\boldsymbol{\mathrm{Z}}\$}.
-$C^{\mathrm{I}}_{j,\delta}\left(d,r\right):=d^{\#\mathrm{c}\left(j\right)\delta}{{\mathrm{log}}^{\#\mathrm{a}\left(j\right)\left(1+\delta\right)}r\ },C^{\mathrm{II}}_{j,\delta}\left(d\right):=d^{\#\mathrm{c}\left(j\right)\delta+n\#\mathrm{a}\left(j\right)\left(1+\delta\right)}$
-$C^{\mathrm{III}}_{j,\delta}\left(d,r\right):=d^{j\delta}r^{\overline{C}\#\mathrm{a}\left(j\right){\delta}_{m}},C^{\mathrm{IV}}_{j,\delta}\left(d,r\right):=C^{\mathrm{III}}_{j,\delta}\left(d,r\right),A_{j}:=3^{-\#\mathrm{a}\left(j\right)}A$
-Property. At the $j$-th interation,
-1. The mass of ${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}$ is concentrated over the $O_{j}\in{\mathcal{O}}_{j}$:
-${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\leq C^{\mathrm{I}}_{j,\delta}\left(d,r\right)\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$ (3.1)
-2. The function $f_{O_{j}}$ satisfies
-$\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|f_{O_{j}}\right\|}^{2}_{2}}\leq C^{\mathrm{II}}_{j,\delta}\left(d\right)d^{\#\mathrm{c}\left(j\right)}{\left\|f\right\|}^{2}_{2}$ (3.2)
-3. Each $f_{O_{j}}$ satisfies for all $1\leq\rho\leq{\rho}_{j}$.
-${\left\|f_{O_{j}}\right\|}^{2}_{2}\leq C^{\mathrm{III}}_{j,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}d^{-\#\mathrm{c}\left(j\right)\left(m-1\right)}{\left\|f\right\|}^{2}_{2}$ (3.3)
-${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq C^{\mathrm{III}}_{j,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ },\ \forall\rho\in\left[1,{\rho}_{j}\right]$ (3.4)
-4. For $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j}\right]$ and $O_{j}\in{\mathcal{O}}_{j}$,
-${\left\|f_{O_{j},\ \mathcal{T}}\right\|}^{2}_{2}\leq C^{\mathrm{IV}}_{j,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\mathcal{T}}\right\|}^{2}_{2}$ (3.5)
-here $\widetilde{f_{O_{j},\ \mathcal{T}}}:={\left(\widetilde{f_{O_{j}}}\right)}_{\mathcal{T}}$ and ${\uparrow}^{j}\mathcal{T}$ denotes the set of wave packets $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ satisfying
-${\mathrm{dist}}_{H}\left(\theta,\zeta\right)\leq c_{j}{\rho}^{-1/2}_{j},\ \ {\mathrm{dist}}_{H}\left(T^{L}_{\zeta,w}\left(y_{O_{j}}\right),T^{l}_{\theta,v}\left(y\right)\cap B_{O_{j}}\right)\leq c_{j}r^{1/2+\delta}$
-for some $\left(\zeta,w,L\right)\in\mathcal{T}$. Here
-$c_{j}:=\sum^{j}_{i=1}{2^{-\left(j-1\right)/2}C}$
-for some absolute constant $C$.
-If $j=0$, these properties are trivial. For $j\geq 1$, we will verify these properties by induction on $j$.
+3. Finding Polynomial Structure
+In this section, we reformulate the induction arguments of [15] as a recursive procedure, combining ideas from [11, 12]. The argument consists of two algorithms.
+In Algorithm 1, we iteratively apply polynomial partitioning until the tangential term dominates.
+This procedure reduces the problem from an $m$-dimensional to an (m1)(m-1)-dimensional setting and corresponds to the induction on scale.
+Algorithm 2 is obtained by repeatedly applying Algorithm 1, and the iteration terminates once the minimal dimension is reached.
+Throughout the iteration, we assume without loss of generality that the error terms $\mathrm{RapDec}(R)\|f\|_{2}$ are negligible.
+3.1. Algorithm 1
+This algorithm proceeds via iterated polynomial partitioning.
+At each step, one of three cases occurs: the cellular case, the transverse case, or the tangential case.
+The iteration terminates either when the tangential case occurs or when the scale becomes sufficiently small.
+Input. Algorithm 1 takes the following data as input:
+  • A grain $(\mathbf{Z},B(y,r))$ of dimension $m$.
+  • A function $f\in L^{\infty}\!\left(2\overline{B}^{\,n-1}\setminus B^{n-1}\right)$ that is tangent to $(\mathbf{Z},B(y,r))$.
+  • An admissible integer $A\in\mathbb{N}$.
+At the initial step, we set
+$O_{0}:=\{\,B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Z}\,\},\qquad f_{O_{0}}:=f,\qquad\mathfrak{h}_{0}:=\emptyset,\qquad\rho_{0}:=r.$
+Output. At the $j$-th iteration, Algorithm 1 outputs:
+  • History.
+A word $\mathfrak{h}_{j}$ of length $j$ in the alphabet $\{\mathrm{a},\mathrm{c}\}$, recording how each cell $O_{j}\in\mathcal{O}_{j}$ is produced by the iterated polynomial partitioning.
+Let $\mathrm{a}(j)$ and $\mathrm{c}(j)$ denote the number of letters $\mathrm{a}$ and $\mathrm{c}$ in $\mathfrak{h}_{j}$, respectively.
+  • Spatial scale $\rho_{j}$. The scale $\rho_{j}$ is determined by the initial scale $r$ and the history $\mathfrak{h}_{j}$. Define an auxiliary parameter $\widetilde{\delta}_{m-1}$ by
+    $(1-\widetilde{\delta}_{m-1})\!\left(\frac{1}{2}+\delta_{m-1}\right)=\frac{1}{2}+\delta_{m},$
+    which implies
+    $\frac{\delta_{m-1}}{2}\leq\widetilde{\delta}_{m-1}\leq 2\delta_{m-1}.$
+  • Cells. A family of subsets $\mathcal{O}_{j}\subset\mathbb{R}^{n}$. Each cell $O_{j}\in\mathcal{O}_{j}$ is contained in a ball $B_{O_{j}}=B(y_{O_{j}},\rho_{j})$.
+  • Localized functions.
+A collection of functions $\{f_{O_{j}}\}_{O_{j}\in\mathcal{O}_{j}}$.
+For each cell $O_{j}$ there exists a translate $\mathbf{Z}_{O_{j}}:=\mathbf{Z}+x_{O_{j}}$ such that $f_{O_{j}}$ is tangent to the grain $(\mathbf{Z}_{O_{j}},B_{O_{j}})$.
+  • Auxiliary parameters. A large integer $d$ depending only on the admissible parameters and $\deg\mathbf{Z}$, together with
+    $C^{\mathrm{I}}_{j,\delta}(d,r):=d^{\#\mathrm{c}(j)\delta}\,(\log r)^{\#\mathrm{a}(j)(1+\delta)},\qquad C^{\mathrm{II}}_{j,\delta}(d):=d^{\#\mathrm{c}(j)\delta+n\#\mathrm{a}(j)(1+\delta)},$
+    $C^{\mathrm{III}}_{j,\delta}(d,r):=d^{j\delta}r^{\overline{C}\#\mathrm{a}(j)\delta_{m}},\qquad C^{\mathrm{IV}}_{j,\delta}(d,r):=C^{\mathrm{III}}_{j,\delta}(d,r),\qquad A_{j}:=3^{-\#\mathrm{a}(j)}A.$
+Properties. At the $j$-th iteration:
+  1. (1)
+    The mass of $\|Ef\|^{p}_{BL^{p}_{k,A}(B_{r})}$ is concentrated on the cells $O_{j}\in\mathcal{O}_{j}$:
+    $\|Ef\|^{p}_{BL^{p}_{k,A}(B_{r})}\leq C^{\mathrm{I}}_{j,\delta}(d,r)\sum_{O_{j}\in\mathcal{O}_{j}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}.$ (3.1)
+  2. (2)
+    The functions $f_{O_{j}}$ satisfy
+    $\sum_{O_{j}\in\mathcal{O}_{j}}\|f_{O_{j}}\|_{2}^{2}\leq C^{\mathrm{II}}_{j,\delta}(d)\,d^{\#\mathrm{c}(j)}\|f\|_{2}^{2}.$ (3.2)
+  3. (3)
+    Each $f_{O_{j}}$ satisfies, for all $1\leq\rho\leq\rho_{j}$,
+    $\|f_{O_{j}}\|_{2}^{2}\leq C^{\mathrm{III}}_{j,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}d^{-\#\mathrm{c}(j)(m-1)}\|f\|_{2}^{2},$ (3.3)
+    and for all $\rho\in[1,\rho_{j}]$,
+    $\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f_{O_{j}}\|_{L^{2}_{\mathrm{avg}}(\theta)}^{2}\leq\\ C^{\mathrm{III}}_{j,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f\|_{L^{2}_{\mathrm{avg}}(\theta)}^{2}$ (3.4)
+  4. (4)
+    For $\mathcal{T}\subseteq\mathbb{T}[\rho_{j}]$ and $O_{j}\in\mathcal{O}_{j}$,
+    $\|f_{O_{j},\mathcal{T}}\|_{2}^{2}\leq C^{\mathrm{IV}}_{j,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\|f_{\uparrow^{j}\mathcal{T}}\|_{2}^{2},$ (3.5)
+    where $(f_{O_{j},\mathcal{T}})^{\widetilde{}}:=({f_{O_{j}}})^{\widetilde{}}_{\mathcal{T}}$, and $\uparrow^{j}\mathcal{T}$ denotes the set of wave packets $(\theta,v,l)\in\mathbb{T}[r]$ such that
+    $\mathrm{dist}_{H}(\theta,\zeta)\leq c_{j}\rho_{j}^{-1/2}$
+    $\mathrm{dist}_{H}\!\left(T_{\zeta,w}^{L}(y_{O_{j}}),\,T_{\theta,v}^{l}(y)\cap B_{O_{j}}\right)\leq c_{j}r^{1/2+\delta}$
+    for some $(\zeta,w,L)\in\mathcal{T}$, where
+    $c_{j}:=\sum_{i=1}^{j}2^{-(j-1)/2}C$
+    for an absolute constant $C$.
 Recursion Steps.
 If $\rho_{j}\leq r^{\widetilde{\delta}_{m-1}}$, then the algorithm terminates.
-We set $J:=j$ and $\mathcal{O}:={\mathcal{O}}_{j}$.
-We mark this situation as [tiny].
-Otherwise, for each $O_{j}\in{\mathcal{O}}_{j}$ we apply the polynomial partitioning, Theorem 2.29, with degree $d$ to ${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}$.
-Let ${\mathcal{O}}_{j\mathrm{,cell}}\subseteq{\mathcal{O}}_{j}$ denote the set of all cells for which the cellular case holds and ${\mathcal{O}}_{j\mathrm{,alg}}:={\mathcal{O}}_{j}\backslash{\mathcal{O}}_{j\mathrm{,cell}}$.
+In this case we set $J:=j$ and $\mathcal{O}:=\mathcal{O}_{j}$, and label this outcome as [tiny].
+Otherwise, for each $O_{j}\in\mathcal{O}_{j}$ we apply polynomial partitioning with degree $d$ to $\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}$.
+Let $\mathcal{O}_{j,\mathrm{cell}}\subseteq\mathcal{O}_{j}$ denote the collection of cells for which the cellular case occurs, and set
+$\mathcal{O}_{j,\mathrm{alg}}:=\mathcal{O}_{j}\setminus\mathcal{O}_{j,\mathrm{cell}}.$
 then one of the following cases holds:
-[…]
-Cellular-dominant case Suppose
-$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}\leq\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$ (3.6)
-Define ${\hslash}_{j+1}$ by appending the $\mathrm{c}$ to ${\hslash}_{j}$. Now
-$\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right),\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right)+1,\ \ A_{j+1}=A_{j}$
-$C^{\mathrm{I}}_{j+1,\delta}\left(d,r\right)=d^{\delta}C^{\mathrm{I}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{II}}_{j+1,\delta}\left(d\right)=d^{\delta}C^{\mathrm{II}}_{j,\delta}\left(d\right)$
-$C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right)=d^{\delta}C^{\mathrm{III}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right)=d^{\delta}C^{\mathrm{IV}}_{j,\delta}\left(d,r\right)$
-Also, let ${\rho}_{j+1}:={\rho}_{j}/2$. Since the cellular case happens, for each $O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}$, there exists a polynomial $P$ with degree $O\left(d\right)$ such that
-1. $\mathcal{O}\left(O_{j}\right)\ \sim\ d^{m}$
-2. For all $\in\mathcal{O}, OO$ is included in a ${\rho}_{j+1}$-ball $B_{O}:=B\left(y_{O},{\rho}_{j+1}\right)$.
-3. ${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O\right)}\sim d^{-m}{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)},\ \ \forall O\in\mathcal{O}$
-where $\mathcal{O}\left(O_{j}\right)$ is a refinement of the set of connected components of $O_{j}\backslash N_{{\rho}^{1/2+{\delta}_{m}}_{j}}Z\left(P\right)$.
-For $O\in\mathcal{O}\left(O_{j}\right)$, define
-$f_{O}:=\sum_{\begin{array}[]{c}\left(\theta,v,l\right)\in\mathbb{T}\left[{\rho}_{j}\right]\\ T^{l}_{\theta,v}\cap O\neq\emptyset\end{array}}{{\left(f_{O_{j}}\right)}^{l}_{\theta,v}}$
-By Lemma 2.30, each $T^{l}_{\theta,v}$ intersects $O\left(d\right)$ cells, and thus Lemma 2.1 gives
-$\sum_{O\in\mathcal{O}\left(O_{j}\right)}{{\left\|f_{O}\right\|}^{2}_{2}}\lesssim d{\left\|f_{O_{j}}\right\|}^{2}_{2}$
-By pigeonholing, there exists a refinement ${\mathcal{O}}^{\prime}\left(O_{j}\right)\subseteq\mathcal{O}\left(O_{j}\right)$ with ${\mathcal{O}}^{\prime}\ \sim\ d^{m}$ such that
-${\left\|f_{O}\right\|}^{2}_{2}\lesssim d^{-m+1}{\left\|f_{O_{j}}\right\|}^{2}_{2},\ \ \forall O\in{\mathcal{O}}^{\prime}\left(O_{j}\right)$
-Since ${\rho}_{j+1}\sim{\rho}_{j}$, there exists $\mathcal{B}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ with # $\mathcal{B}\ \sim\ 1$ such that
-$N_{{\rho}^{1/2+{\delta}_{m}}_{j}}{\boldsymbol{\mathrm{Z}}}_{O_{j}}\subseteq\bigcup_{b\in\mathcal{B}}{N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{+}b\right)}$
-For $O\in{\mathcal{O}}^{\prime}\left(O_{j}\right)$ and $b\in\mathcal{B}$, define
-$O_{b}:=O\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{+}b\right),\ \ \widetilde{f_{O_{b}}}:=\sum_{\begin{array}[]{c}\left(\zeta,w,L\right)\in\mathbb{T}\left[{\rho}_{j+1}\right]\\ T^{L}_{\zeta,w}\left(y_{O}\right)\cap O_{b}\neq\emptyset\end{array}}{{\left(\widetilde{f_{O}}\right)}^{L}_{\zeta,w}}$
-here $\widetilde{f_{O}}\left(\xi\right):=e^{i{\psi}_{y_{O}}\left(\xi\right)}f_{O}\left(\xi\right)$.
-By Lemma 2.6, $f_{O_{b}}$ is tangent to the grain $\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{+}b,B_{O}\right)$.
-Let
-${\mathcal{O}}_{j+1}:=\bigcup_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{\bigcup_{O\in{\mathcal{O}}^{\prime}\left(O_{j}\right)}{\left\{O_{b}:\ b\in\mathcal{B}\right\}}}$
-Now we verify the properties. For Property 1, 2, and 3, see [11] for details. Here we verify Property 4. By Lemma 2.1 and 2.6, for $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$,
-${\left\|{\tilde{f}}_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$ (3.7)
-where $\uparrow\mathcal{T}$ denotes set the $\left(\theta,v,l\right)\in\mathbb{T}\left[{\rho}_{j}\right]$ such that
-${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq C{\rho}^{-\frac{1}{2}}_{j+1},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y_{O_{j+1}}\right),\ T^{l}_{\theta,v}\cap B\left(y_{O_{j+1}},{\rho}_{j+1}\right)\right)\leq C{\rho}^{1/2+\delta}_{j}$
-for some $\left(\zeta,w,L\right)\in\mathcal{T}$. Here $C$ is the absolute constant that appears in Property 4. Combine (3.5) and (3.7),
-${\left\|f_{O_{j+1},\ \mathcal{T}}\right\|}^{2}_{2}\lesssim d^{-\delta}C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
-Since ${\rho}_{j+1}\ \sim\ {\rho}_{j}$, by choosing $d$ large enough, one has
-${\left\|f_{O_{j+1},\ \mathcal{T}}\right\|}^{2}_{2}\leq C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
-[…]
+Cellular-dominant case. Suppose
+$\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{alg}}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}\leq\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{cell}}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}.$ (3.6)
+Append the letter $\mathrm{c}$ to $\mathfrak{h}_{j}$ to obtain $\mathfrak{h}_{j+1}$. Then
+$\mathrm{a}(j+1)=\#\mathrm{a}(j),\qquad\#\mathrm{c}(j+1)=\#\mathrm{c}(j)+1,\qquad A_{j+1}=A_{j},$
+and
+$C^{\mathrm{I}}_{j+1,\delta}=d^{\delta}C^{\mathrm{I}}_{j,\delta},\qquad C^{\mathrm{II}}_{j+1,\delta}=d^{\delta}C^{\mathrm{II}}_{j,\delta},$
+$C^{\mathrm{III}}_{j+1,\delta}=d^{\delta}C^{\mathrm{III}}_{j,\delta},\qquad C^{\mathrm{IV}}_{j+1,\delta}=d^{\delta}C^{\mathrm{IV}}_{j,\delta}.$
+Set $\rho_{j+1}:=\rho_{j}/2$. For each $O_{j}\in\mathcal{O}_{j,\mathrm{cell}}$, the cellular case yields a polynomial $P$ of degree $O(d)$ such that:
+  1. (1)
+    $\mathcal{O}(O_{j})\sim d^{m}$,
+  2. (2)
+    every $O\in\mathcal{O}(O_{j})$ lies in a ball $B_{O}=B(y_{O},\rho_{j+1})$,
+  3. (3)
+    $\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O)}\sim d^{-m}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O_{j})}$.
+Here $\mathcal{O}(O_{j})$ refines the connected components of $O_{j}\setminus N_{\rho_{j}^{1/2+\delta_{m}}}Z(P)$. Define
+$f_{O}:=\sum_{\begin{subarray}{c}(\theta,v,l)\in\mathbb{T}[\rho_{j}]\\ T_{\theta,v}^{l}\cap O\neq\emptyset\end{subarray}}(f_{O_{j}})_{\theta,v}^{l}.$
+By Lemma 2.21 and Lemma 2.1,
+$\sum_{O\in\mathcal{O}(O_{j})}\|f_{O}\|_{2}^{2}\lesssim d\,\|f_{O_{j}}\|_{2}^{2}.$
+Pigeonholing gives a refinement $\mathcal{O}^{\prime}(O_{j})$ with $\mathcal{O}^{\prime}\sim d^{m}$ such that
+$\|f_{O}\|_{2}^{2}\lesssim d^{-m+1}\|f_{O_{j}}\|_{2}^{2}.$
+Since $\rho_{j+1}\sim\rho_{j}$, there exists a finite set $\mathcal{B}\subset B(0,\rho_{j}^{1/2+\delta_{m}})$ with $\mathcal{B}\sim 1$ such that
+$N_{\rho_{j}^{1/2+\delta_{m}}}\mathbf{Z}_{O_{j}}\subset\bigcup_{b\in\mathcal{B}}N_{\rho_{j+1}^{1/2+\delta_{m}}}(\mathbf{Z}_{O_{j}}+b).$
+For $O\in\mathcal{O}^{\prime}(O_{j})$ and $b\in\mathcal{B}$ define
+$O_{b}:=O\cap N_{\rho_{j+1}^{1/2+\delta_{m}}}(\mathbf{Z}_{O_{j}}+b),$
+$\widetilde{f_{O_{b}}}:=\sum_{\begin{subarray}{c}(\zeta,w,L)\in\mathbb{T}[\rho_{j+1}]\\ T_{\zeta,w}^{L}(y_{O})\cap O_{b}\neq\emptyset\end{subarray}}(\widetilde{f_{O}})_{\zeta,w}^{L},\qquad\widetilde{f_{O}}(\xi)=e^{i\psi_{y_{O}}(\xi)}f_{O}(\xi).$
+By Lemma 2.6, $f_{O_{b}}$ is tangent to $(\mathbf{Z}_{O_{j}}+b,B_{O})$. Set
+$\mathcal{O}_{j+1}:=\bigcup_{O_{j}\in\mathcal{O}_{j,\mathrm{cell}}}\bigcup_{O\in\mathcal{O}^{\prime}(O_{j})}\{O_{b}:\,b\in\mathcal{B}\}.$
+Properties 1–3 follow from [11]. We verify Property 4. By Lemma 2.1 and Lemma 2.6, for $\mathcal{T}\subseteq\mathbb{T}[\rho_{j+1}]$,
+$\|\widetilde{f}_{O_{j+1},\mathcal{T}}\|_{2}^{2}\lesssim\|f_{O_{j},\uparrow\mathcal{T}}\|_{2}^{2}.$ (3.7)
+Here $\uparrow\mathcal{T}$ denotes the set of wave packets $(\theta,v,l)\in\mathbb{T}[\rho_{j}]$ such that
+$\mathrm{Dist}_{H}(\theta,\zeta)\leq C\rho_{j+1}^{-1/2},\qquad\mathrm{Dist}_{H}\!\left(T_{\zeta,w}^{L}(y_{O_{j+1}}),\,T_{\theta,v}^{l}\cap B(y_{O_{j+1}},\rho_{j+1})\right)\leq C\rho_{j}^{1/2+\delta}$
+for some $(\zeta,w,L)\in\mathcal{T}$, where $C$ is the absolute constant from Property 4.Combining (3.5) and (3.7), we obtain
+$\|f_{O_{j+1},\mathcal{T}}\|_{2}^{2}\lesssim d^{-\delta}C^{\mathrm{IV}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\|f_{\uparrow^{j}(\uparrow\mathcal{T})}\|_{2}^{2}.$
+Since $\rho_{j+1}\sim\rho_{j}$, choosing $d$ sufficiently large yields
+$\|f_{O_{j+1},\mathcal{T}}\|_{2}^{2}\leq C^{\mathrm{IV}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j+1}}\right)^{-(n-m)/2}\|f_{\uparrow^{j}(\uparrow\mathcal{T})}\|_{2}^{2}.$
+From [12] we see ${\uparrow}^{j}\left(\uparrow\mathcal{T}\right)\subseteq{\uparrow}^{j+1}\mathcal{T}$. This completes the verification.
+Algebraic-dominant Case. If (3.6) fails, then
 $\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}\leq\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$
-Define ${\hslash}_{j+1}$ by appending the $\mathrm{a}$ to ${\hslash}_{j}$. Now
+Append the letter $\mathrm{a}$ to $\mathfrak{h}_{j}$ to obtain $\mathfrak{h}_{j+1}$. Now we have
 $\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right)+1,\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right),\ \ A_{j+1}=A_{j}/3$
@@
 $C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right)=r^{\overline{C}{\delta}_{m}}d^{\delta}C^{\mathrm{III}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right)=r^{\overline{C}{\delta}_{m}}d^{\delta}C^{\mathrm{IV}}_{j,\delta}\left(d,r\right)$
-Let ${\rho}_{j+1}:={\rho}^{1-{\widetilde{\delta}}_{m-1}}_{j}$.
-Since the algebraic case happens, for each $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$, there exists a $\left(m-1\right)$-dimensional transverse complete intersection $\boldsymbol{\mathrm{Y}}\subseteq{\boldsymbol{\mathrm{Z}}}_{O_{j}}$ of degree $C_{\mathrm{alg}}d$ such that
-${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}\lesssim{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\boldsymbol{\mathrm{Y}}\right)}$
-where $C_{\mathrm{alg}}$ is large constant.
-Cover $O_{j}$ by finitely overlapping ${\rho}_{j+1}$-balls $B$.
-Let ${\mathcal{B}}_{O_{j}}$ denotes the set of such balls.
-For each $B\in{\mathcal{B}}_{O_{j}}$, let
-${\mathbb{T}}_{B}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[{\rho}_{j}\right]:T^{l}_{\theta,v}\cap B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\boldsymbol{\mathrm{Y}}\neq\emptyset\right\},f_{B}:={\left(f_{O_{j}}\right)}_{{\mathbb{T}}_{B}}$
-Let ${\mathbb{T}}_{B\mathrm{,tang}}\subseteq{\mathbb{T}}_{B}$ be the set of ${\rho}_{j}$-scale wave packets $\left(\theta,v,l\right)$ such that
-1. $T^{l}_{\theta,v}\cap 2B\subseteq N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\boldsymbol{\mathrm{Y}}\boldsymbol{=}N_{{\rho}^{1/2+{\delta}_{m-1}}_{j+1}}\boldsymbol{\mathrm{Y}}$
-2. If $x\in T^{l}_{\theta,v}$ and $y\in\boldsymbol{\mathrm{Y}}\mathrm{\cap}2B$ satisfy $\left|y-x\right|\lesssim{\rho}^{1/2+{\delta}_{m}}_{j}=\ {\rho}^{1/2+{\delta}_{m-1}}_{j+1}$, then
-$\angle\left(L\left(\theta\right),T_{y}\boldsymbol{\mathrm{Y}}\right)\lesssim{\rho}^{-1/2+{\delta}_{m-1}}_{j-1}$
-Let ${\mathbb{T}}_{B\mathrm{,trans}}:={\mathbb{T}}_{B}\backslash{\mathbb{T}}_{B\mathrm{,tang}}$ and
-$f_{B,\mathrm{tang}}:={\left(f_{O_{j}}\right)}_{{\mathbb{T}}_{B\mathrm{,tang}}},\ \ f_{B,\mathrm{trans}}:={\left(f_{O_{j}}\right)}_{{\mathbb{T}}_{B\mathrm{,trans}}}$
+Set $\rho_{j+1}:=\rho_{j}^{\,1-\widetilde{\delta}_{m-1}}$.
+Since the algebraic case occurs, for each $O_{j}\in\mathcal{O}_{j,\mathrm{alg}}$ there exists an (m1)(m-1)-dimensional transverse complete intersection $\mathbf{Y}\subseteq\mathbf{Z}_{O_{j}}$ of degree $C_{\mathrm{alg}}d$ such that
+${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}\lesssim{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\bm{\mathrm{Y}}\right)}$
+where $C_{\mathrm{alg}}$ is a sufficiently large constant. Let $\mathcal{B}_{O_{j}}$ denote the collection of such balls. For each $B\in\mathcal{B}_{O_{j}}$, define
+$\mathbb{T}_{B}:=\left\{(\theta,v,l)\in\mathbb{T}[\rho_{j}]:T_{\theta,v}^{l}\cap B\cap N_{\rho_{j}^{1/2+\delta_{m}}}\mathbf{Y}\neq\emptyset\right\},\qquad f_{B}:=(f_{O_{j}})_{\mathbb{T}_{B}}.$
+Let $\mathbb{T}_{B,\mathrm{tang}}\subseteq\mathbb{T}_{B}$ be the set of $\rho_{j}$-scale wave packets $(\theta,v,l)$ satisfying:
+  1. (1)
+    $T_{\theta,v}^{l}\cap 2B\subseteq N_{\rho_{j}^{1/2+\delta_{m}}}\mathbf{Y}=N_{\rho_{j+1}^{1/2+\delta_{m-1}}}\mathbf{Y}$;
+  2. (2)
+    whenever $x\in T_{\theta,v}^{l}$ and $y\in\mathbf{Y}\cap 2B$ satisfy $|x-y|\lesssim\rho_{j}^{1/2+\delta_{m}}=\rho_{j+1}^{1/2+\delta_{m-1}}$, then
+    $\angle\!\big(L(\theta),T_{y}\mathbf{Y}\big)\lesssim\rho_{j+1}^{-1/2+\delta_{m-1}}.$
+Define $\mathbb{T}_{B,\mathrm{trans}}:=\mathbb{T}_{B}\setminus\mathbb{T}_{B,\mathrm{tang}}$ and
+$f_{B,\mathrm{tang}}:=(f_{O_{j}})_{\mathbb{T}_{B,\mathrm{tang}}},\qquad f_{B,\mathrm{trans}}:=(f_{O_{j}})_{\mathbb{T}_{B,\mathrm{trans}}}.$
 By Lemma 2.12,
-${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}\ }\left(O_{j}\right)}\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(B\right)}}\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(B\right)}}$
-$\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{tang}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}+\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{trans}}\right\|}^{p}_{BL^{p}_{k,2A_{j+1}}\left(B\right)}}$ (3.8)
-By pigeonholing, at least one of the following situations holds.
-Tangential subcase: Let $C_{\mathrm{tang}}$ be a large constant. If
-$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\leq\left(C_{\mathrm{tang}}/2\right)\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{tang}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}}$ (3.9)
-then the algorithm terminates, and we set $J=j+1$. We mark this situation as [tang]. Let $\mathcal{S}$ be a collection of grains, which is defined by
-$\mathcal{S}:=\left\{\left(\boldsymbol{\mathrm{Y}},B\right):B\in{\mathcal{B}}_{O_{j}}\mathrm{\ for\ some\ }O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}\right\}$
-For $S=\left(\boldsymbol{\mathrm{Y}},B\right)\in\mathcal{S}$, define $f_{S}:=f_{B,\mathrm{tang}}$ and $B_{S}:=B$. Then Lemma 2.6 implies that $f_{S}$ is tangent to $S$. Since (3.6) fails, (3.9) gives
-$\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\leq C_{\mathrm{tang}}\sum_{S\in\mathcal{S}}{{\left\|Ef_{S}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B_{S}\right)}}$ (3.10)
-By Lemma 2.1 and Corollary 2.3, for all $1\leq\rho\leq{\rho}_{j+1}$, there holds
+${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}\ }\left(O_{j}\right)}\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{tang}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}+\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{trans}}\right\|}^{p}_{BL^{p}_{k,2A_{j+1}}\left(B\right)}}$ (3.8)
+By a pigeonholing argument, at least one of the following cases holds.
+Tangential Subcase. Let $C_{\mathrm{tang}}$ be a large constant. If
+$\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{alg}}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O_{j})}\leq\frac{C_{\mathrm{tang}}}{2}\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{alg}}}\sum_{B\in\mathcal{B}_{O_{j}}}\|Ef_{B,\mathrm{tang}}\|^{p}_{BL^{p}_{k,A_{j+1}}(B)},$ (3.9)
+then the algorithm terminates. We set $J=j+1$ and label this outcome [tang]. Define the collection of grains
+$\mathcal{S}:=\left\{(\mathbf{Y},B):B\in\mathcal{B}_{O_{j}}\ \text{for some}\ O_{j}\in\mathcal{O}_{j,\mathrm{alg}}\right\}.$
+For each $S=(\mathbf{Y},B)\in\mathcal{S}$, set $f_{S}:=f_{B,\mathrm{tang}}$ and $B_{S}:=B$. Lemma 2.6 implies that $f_{S}$ is tangent to $S$. Since (3.6) fails, inequality (3.9) yields
+$\sum_{O_{j}\in\mathcal{O}_{j}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O_{j})}\leq C_{\mathrm{tang}}\sum_{S\in\mathcal{S}}\|Ef_{S}\|^{p}_{BL^{p}_{k,A_{j+1}}(B_{S})}.$ (3.10)
+By Lemma 2.1 and Corollary 2.3, for all $1\leq\rho\leq\rho_{j+1}$, we have
 ${\mathop{\mathrm{max}}_{S\in\mathcal{S}}{\left\|f_{S}\right\|}^{2}_{2}\ }\leq C_{\mathrm{tang}}\mathop{\mathrm{max}}_{O_{j}\in{\mathcal{O}}_{j}}{\left\|f_{O_{j}}\right\|}^{2}_{2}$ (3.11)
@@
 $\sum_{S\in\mathcal{S}}{{\left\|f_{S}\right\|}^{2}_{2}}\leq C_{\mathrm{tang}}r^{n{\widetilde{\delta}}_{m-1}}\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|f_{O_{j}}\right\|}^{2}_{2}}$ (3.13)
-Also, for $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$ and $S=\left(Y,B\right)\in\mathcal{S}$ with $B\in{\mathcal{B}}_{O_{j}}$, Lemma 2.1 and Corollary 2.7 gives
+Moreover, for $\mathcal{T}\subseteq\mathbb{T}[\rho_{j+1}]$ and $S=(\mathbf{Y},B)\in\mathcal{S}$ with $B\in\mathcal{B}_{O_{j}}$, Lemma 2.1 together with Corollary 2.7 yields
 ${\left\|f_{S,\mathcal{T}}\right\|}^{2}_{2}\leq C_{\mathrm{tang}}{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$ (3.14)
-where $\widetilde{f_{S,\mathcal{T}}}:={\left(\widetilde{f_{S}}\right)}_{\mathcal{T}}$
+where $\big(f_{S,\mathcal{T}}\big)^{\widetilde{}}:={\left(f_{S}\right)}_{\mathcal{T}}^{\widetilde{}}$
 Transverse Subcase: If (3.9) fails, then (3.8) gives
 $\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\lesssim\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{t}\mathrm{rans}}\right\|}^{p}_{BL^{p}_{k,2A_{j+1}}\left(B\right)}}}$
-Let $f_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}:={\left(f_{B,\mathrm{t}\mathrm{rans}}\right)}_{\mathrm{ess}}$. From (2.12),
+Let $f_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}:={\left(f_{B,\mathrm{t}\mathrm{rans}}\right)}_{\mathrm{ess}}$. By (2.5),
 $\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\lesssim\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}}$
 Fix $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$ and $B\in{\mathcal{B}}_{O_{j}}$.
-By Lemma 2.22, there exist ${\mathfrak{b}}_{B,O_{j}}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ and $g_{b}$ tangent to $\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b,B\right)$ for each $b\in{\mathfrak{b}}_{B,O_{j}}$ such that
-${\left\|Ef_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}\lesssim{\mathrm{log}r\ }\sum_{b\in{\mathfrak{b}}_{B,O_{j}}}{{\left\|Eg_{b}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b\right)\right)}}$
+Arguing as in [6], for each $b\in{\mathfrak{b}}_{B,O_{j}}$ there exist ${\mathfrak{b}}_{B,O_{j}}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ and $g_{b}$ tangent to $\left({\bm{\mathrm{Z}}}_{O_{j}}+b,B\right)$ such that
+${\left\|Ef_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}\lesssim{\mathrm{log}r\ }\sum_{b\in{\mathfrak{b}}_{B,O_{j}}}{{\left\|Eg_{b}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\left({\bm{\mathrm{Z}}}_{O_{j}}+b\right)\right)}}$
 $\sum_{b\in{\mathfrak{b}}_{B,O_{j}}}{{\left\|g_{b}\right\|}^{2}_{2}}\lesssim{\left\|f_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{2}_{2}$
 Let
-$\mathcal{O}\left(O_{j}\right):=\left\{B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b\right):b\in{\mathfrak{b}}_{B,O_{j}},\ B\in{\mathcal{B}}_{O_{j}}\right\}$
-For $O=B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\boldsymbol{\mathrm{Z}}}_{O_{j}}+b\right)\in\mathcal{O}\left(O_{j}\right)$, define $f_{O}:=g_{b}$. Finally, let
+$\mathcal{O}\left(O_{j}\right):=\left\{B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\bm{\mathrm{Z}}}_{O_{j}}+b\right):b\in{\mathfrak{b}}_{B,O_{j}},\ B\in{\mathcal{B}}_{O_{j}}\right\}$
+For $O=B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\bm{\mathrm{Z}}}_{O_{j}}+b\right)\in\mathcal{O}\left(O_{j}\right)$, define $f_{O}:=g_{b}$. Finally, set
 ${\mathcal{O}}_{j+1}:=\bigcup_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\mathcal{O}\left(O_{j}\right)}$
-[…]
-${\left\|f_{O_{j+1}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j}}\right\|}^{2}_{2}$ (3.15)
-Combine (3.15) and (3.3),
-${\left\|f_{O_{j+1}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)-\overline{C}{\delta}_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}d^{-\#\mathrm{c}\left(j+1\right)\left(m-1\right)}{\left\|f\right\|}^{2}_{2}$
-The induction closes if $\overline{C}$ and $d$ is large enough. For a ${\rho}^{-1/2}$-sector $\theta$ with $1\leq\rho\leq{\rho}_{j+1}$, Lemma 2.6, Lemma 2.27, and Corollary 2.2 gives
+We now verify the properties. Properties 1 and 2 follow from [11]; here we verify Properties 3 and 4.
+Property 3. By Lemma 2.18,
+${\left\|f_{O_{j+1}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j}}\right\|}^{2}_{2}$
+Combining this with (3.3), we obtain
+$\|f_{O_{j+1}}\|_{2}^{2}\lesssim r^{O(\delta_{m})-\overline{C}\delta_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}d^{-\#\mathrm{c}(j+1)(m-1)}\|f\|_{2}^{2}.$
+The induction closes provided $\overline{C}$ and $d$ are chosen sufficiently large. For a $\rho^{-1/2}$-sector $\theta$ with $1\leq\rho\leq\rho_{j+1}$, Lemma 2.6, Lemma 2.19, and Corollary 2.2 yield
 ${\left\|f_{O_{j+1}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}\sum_{\begin{array}[]{c}{\theta}^{\prime}:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}\\ {\theta}^{\prime}\cap C\theta\neq\emptyset\end{array}}{{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left({\theta}^{\prime}\right)}}$
 By pigeonholing, there exists a $\rho^{-1/2}$-sector $\theta^{\prime}$ such that $\theta^{\prime}\cap C\theta\neq\emptyset$ and
-${\left\|f_{O_{j+1}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left({\theta}^{\prime}\right)}$ (3.16)
-Combine (3.16) and (3.4),
-${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq r^{O\left({\delta}_{m}\right)-\overline{C}{\delta}_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$
-[…]
-${\left\|{\tilde{f}}_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j}/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$ (3.17)
-Combine (3.5) and (3.17)
+$\|f_{O_{j+1}}\|^{2}_{L^{2}_{\mathrm{avg}}(\theta)}\lesssim r^{O(\delta_{m})}\left(\frac{\rho_{j+1}}{\rho_{j}}\right)^{-(n-m)/2}\|f_{O_{j}}\|^{2}_{L^{2}_{\mathrm{avg}}(\theta^{\prime})}$
+Combining this and (3.4), we obtain
+$\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f_{O_{j+1}}\|^{2}_{L^{2}_{\mathrm{avg}}(\theta)}\lesssim\\ r^{O(\delta_{m})-\overline{C}\delta_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f\|^{2}_{L^{2}_{\mathrm{avg}}(\theta)}$
+Since $\overline{C}$ and $d$ are chosen sufficiently large, this verifies Property 3.
+Property 4. For $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$, Lemma 2.19 implies
+${\left\|{\tilde{f}}_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j}/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$
+Combining this with (3.5) we have
 ${\left\|f_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)-\overline{C}{\delta}_{m}}d^{-\delta}{\left({\rho}_{j}/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
-Since $\overline{C}$ and $d$ is large enough, one has
+Since $\overline{C}$ and $d$ are chosen sufficiently large, one has
 ${\left\|f_{O_{j+1},\ \mathcal{T}}\right\|}^{2}_{2}\leq C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
-Like the cellular case, ${\uparrow}^{j}\left(\uparrow\mathcal{T}\right)\subseteq{\uparrow}^{j+1}\mathcal{T}$, and thus the induction closes.
-Remark 3.1.
-A trivial observation is
-$\mathrm{c}\left(J\right)\lesssim{\mathrm{log}r\ },\ \ \#\mathrm{a}\left(J\right)\lesssim{\delta}^{-1}_{m-1}{\mathrm{log}{\delta}^{-1}_{m-1}\ }$ (3.18)
-Hence, the accumulation of error is still negligible after the algorithm terminates. We omit the details.
-Let $D:=d^{c\left(J\right)}$. If [tang] happens, then (3.1), (3.2), (3.3), (3.4), (3.5), (3.10), (3.11), (3.12), (3.13), and (3.14) gives
-${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox D^{\delta}\sum_{S\in\mathcal{S}}{{\left\|Ef_{S}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(B_{S}\right)}}$ (3.19)
-$\sum_{S\in\mathcal{S}}{{\left\|f_{S}\right\|}^{2}_{2}}\lessapprox D^{1+\delta}{\left\|f\right\|}^{2}_{2}$ (3.20)
-${\mathop{\mathrm{max}}_{S\in\mathcal{S}}{\left\|f_{S}\right\|}^{2}_{2}\ }\lessapprox D^{-\left(m-1\right)+\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.21)
-${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{S}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox D^{\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (3.22)
-${\left\|f_{S,\ \mathcal{T}}\right\|}^{2}_{2}\lessapprox D^{\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{J}\mathcal{T}}\right\|}^{2}_{2}$ (3.23)
+As in the cellular case, we have $\uparrow^{j}(\uparrow\mathcal{T})\subseteq\uparrow^{j+1}\mathcal{T}$, and thus the induction closes.
+Summary. We now summarize the output in a concise form. Since
+$\mathrm{c}\left(J\right)\lesssim\log{r},\ \ \#\mathrm{a}\left(J\right)\lesssim{\delta}^{-1}_{m-1}\log{{\delta}^{-1}_{m-1}}$ (3.15)
+the accumulated error remains negligible after the algorithm terminates.
+Let $D:=d^{c\left(J\right)}$.
+If [tang] happens, then (3.1), (3.2), (3.3), (3.4), (3.5), (3.10), (3.11), (3.12), (3.13), and (3.14) gives
+${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox\sum_{S\in\mathcal{S}}{{\left\|Ef_{S}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(B_{S}\right)}}$ (3.16)
+$\sum_{S\in\mathcal{S}}{{\left\|f_{S}\right\|}^{2}_{2}}\lessapprox D{\left\|f\right\|}^{2}_{2}$ (3.17)
+${\mathop{\mathrm{max}}_{S\in\mathcal{S}}{\left\|f_{S}\right\|}^{2}_{2}\ }\lessapprox D^{-(m-1)}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.18)
+${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{S}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (3.19)
+${\left\|f_{S,\ \mathcal{T}}\right\|}^{2}_{2}\lessapprox{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{J}\mathcal{T}}\right\|}^{2}_{2}$ (3.20)
 If [tiny] happens, then (3.1), (3.2), (3.3), (3.4), and (3.5) gives
-${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox D^{\delta}\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(O\right)}}$ (3.24)
-$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox D^{1+\delta}{\left\|f\right\|}^{2}_{2}$ (3.25)
-${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox D^{-\left(m-1\right)+\delta}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.26)
-3.2 The second algorithm
-[alg 2] is established by repeated application of [alg 1] and terminates if [tiny] term dominates.
+${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(O\right)}}$ (3.21)
+$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox D{\left\|f\right\|}^{2}_{2}$ (3.22)
+${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox D^{-(m-1)}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.23)
+3.2. Algorithm 2
+Algorithm 2 is obtained by repeatedly applying Algorithm 1 and terminates when the [tiny] case dominates.
 Consider a family of Lebesgue exponents $p_{i}$ for $k\leq l\leq n$, which will be determined in Section 4, satisfying
-$p_{k}\geq p_{k+1}\geq\dots\geq p_{n}=:p>2$
-and define $0\leq{\alpha}_{i},{\beta}_{i}\leq 1$ in term of the $p_{i}$ by ${\alpha}_{n}={\beta}_{n}:=1$ and
+$p_{k}\geq p_{k+1}\geq\dots\geq p_{n}=:p>2$ (3.24)
+and define $0\leq{\alpha}_{i},{\beta}_{i}\leq 1$ by ${\alpha}_{n}={\beta}_{n}:=1$ and
 ${\alpha}_{i}:=\left(\frac{1}{2}-\frac{1}{p_{i+1}}\right){\left(\frac{1}{2}-\frac{1}{p_{i}}\right)}^{-1},\ \ {\beta}_{i}:=\left(\frac{1}{2}-\frac{1}{p}\right){\left(\frac{1}{2}-\frac{1}{p_{i}}\right)}^{-1}\ \mathrm{for}\ k\leq i\leq n-1$ (3.25)
@@
 ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\gtrsim R^{\varepsilon}{\left\|f\right\|}_{2}$ (3.26)
-Here $A\in\mathbb{N}$ is a large constant.
-Output: The algorithm outputs the following objects.
-$\bullet$ A collection $\mathcal{O}$ of cells such that each $O\in\mathcal{O}$ is included in a $R^{O\left({\delta}_{0}\right)}$-ball.
-$\bullet$ For each $O\in\mathcal{O}$ a function $f_{O}$.
-$\bullet$ Scales $\overrightarrow{r}=\left(r_{n},\dots,r_{m-1}\right)$ satisfying $R=r_{n}>r_{n-1}>\dots>r_{m-1}=1$.
-$\bullet$ Integers $\overrightarrow{A}=\left(A_{n},\dots,A_{m-1}\right)$ satisfying $A=A_{n}>A_{n-1}>\dots>A_{m-1}$.
-$\bullet$ A tuple of parameters $\overrightarrow{D}=\left(D_{n},\dots,D_{m-1}\right)$.
-$\bullet$ For $m\leq l\leq n$ a family ${\overrightarrow{\mathcal{S}}}_{l}$ of level $\left(n-l+1\right)$ multigrain.
-Each ${\overrightarrow{S}}_{l}=\left(\left(Z_{l},B_{r_{l}}\right),\dots,\left(Z_{n},B_{r_{n}}\right)\right)\in{\overrightarrow{\mathcal{S}}}_{l}$ has a multiscale ${\overrightarrow{r}}_{l}=\left(r_{l},\dots,r_{n}\right)$ and a bounded degree.
-Also, for $l\leq s\leq t\leq n$:
-$Z_{s}=Z^{\prime}_{s}+b\mathrm{,\ \ for\ some\ }b\mathrm{\ with\ }\left|b\right|\lesssim r^{1/2+{\delta}_{t}}_{t}\ \ \mathrm{and\ }Z^{\prime}_{s}\subseteq Z_{t}$
-$\bullet$ For $m\leq l\leq n$ an assignment of a function $f_{{\overrightarrow{S}}_{l}}$ to each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}$.
-Each $f_{{\overrightarrow{S}}_{l}}$ is tangent to $S_{l}=\left(Z_{l},B_{r_{l}}\right)$, the first component of ${\overrightarrow{S}}_{l}$.
+Here, $A\in\mathbb{N}$ is chosen to be sufficiently large.
+Output. The algorithm produces the following objects:
+  • A collection $\mathcal{O}$ of cells, each contained in a $R^{O(\delta_{0})}$-ball.
+  • For each $O\in\mathcal{O}$, a function $f_{O}$.
+  • Scales $\overrightarrow{r}=(r_{n},\dots,r_{m-1})$ satisfying $R=r_{n}>r_{n-1}>\cdots>r_{m-1}=1$.
+  • Integers $\overrightarrow{A}=(A_{n},\dots,A_{m-1})$ satisfying $A=A_{n}>A_{n-1}>\cdots>A_{m-1}$.
+  • A tuple of parameters $\overrightarrow{D}=(D_{n},\dots,D_{m-1})$.
+  • For $m\leq l\leq n$, a family $\overrightarrow{\mathcal{S}}_{l}$ of level (nl+1)(n-l+1) multigrains.
+Each $\overrightarrow{S}_{l}=((Z_{l},B_{r_{l}}),\dots,(Z_{n},B_{r_{n}}))\in\overrightarrow{\mathcal{S}}_{l}$ has multiscale $\overrightarrow{r}_{l}=(r_{l},\dots,r_{n})$ and bounded degree.
+Moreover, for $l\leq s\leq t\leq n$,
+    $Z_{s}=Z^{\prime}_{s}+b\quad\text{for some }b\text{ with }|b|\lesssim r_{t}^{1/2+\delta_{t}},\qquad Z^{\prime}_{s}\subseteq Z_{t}.$
+  • For $m\leq l\leq n$, an assignment of a function $f_{\overrightarrow{S}_{l}}$ to each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}_{l}$.
+Each $f_{\overrightarrow{S}_{l}}$ is tangent to $S_{l}=(Z_{l},B_{r_{l}})$, the first component of $\overrightarrow{S}_{l}$.
 Property:
@@
 ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p_{m}}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}}\right)}^{{\beta}_{m}/p_{m}\ }$ (3.27)
-holds for
+holds, where
 $M\left(\overrightarrow{r},\overrightarrow{D}\right):={D^{\delta}_{m-1}\left(\prod^{n-1}_{i=m}{D_{i}}\right)}^{\left(n-m\right)\delta}\left(\prod^{n-1}_{i=m}{r^{\left({\beta}_{i+1}-{\beta}_{i}\right)/2}_{i}D^{\left({\beta}_{i+1}-{\beta}_{m}\right)/2}_{i}}\right)$ (3.28)
-2.There holds
-$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox\left(\prod^{n-1}_{i=m-1}{D^{1+\delta}_{i}}\right){\left\|f\right\|}^{2}_{2}$ (3.31)
+2.The inequality
+$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox\left(\prod^{n-1}_{i=m-1}{D_{i}}\right){\left\|f\right\|}^{2}_{2}$ (3.29)
+holds.
 3.For $m\leq l\leq n$,
-${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox r^{-\frac{n-l}{2}}_{l}\left(\prod^{l-1}_{i=m-1}{r^{-1/2}_{i}D^{-i+\delta}_{i}}\right){\mathop{\mathrm{max}}_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\ }$ (3.32)
+${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox r^{-\frac{n-l}{2}}_{l}\left(\prod^{l-1}_{i=m-1}{r^{-1/2}_{i}D^{-i}_{i}}\right){\mathop{\mathrm{max}}_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\ }$ (3.30)
 […]
-${\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-1/2}_{j}D^{\delta}_{j-1}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}$ (3.33)
+${\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-1/2}_{j}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}$ (3.31)
 where $f^{\#}_{{\overrightarrow{S}}_{l}}:={\left(f_{{\overrightarrow{S}}_{n-1}}\right)}_{\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{l}\right]}$.
@@
 ${\overrightarrow{\mathcal{S}}}_{n}:=\left\{{\overrightarrow{S}}_{n}\right\},{\overrightarrow{S}}_{n}:=\left(S_{n}\right),S_{n}=\left({\mathbb{R}}^{n},\ B_{R}\right),f_{{\overrightarrow{S}}_{n}}:=f$
-$\left(n+2-l\right)$-th step. Suppose the algorithm has run through $\left(n+1-l\right)$ steps and the following things have been determined.
+$\left(n+2-l\right)$-th step. Suppose the algorithm has run for (n+1l)(n+1-l) steps and the following things have been determined.
 ${\overrightarrow{r}}_{l}=\left(r_{n},\dots,r_{l}\right),\ \ {\overrightarrow{A}}_{l}=\left(A_{n},\dots,A_{l}\right),\ \ {\overrightarrow{D}}_{l}=\left(D_{n},\dots,D_{l}\right),\ \ {\overrightarrow{\mathcal{S}}}_{l}$
-For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}$ we apply [alg 1] to the $k$-board norm ${\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}$.
-Let $\mathrm{a}\left[{\overrightarrow{S}}_{l}\right]$ and $\mathrm{c}\left[{\overrightarrow{S}}_{l}\right]$ denote $\mathrm{a}\left(J\right)$ and $\mathrm{c}\left(J\right)$ output by [alg 1] respectively.
-Let $A\left[{\overrightarrow{S}}_{l}\right]$ and $r\left[{\overrightarrow{S}}_{l}\right]$ denote $A_{J}$ and $r_{J}$ output by [alg 1] respectively.
+For each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}_{l}$, we apply Algorithm 1 to the $k$-broad norm $\|Ef_{\overrightarrow{S}_{l}}\|_{BL^{p_{l}}_{k,A_{l}}(B_{r_{l}})}$.
+Let $r[\overrightarrow{S}_{l}]$, $D[\overrightarrow{S}_{l}]$, and $A[\overrightarrow{S}_{l}]$ denote the quantities $r_{J}$, $D$, and $A_{J}$ returned by Algorithm 1, respectively.
 Let
@@
 ${\overrightarrow{\mathcal{S}}}_{l\mathrm{,tang}}:=\left\{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}\mathrm{:}\mathrm{[alg\ 1]\ terminates\ due\ to\ [t}\mathrm{ang}\mathrm{]}\right\}$
-[…]
+Consequently, at least one of the following two cases must hold:
+Tiny-dominated Case. We say that the tiny-dominated case happens if
 $\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\leq 2\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.32)
-By (3.18) and (3.34), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$ and $A_{l-1}\in\mathbb{N}$ such that $A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1}$ for all ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}$ and
+By (3.15), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$ and $A_{l-1}\in\mathbb{N}$ such that $A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1}$ and $D\left[{\overrightarrow{S}}_{l}\right]=D_{l-1}$ for all ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}$ and
 $\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\lessapprox\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.33)
-For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}$, let $\mathcal{O}\left[{\overrightarrow{S}}_{l}\right]$ denote the final cells output by [alg 1].
-For each $O\in\mathcal{O}\left[{\overrightarrow{S}}_{l}\right]$ let $f_{O}$ denote the function corresponding to $O$ output by [alg 1].
-Finally, [alg 2] outputs the following additional things, and then terminates.
-$m=l,\ \ r_{m-1}:=1,\ \ \mathcal{O}:=\bigcup_{{\overrightarrow{S}}_{m}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{m\mathrm{,tiny}}}{\mathcal{O}\left[{\overrightarrow{S}}_{m}\right]}$
-[…]
-Tangent-dominated case If (3.34) fails, then
-$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\leq 2\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tang}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.36)
-By (3.18) and (3.36), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$, $A_{l-1}\in\mathbb{N}$, and $r_{l-1}\in\left(0,R^{O\left({\delta}_{0}\right)}\right]$ such that
+For each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}^{\prime}_{l,\mathrm{tiny}}$, let $\mathcal{O}[\overrightarrow{S}_{l}]$ denote the final cells produced by Algorithm 1.
+For each $O\in\mathcal{O}[\overrightarrow{S}_{l}]$, let $f_{O}$ be the corresponding function returned by Algorithm 1.
+Algorithm 2 then outputs the following additional data and terminates:
+$m=l,\qquad r_{m-1}:=1,\qquad\mathcal{O}:=\bigcup_{\overrightarrow{S}_{m}\in\overrightarrow{\mathcal{S}}^{\prime}_{m,\mathrm{tiny}}}\mathcal{O}[\overrightarrow{S}_{m}].$
+Tangent-dominated Case. If (3.32) fails, then
+$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\leq 2\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tang}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$
+By (3.15), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$, $A_{l-1}\in\mathbb{N}$, and $r_{l-1}\in\left(0,R^{O\left({\delta}_{0}\right)}\right]$ such that
 $\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\lessapprox\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.34)
 and
-$A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1},\ \ r\left[{\overrightarrow{S}}_{l}\right]=r_{l-1},\ \ \forall{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}$
-For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}$, let $S\left[{\overrightarrow{S}}_{l}\right]$ denote the grains output by [alg 1].
-For each $S_{l\mathrm{-}\mathrm{1}}\in S\left[{\overrightarrow{S}}_{l}\right]$, let ${\overrightarrow{S}}_{l-1}:=\left(S_{l\mathrm{-}\mathrm{1}},{\overrightarrow{S}}_{l}\right)$ and $f_{{\overrightarrow{S}}_{l-1}}$ denote the function corresponding to $S_{l\mathrm{-}\mathrm{1}}$ output by [alg 1].
-Finally, we define
+$A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1},\ \ r\left[{\overrightarrow{S}}_{l}\right]=r_{l-1},\ \ D\left[{\overrightarrow{S}}_{l}\right]=D_{l-1},\ \ \forall{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}$
+For each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}^{\prime}_{l,\mathrm{tang}}$, let $S[\overrightarrow{S}_{l}]$ denote the collection of grains produced by Algorithm 1.
+For each $S_{l-1}\in S[\overrightarrow{S}_{l}]$, define
+$\overrightarrow{S}_{l-1}:=(S_{l-1},\overrightarrow{S}_{l}),$
+and let $f_{\overrightarrow{S}_{l-1}}$ denote the corresponding function returned by Algorithm 1. Finally, we define
 ${\overrightarrow{\mathcal{S}}}_{l\mathrm{-}\mathrm{1}}:=\bigcup_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}}{\left\{{\overrightarrow{S}}_{l-1}:S_{l\mathrm{-}\mathrm{1}}\in S\left[{\overrightarrow{S}}_{l}\right]\right\}}$
-Remark 3.2.
-If $m<k$, then Lemma 2.21 implies
-${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lesssim\mathrm{RapDec}\left(R\right){\left\|f\right\|}_{2}$
-which contradicts (3.28). Thus, $m\geq k$.
-[…]
-Verify the properties Property 2 follows from (3.25) and the repeated application of (3.20).
-Property 3 follows from (3.26) and the repeated application of (3.21).
-Property 1 follows from (3.19), (3.37), (3.24), (3.35), and Lemma 2.19.
-See [11] for details.
-Here we verify Property 4.
-Apply (3.23) repeatedly, one has
-${\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox\prod^{n-2}_{i=l}{D^{\delta}_{i}{\left(\frac{r_{i+1}}{r_{i}}\right)}^{-\frac{n-i-1}{2}}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}=r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-1/2}_{j}D^{\delta}_{j-1}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}$
-This gives Property 4.
-3.3 Establish the $k$-broad estimates
-The task now is how to apply the properties guaranteed by the algorithms [alg 2]. Since each $O\in\mathcal{O}$ has diameter at most $R^{O\left({\delta}_{0}\right)}$, trivially one may bound
-${\left\|Ef_{O}\right\|}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}\lessapprox{\left\|f_{O}\right\|}_{2}$ (3.38)
-Combine (3.29) and (3.38),
-${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p_{m}}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}}\right)}^{{\beta}_{m}/p_{m}\ }$
-$\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\right)}^{{\beta}_{m}/p_{m}\ }{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2{\beta}_{m}\left(\frac{1}{2}-\frac{1}{p_{m}}\right)}_{2}\ }$ (3.39)
-Combine (3.30), (3.27), and (3.29),
-${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{i=m-1}{r^{\frac{{\beta}_{i+1}-{\beta}_{i}}{2}}_{i}D^{\frac{{\beta}_{i+1}}{2}-\left(\frac{1}{2}-\frac{1}{p}\right)+O\left(\delta\right)}_{i}}{\left\|f\right\|}^{\frac{2}{p}}_{2}{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{1-\frac{2}{p}}_{2}\ }$ (3.40)
-By taking $l=n$ in (3.32),
-${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{i=m-1}{r^{-1/2}_{i}D^{-i+\delta}_{i}}\right){\left\|f\right\|}^{2}_{2}$ (3.41)
-Combine (3.40) and (3.41),
-${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{i=m-1}{r^{\frac{{\beta}_{i+1}-{\beta}_{i}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{i}D^{\frac{{\beta}_{i+1}}{2}-\left(i+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)+O\left(\delta\right)}_{i}}{\left\|f\right\|}^{2}_{2}$ (3.42)
-By choosing suitable parameters $p_{i}$ for $k\leq l\leq n$, one can guarantee
-$\frac{{\beta}_{i+1}-{\beta}_{i}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)\leq 0,\ \ \ \ \ \ \ \ \frac{{\beta}_{i+1}}{2}-\left(i+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)<0,\ \ \forall m\leq i\leq n-1$
-This gives Theorem 1.3. See [11] for details.
-4 Proof of Theorem 2.23
-In this section we adapt the method provided in [12] to prove Theorem 2.23.
-Let $f:2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$ and $A$ is a large integer.
-Without loss of generality, we assume that $f$ satisfies (3.27).
-Applying [alg 2] to decompose the $k$-broad norm ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}$, we get the outputs described in Section 3.2.
-First, we improve (3.42) by Corollary 2.17, the nested polynomial Wolff axioms.
+Verification of the Properties.
+Property 2 follows from (3.22) together with repeated applications of (3.17).
+Property 3 follows from (3.23) and repeated applications of (3.18).
+Property 1 follows from (3.16), (3.34), (3.21), (3.33), Lemmas 2.12 and 2.13; see [11] for details.
+We now verify Property 4.
+Applying (3.20) repeatedly yields
+$\|f_{\overrightarrow{S}_{l}}\|_{2}^{2}\lessapprox\prod_{i=l}^{n-2}\left(\frac{r_{i+1}}{r_{i}}\right)^{-(n-i-1)/2}\|f^{\#}_{\overrightarrow{S}_{l}}\|_{2}^{2}=r_{l}^{(n-l-1)/2}\prod_{j=l+1}^{n-1}r_{j}^{-1/2}\|f^{\#}_{\overrightarrow{S}_{l}}\|_{2}^{2}.$
+This establishes Property 4.
+3.3. Establish the Broad Estimates
+We now apply the properties provided by Algorithm 2. Since each $O\in\mathcal{O}$ has diameter at most $R^{O(\delta_{0})}$, we may trivially bound
+${\left\|Ef_{O}\right\|}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}\lessapprox{\left\|f_{O}\right\|}_{2}$
+Together with (3.27), this yields
+${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\right)}^{{\beta}_{m}/p_{m}\ }{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2{\beta}_{m}\left(\frac{1}{2}-\frac{1}{p_{m}}\right)}_{2}\ }$ (3.35)
+Combining this with (3.28) and (3.25), we obtain
+${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{i=m-1}{r^{\frac{{\beta}_{i+1}-{\beta}_{i}}{2}}_{i}D^{\frac{{\beta}_{i+1}}{2}-\left(\frac{1}{2}-\frac{1}{p}\right)}_{i}}{\left\|f\right\|}^{\frac{2}{p}}_{2}{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{1-\frac{2}{p}}_{2}\ }$ (3.36)
+4. Complete the proof
+In this section, we complete the proof of Proposition 2.16. For the reader’s convenience, we restate the proposition below.
+Proposition 4.1.
+Let $n\geq 3$ and $k\in[2,n-2]$. For
+$p=2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}},$
+$\frac{1}{q}=\frac{1}{p}+\frac{1}{4n-8}\big(\frac{6}{p}-1\big),$
+and some integer $A\lesssim_{\varepsilon,n}1$, the estimate
+$\|Ef\|_{BL_{k,A}^{p}(B_{R})}\lesssim_{p,q,\varepsilon}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{q}}\|f\|_{L^{\infty}(\mathbb{R}^{n-1})}^{1-\frac{2}{q}}$
+holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
+Suppose that, after applying Algorithm 2 to decompose the $k$-broad norm $\|Ef\|_{BL^{p}_{k,A}(B_{R})}$, we obtain the outputs described in Section 3.2.
+By Lemma 2.14, we have $m\geq k$.
+Consequently, we divide the proof into cases $n-1\leq m\leq n$ and $k\leq m\leq n-2$.
+Case $n-1\leq m\leq n$ . Taking $l=n$ in (3.30), we obtain
+${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{j=m-1}{r^{-1/2}_{j}D^{-j}_{j}}\right){\left\|f\right\|}^{2}_{2}$ (4.1)
+Combining this with (3.36), we have
+${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{j=m-1}{r^{\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}D^{\frac{{\beta}_{j+1}}{2}-\left(j+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}{\left\|f\right\|}^{2}_{2}$ (4.2)
+Set
+$\begin{cases}p=p_{n}=\frac{2n}{n-1}&\text{if $m=n$},\\[4.0pt] p_{n-1}=\frac{2(n-1)}{n-2},\quad p=p_{n}=\frac{2(2n-1)}{2n-3}&\text{if $m=n-1$}.\end{cases}$ (4.3)
+In both cases, $p<p(n,k)$ and
+$\frac{{\beta}_{i+1}-{\beta}_{i}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)=0,\quad\forall j\in[m-1,n-1]$
+$\frac{{\beta}_{i+1}}{2}-\left(i+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)\leq 0,\quad\forall j\in[m-1,n-1]$
+This finishes the proof in this case.
+Case $k\leq m\leq n-2$ . Applying the nested polynomial Wolff axioms, we have
 Lemma 4.2.
-For $m\leq l<n-1$,
-${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox D^{\delta}_{n-1}\left(\prod^{n-2}_{i=l}{r^{-\frac{1}{2}}_{i}}\right){\left\|f\right\|}^{2}_{\infty}$
+For $m\leq l\leq n-2$,
+${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox\left(\prod^{n-2}_{i=l}{r^{-\frac{1}{2}}_{i}}\right){\left\|f\right\|}^{2}_{\infty}$
 Proof.
-Since $f^{\#}_{{\overrightarrow{S}}_{l}}$ is concentrated on the $r_{n-1}$-scale wave packets $\left(\theta,v,l\right)$ with $\theta\in\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right$], Corollary 2.17 implies
-${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lesssim\#\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right]r^{-\frac{n-2}{2}}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox\left(\prod^{n-2}_{j=m}{r^{-\frac{1}{2}}_{j}}\right){\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (4.1)
+Since $f^{\#}_{\overrightarrow{S}_{l}}$ is concentrated on the $r_{n-1}$-scale wave packets $(\theta,v,l)$ with $\theta\in\Theta[\overrightarrow{S}^{\#}_{l}$], Corollary 2.27 implies
+${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lesssim\#\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right]r^{-\frac{n-2}{2}}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\\ \lessapprox\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right){\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (4.4)
 By Corollary 2.3,
 ${\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lesssim{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f_{{\overrightarrow{S}}_{n-1}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$
-Put (3.22) into this bound,
-${\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox D^{\delta}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq D^{\delta}_{n-1}{\left\|f\right\|}^{2}_{\infty}$ (4.2)
-Now Lemma (4.1) follows from (4.1) and (4.2). ∎
-Combining (3.32), (3.33), and Lemma 4.1, we see that for $m\leq l<n-1$,
-${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox r^{-\frac{n-l}{2}}_{l}\left(\prod^{l-1}_{i=m-1}{r^{-\frac{1}{2}}_{i}D^{-i+\delta}_{i}}\right)r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-\frac{1}{2}}_{j}D^{\delta}_{j-1}}\left(\prod^{n-2}_{i=l}{r^{-\frac{1}{2}}_{i}}\right)D^{\delta}_{n-1}{\left\|f\right\|}^{2}_{\infty}$
-$\left(\prod^{n-1}_{j=m\ }{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{l-1}_{j=m-1}{D^{-j}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$ (4.3)
-Let ${\gamma}_{m},{\gamma}_{m+1},\dots,{\gamma}_{n}$ be positive real numbers satisfying ${\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n}=1$, and
-${\sigma}_{j}:=\sum^{j}_{i=m}{{\gamma}_{i}},\ \ {\tau}_{j}:=\left\{\begin{array}[]{c}{\sigma}_{j}\mathrm{,\ \ if\ }j\leq n-2\\ 0,\ \ \mathrm{if\ }j=n-1\end{array}\right$.
-Take the geometric mean of (3.41) and (4.3) with weights $\gamma_{i}$,
-${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{j=m}{r^{-\frac{1+{\tau}_{j}}{2}}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{-j\left(1-{\sigma}_{j}\right)+\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$
-Put this bound into (3.40),
-${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{j=m}{r^{\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\tau}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}\prod^{n-1}_{j=m-1}{D^{\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)+O\left(\delta\right)}_{j}}{\left\|f\right\|}^{\frac{2}{p}}_{2}{\left\|f\right\|}^{1-\frac{2}{p}}_{\infty}$
-It just remains to find suitable parameters $p_{i}$ for $m\leq i\leq n$ such that
-$X_{j}:=\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\tau}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)\leq 0,\ \ \mathrm{for\ }m\leq j\leq n-1$
-$Y_{j}:=\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)<0,\ \ \mathrm{for\ }m-1\leq j\leq n-1$
-To find the critical values, we assume that $X_{j}=0$ for $m\leq j\leq n-1$ and $Y_{j}=0$ for $m-1\leq j\leq n-1$. By (3.27),
-$\frac{1}{2}{\left(\frac{1}{2}-\frac{1}{p_{j+1}}\right)}^{-1}-\frac{1}{2}{\left(\frac{1}{2}-\frac{1}{p_{j}}\right)}^{-1}=\frac{1+{\tau}_{j}}{2},\ \ \mathrm{for\ }m\leq j\leq n-1$ (4.4)
-$\frac{1}{2}{\left(\frac{1}{2}-\frac{1}{p_{j+1}}\right)}^{-1}=1+j\left(1-{\sigma}_{j}\right),\ \ \mathrm{for\ }m-1\leq j\leq n-1$ (4.5)
-Putting (4.5) into (4.4),
-$j\left(1-{\sigma}_{j}\right)-\left(j-1\right)\left(1-{\sigma}_{j-1}\right)=\frac{1+{\tau}_{j}}{2},\ \ \mathrm{for\ }m\leq j\leq n-1$ (4.6)
-Thus, ${\gamma}_{m},{\gamma}_{m+1}\dots,{\gamma}_{n-1}$ can be determined recursively. Since the initial step and the recursive steps for $j\leq n-2$ are same as that in [12], we have
-${\gamma}_{j}=\frac{m-1}{2}\frac{1}{j\left(j-1\right)}\prod^{j}_{i=m}{\frac{2i}{2i+1}},\ \ \mathrm{for\ }m\leq j\leq n-2$ (4.7)
-Since the relationship (4.5) and ${\gamma}_{m},{\gamma}_{m+1}\dots,{\gamma}_{n-2}$ are same as that in [12], one has
-$p_{j}=2+\frac{6}{2\left(j-1\right)+\left(m-1\right)\prod^{j-1}_{i=m}{\frac{2i}{2i+1}}}\ ,\ \ \mathrm{for\ }m\leq j\leq n-1$ (4.8)
-The condition $X_{n-1}=0$ is equivalent to
-${\left(\frac{1}{2}-\frac{1}{p_{n}}\right)}^{-1}-{\left(\frac{1}{2}-\frac{1}{p_{n-1}}\right)}^{-1}=1$ (4.9)
-By putting (4.8) into (4.9), one has
-$p_{n}=2+\frac{12}{4n-5+2\left(m-1\right)\prod^{n-2}_{i=m}{\frac{2i}{2i+1}}}$
-By remark 3.2, $m\geq k$. Thus, if we take
-$p>2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}}$
-Then the broad estimate always (2.7) holds. It just remains to check ${\sigma}_{n-1}={\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n-1}\leq 1$. By (4.5),
-${\sigma}_{n-2}={\gamma}_{m}+{\gamma}_{m+1}+\dots+{\gamma}_{n-2}<\frac{m-1}{2}\sum_{j\geq m}{\frac{1}{j\left(j-1\right)}}=\frac{1}{2}$ (4.10)
-Take j=n1j=n-1 in (4.6),
-$\left(n-1\right)\left(1-{\sigma}_{n-1}\right)-\left(n-2\right)\left(1-{\sigma}_{n-2}\right)=1/2\ \Longrightarrow{\gamma}_{n-1}=\frac{1/2-{\sigma}_{n-2}}{n-1}$
-By (4.10),
-${\gamma}_{n-1}<\frac{1}{2\left(n-1\right)}$ (4.11)
-Combine (4.10) and (4.11), we see that ${\sigma}_{n-1}<1$.
+Together with (3.19) we have
+${\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (4.5)
+Lemma (4.2) now follows from (4.4) and (4.5). ∎
+Combining (3.30), (3.31), and Lemma 4.2, we obtain that for $m\leq l\leq n-2$,
+${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{j=m\ }{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{l-1}_{j=m-1}{D^{-j}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$ (4.6)
+Define $p_{j}$ by
+$p_{j}:=\begin{cases}2+\frac{6}{2\left(j-1\right)+\left(m-1\right)\prod^{j-1}_{i=m}{\frac{2i}{2i+1}}}&\text{if $m\leq j\leq n-1$},\\[4.0pt] p(n,k)&\text{if $j=n$}.\end{cases}$ (4.7)
+For $m\leq j\leq n-1$, we define $\gamma_{j}\in[0,1]$ by
+$\gamma_{j}:=\begin{cases}\frac{m-1}{2}\frac{1}{j\left(j-1\right)}\prod^{j}_{i=m}{\frac{2i}{2i+1}}&\text{if $m\leq j\leq n-2$},\\[4.0pt] 1-\sigma_{n-2}&\text{if $j=n-1$}.\end{cases}$ (4.8)
+where
+${\sigma}_{j}:=\sum^{j}_{i=m}{{\gamma}_{i}}$
+Taking the geometric mean of (4.6) with weights $\gamma_{l}$ and (4.1) with weight $\gamma_{n-1}$, we obtain
+${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\\ {r_{n-1}^{-\frac{1}{2}}}D_{n-1}^{-\gamma_{n-1}(n-1)}\left(\prod^{n-2}_{j=m}{r^{-\frac{1+{\sigma}_{j}}{2}}_{j}}\right)\left(\prod^{n-2}_{j=m-1}{D^{-j\left(1-{\sigma}_{j}\right)}_{j}}\right){\left\|f\right\|}^{2\gamma_{n-1}}_{2}{\left\|f\right\|}^{2\sigma_{n-2}}_{\infty}$
+Together with (3.36), we see that
+${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox{r^{\frac{{\beta}_{n}-{\beta}_{n-1}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{n-1}}D^{\frac{{\beta}_{n}}{2}-\left(1+\gamma_{n-1}(n-1)\right)\left(\frac{1}{2}-\frac{1}{p}\right)}_{n-1}\\ \prod^{n-2}_{j=m}{r^{\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\sigma}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}\prod^{n-2}_{j=m-1}{D^{\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}{\left\|f\right\|}^{\frac{2}{q}}_{2}{\left\|f\right\|}^{1-\frac{2}{q}}_{\infty}$
+where $q$ is defined by
+$\frac{1}{q}=\frac{1}{p}+\frac{1}{4n-8}\big(\frac{6}{p}-1\big)$
+Arguing as in [12], we see that
+$\frac{{\beta}_{n}-{\beta}_{n-1}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)=0$
+$\frac{{\beta}_{n}}{2}-\left(1+\gamma_{n-1}(n-1)\right)\left(\frac{1}{2}-\frac{1}{p}\right)<0$
+$\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\sigma}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)=0,\qquad\forall j\in[m,n-2]$
+$\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)=0,\qquad\forall j\in[m-1,n-2]$
+This concludes the proof since $k\leq m$.